### v1

# Exergetic Port-Hamiltonian Systems Modeling Language

Abstract
Mathematical modeling of real-world physical systems requires the consistent combination of a multitude of physical laws and phenomenological models. This challenging task can be greatly simplified by hierarchically decomposing systems. Moreover, the use of diagrams for expressing such decompositions helps make the process more intuitive and facilitates communication, even with non-experts. As an important requirement, models have to respect fundamental physical laws such as the first and the second law of thermodynamics. While some existing modeling frameworks can make such guarantees based on structural properties of their models, they lack a formal graphical syntax. We present a compositional and thermodynamically consistent modeling language with a graphical syntax. As its semantics, port-Hamiltonian systems are endowed with further structural properties and a fixed physical interpretation such that thermodynamic consistency is ensured in a way that is closely related to the GENERIC framework. While port-Hamiltonian systems are inspired by graphical modeling with bond graphs, neither the link between the two, nor bond graphs themselves, can be easily formalized. In contrast, our syntax is based on a refinement of the well-studied operad of undirected wiring diagrams. The language effectively decouples the construction of complex models via the graphical syntax from physical concerns, which are dealt with only at the level of primitive subsystems that represent elementary physical behaviors. As a consequence, reuse of models and substitution of their parts becomes possible. Finally, by construction, systems interact by exchanging exergy, i.e. energy that is available for doing work, so the language is particularly well suited for thermodynamic analysis and optimization.
keywords:
applied category theory, bond graphs, classical electromagnetism, classical mechanics, compositionality, exergy, GENERIC, graphical syntax, local thermodynamic equilibrium, multiphysics, nonequilibrium thermodynamics, port-Hamiltonian systems
journal: arXiv
1 Introduction
In this introduction we want to answer the following questions:
  1. 1.
    What are Exergetic Port-Hamiltonian Systems (EPHS)?
  2. 2.
    How do they enable compositional modeling?
  3. 3.
    Why is this practically relevant?
In addition, we briefly summarize key ideas of nonequilibrium thermodynamics and we discuss related work that inspired our developments. The introduction finishes with an outline of the main part of the paper.
1.1 What?
EPHS provide a compositional and thermodynamically consistent language for expressing mathematical models of multiphysical systems at macroscopic scales. Such systems may combine
  1. 1.
    classical mechanics
    (e.g. spring-mass systems, multibody dynamics, fluid mechanics),
  2. 2.
    classical electromagnetism
    (e.g. LC circuits, electromagnetic wave propagation)
  3. 3.
    and irreversible processes with local thermodynamic equilibrium
    (e.g. friction/viscosity, electrical resistance/conduction, thermal conduction).
The EPHS language is termed compositional since it enables a completely modular and hierarchical specification of models based on its simple graphical syntax, which is discussed in more detail in the next subsection.
If one flattens any hierarchical nesting of systems, one finds that every system is determined by a power-preserving interconnection of primitive subsystems. There are three main classes of primitive systems:
  1. 1.
    Systems that represent energy storage are characterized by a function which yields the stored energy for each possible state of the primitive system.
  2. 2.
    Systems that represent reversible energy exchange are characterized by a so-called Dirac structure, as known e.g. from (constrained) mechanical systems Courant (1990).
  3. 3.
    Systems that represent irreversible energy exchange are characterized by another type of structure that we call Onsager structure.
Since both classical mechanics and electromagnetism can be expressed using Dirac structures, item 2. corresponds to (a) and (b) above, while item 3. corresponds to (c). Structural properties of the primitive systems guarantee that all models respect the first and the second law of thermodynamics, Onsager’s reciprocal relations, as well as further conservation laws (e.g. for mass). Section 1.4 aims for a very brief introduction to nonequilibrium thermodynamics.
1.2 How?
Figure 1: Graphical representation of an interconnection pattern with different syntactic elements shown in a different color. The outer box shown in orange represents the interface of the resulting composite system. The two protruding lines correspond to its ports called $\mathtt{f}$ and $\mathtt{b_{k}}$. The two inner boxes shown in red represent (the interface of) the subsystems called $\mathtt{ke}$ (storage of kinetic energy) and $\mathtt{sa}$ (self-advection of kinetic energy). Both have ports called $\mathtt{p}$ (momentum) and $\mathtt{m}$ (mass), while $\mathtt{sa}$ has another port called $\mathtt{b_{k}}$ (advection of kinetic energy across the boundary of the spatial domain). The port combinator shown in purple combines the ports $\mathtt{p}$ (fluid momentum) and $\mathtt{m}$ (fluid mass) into the multiport $\mathtt{f}$ (fluid). The black dots represent junctions where power is exchanged among the connected ports. The dashed line indicates that port $\mathtt{m}$ of the port combinator does not partake in power exchange, but merely shares information about the state (fluid mass).
Expressions in the graphical syntax are called interconnection patterns because a pattern specifies how a composite system can be formed through a power-preserving interconnection of subsystems. Figure 1 shows an example from fluid dynamics.
The syntax deals with systems only in terms of their interfaces, i.e. their exposed ports. Each port has two attributes: The first attribute gives the physical quantity that can be exchanged via the port together with the corresponding state space. To rule out ill-defined connections, only ports with the same quantity can be connected. The second attribute is of Boolean type and indicates whether power can be exchanged over the port or merely information about system state, as indicated graphically by a solid or a dashed line, respectively.
An interconnection pattern should be thought of as a function which, for every round inner box, takes as input a system with matching interface and returns a system whose interface matches the outer box. For instance, the pattern shown in Figure 1 represents a function which takes two subsystems whose interfaces are determined by the ports of the boxes $\mathtt{ke}$ and $\mathtt{sa}$, and it returns a composite system whose interface is determined by the ports of the outer box. In the graphical representation, we often draw inner boxes filled with a color that indicates the nature of the subsystems that we consider as inputs. In Figure 1, $\mathtt{ke}$ has a blue filling as it represents energy storage, i.e. the fluid’s kinetic energy, while $\mathtt{sa}$ has a green filling as it represents reversible energy exchange, i.e. the advection of momentum and mass by the fluid flow. Written on top of the outer box, kin is an identifier akin to a function name and the text in parenthesis is just a short description. The port combinator and the corresponding multiport $\mathtt{f}$ are syntactic sugar that allows the composite system to be connected to other systems without drawing parallel connections.
(a) Kinetic energy system
(b) Fluid system
(c) EMHD System
Figure 2: Hierarchical specification of an electro-magneto-hydrodynamics model: The fluid model f shown in (b) comprises the kinetic energy system kin shown in (a) and four further models which are not shown, namely an internal energy system int, a thermal conduction model th, and volume / shear viscosity models vol / shr. The electro-magneto-hydrodynamics model emhd shown in (c) comprises the fluid model f and three further models which are not shown, namely an electro-magnetic system em, an electric conduction model el, and a model for the electro-kinetic coupling ek.
Figure 2 demonstrates the hierarchical specification of a model for electro-magneto-hydrodynamics. For simplicity, a spatial domain with empty or isolated boundary is assumed such that only the in-domain interconnection of the different models needs to be dealt with. The yellow filling of the inner boxes in Figures 1(b) and 1(c) indicates that another composite system, which in turn is specified with its own interconnection pattern, is supplied as input. Recursively substituting into all yellow inner boxes the respective interconnection patterns would yield an equivalent model specification without hierarchical decomposition. The composition of interconnection patterns through substitution is uniquely defined as long as interfaces match. Similarly, one can add a level of nesting by refactoring a number of subsystems into a reusable composite subsystem. In the fluid model, one could for instance factor out the kinetic energy system filling the box $\mathtt{kin}$ and the internal energy subsystem filling the box $\mathtt{int}$ into a reusable ideal compressible fluid model.
1.3 Why?
EPHS provide a formal and intuitive modeling language suited for humans and software tools alike. As shown above, increasingly complex models can be assembled from simpler and ultimately primitive parts. This fosters the reusability of models and encourages the conceptual separation of different levels of detail. Beyond the modeling process, this provides value in interdisciplinary communication and education. Some modeling errors can be completely avoided due to typed interfaces (given by ports with fixed physical quantities) and the built-in thermodynamic consistency. Since interconnection patterns can be used as exergy flow diagrams, EPHS models are particularly useful for thermodynamic analysis and optimization. The compositional nature of EPHS could enable a scalable approach to model transformations, such as discretizations, and also to the integration of scientific machine learning techniques in first-principles-based physical modeling.
1.4 Nonequilibrium thermodynamics in a tiny nutshell
Most physical systems are characterized by a huge number of degrees of freedom but keeping track of them is usually neither possible nor useful. Nonequilibrium thermodynamics aims to provide a framework for coarse-grained modeling, meaning that fast dynamics at small scales are not resolved and a quantity called entropy is introduced to account for the missing information, see e.g. Pavelka et al. (2018). Coarse-graining is possible because, despite of fluctuations at the microscopic scale, one can eventually observe non-fluctuating properties at a more macroscopic scale; at least this is the case for isolated systems. For instance, temperature is a property that does not appear at the microscopic scale. Thermodynamic properties are classified as extensive when they are proportional to the size of the system (e.g. entropy, volume, mass) or as intensive when they do not depend on the size of the system (e.g. temperature, pressure, chemical potential), see e.g. Callen (1985)
A system is said to be in local thermodynamic equilibrium if for every material point there exists a neighborhood large enough to observe thermodynamic properties, yet small enough for the properties to be non-varying inside the neighborhood, see e.g. Öttinger (2005). The size of these neighborhoods presents a macroscopic scale at which the dynamics of the thermodynamic properties can be observed and also modeled.
The first law of thermodynamics just restates a well-known fact from mechanics, namely that energy is a conserved quantity. However, instead of considering directly the mechanical and electromagnetic energy associated to microscopic degrees of freedom, a thermodynamic model with local equilibrium resorts to a phenomenological description called internal energy, which is considered to depend on a number of extensive properties, including entropy, and possibly further mechanical or electromagnetic state variables which are observable on the considered macroscopic scale. These quantities then serve as the state variables of a coarse-grained description. Moreover, temperature is the derivative of the internal energy with respect to entropy. Similarly, pressure is given by the derivative with respect to volume and the chemical potential is the derivative with respect to mass, see e.g. Callen (1985).
Classical mechanics and electromagnetism exhibit reversible dynamics, meaning that one could in principle arrive again at the initial condition by reversing the direction of time (although for many systems this would be impossible due to chaos). Incomplete knowledge about the microscopic state immediately leads to irreversible dynamics and an irreversible dynamics in turn leads to a growth of entropy, as information (e.g. about the initial state of the system) is lost in the process. This is reflected by the second law of thermodynamics which states that the entropy of an isolated system never decreases and that the system evolves towards a state of maximum entropy, see e.g. Öttinger (2005). In other words, the uncertainty about the microscopic state can only grow, not decrease.
For models with local equilibrium, the irreversible dynamics can be expressed in terms of so-called thermodynamic fluxes that determine instantaneous changes of coarse-grained state variables. The fluxes are expressed as a function of thermodynamic forces, which are given by differences in intensive quantities, see e.g. Öttinger (2005); de Groot and Mazur (1984). For instance, temperature differences cause heat flux, according to Fourier’s law, and differences in chemical potential cause a diffusive mass flux, according to Fick’s law. Further, one may observe cross effects such as a diffusive mass flux caused by temperature differences (Soret effect) and likewise a heat flux caused by differences in chemical potential (Dufour effect). Onsager’s reciprocal relations state that the relation between fluxes and forces must possess a certain symmetry Onsager (1931). For instance, the ratio of the mass flux caused by the Soret effect and the respective temperature difference is equal to the ratio of the heat flux caused by the Dufour effect and the respective difference in chemical potential.
1.5 Related work
Figure 3 gives a high-level view of this subsection.
Figure 3: Informal Venn diagram illustrating that EPHS integrates ideas from graphical and energy-based modeling of physical systems with bond graphs, the GENERIC framework for nonequilibrium thermodynamics, the port-Hamiltonian theory of open dissipative systems (PHS), and applied category theory (ACT) research on the formalization of graphical languages and compositional dynamical systems.
1.5.1 Port-Hamiltonian systems
Apparent from their name, EPHS have a lot in common with Port-Hamiltonian Systems (PHS). PHS were originally inspired by relating the dynamical systems defined by generalized bond graphs Breedveld (1982) (i.e. bond graphs using a single kind of energy storage elements together with gyrator elements) to Hamiltonian mechanics Maschke (1991). While many authors make use of explicit input-output port-Hamiltonian systems, PHS constitute an acausal systems theory, inspired also by the behavioral approach of Willems Willems (1974, 2007). The dynamics of networks comprising energy storing elements and reversible coupling elements between them (such as generalized transformers and gyrators) can be formalized, from a geometric viewpoint, as a Hamiltonian system whose Poisson structure reflects the network topology Maschke et al. (1992). The generalization to networks with algebraic constraints and externally controlled ports is achieved by using Dirac structures, which unify Poisson structures and presymplectic structures van der Schaft and Maschke (1995). By specifying a resistive relation on external port variables, the implicit Hamiltonian dynamics can be combined with a dissipative gradient dynamics, allowing for network elements that represent e.g. dampers and resistors. The key properties of PHS are firstly a power-balance equation which states that the sum of the stored power and the dissipated power is equal to the supplied power (passivity) and secondly the possibility to combine port-Hamiltonian systems by composition of their Dirac structures with an interconnecting Dirac structure, see for instance van der Schaft and Jeltsema (2014).
While the dynamical systems defined by EPHS models are essentially PHS, they have further structural properties and a fixed physical interpretation. PHS in general have no uniform thermodynamic interpretation and they lack a formal graphical syntax to express their interconnection. In contrast to EPHS, PHS also lack a built-in mechanism for referring to subsystems and ports.
1.5.2 GENERIC framework
The general equation for reversible-irreversible coupling (GENERIC) provides a framework for combining reversible and irreversible dynamics in a way that respects fundamental thermodynamic principles, namely the first law, the second law and the reciprocal relations discovered by Onsager Grmela and Öttinger (1997). The reversible part is defined by an entropy-conserving Hamiltonian dynamics, while the irreversible part is given by an energy-conserving gradient flow, see e.g. Öttinger (2005). We note that the same holds for closely-related earlier developments in this direction, including metriplectic systems, see Kaufman (1984); Grmela (1984a, b); Morrison (1984). While a large class of well-known dynamical systems from physics can be identified as instances of the GENERIC, see e.g. Pavelka et al. (2018), the framework is especially helpful for the development of new thermodynamic models.
The combination of Hamiltonian and gradient dynamics in both PHS and GENERIC systems led to the initial observation of a relationship between the two frameworks in Badlyan et al. (2018). GENERIC systems with local thermodynamic equilibrium can be seen as port-Hamiltonian systems, where the storage function corresponds to the system’s exergy. This function yields the maximum amount of work that can be extracted by an ideal (i.e. purely reversible) process which may arbitrarily interact with the system’s storage components and a given reference environment Lohmayer et al. (2021).
The dynamical systems defined by (isolated) EPHS models are basically GENERIC systems, since their built-in thermodynamic consistency and associated physical interpretation follow from the same ideas. Beyond the aspect of compositionality, there are two main differences: 1. GENERIC systems are more general in the sense of including systems beyond local thermodynamics equilibrium (where the concept of temperature does not apply). 2. The reversible part of EPHS is defined using Dirac structures, rather than less general Poisson structures (which cannot express kinematic constraints). This is crucial e.g. for modeling multibody systems and electrical circuits.
1.5.3 Bond graph modeling
Bond graph modeling is based on a graphical notation for representing dynamical systems and in particular multi-domain models involving mechanical and electrical elements Paynter (1961); Borutzky (2010). The main idea is that elements interact via bonds that represent links for energy exchange. Associated to each bond, there are two dynamic variables called flow and effort, whose product (or duality pairing) gives the exchanged power. Bonds can be connected via two types of junctions, namely 00-junctions, where the flow variables of all connected ports balance, with all effort variables being equal, and dually defined $1$-junctions, where the effort variables balance, with all flow variables being equal. In both cases, the net power at a junction is zero.
While PHS can be seen as a geometric formalization of the dynamical systems defined by generalized bond graphs, the algebraic formalization of the bond graph notation itself is not straightforward, when compared to e.g. standard multigraphs, see Birkett and Roe (1989); Coya (2017). The inherent complexity of the notation is evident also in practice, since the direction in which exchange of power is counted positive and sometimes also the question which variable is flow and which is effort have to be decided arbitrarily, bond by bond. While bond graphs are frequently used in the literature to illustrate PHS, their use as a formal syntax for model specification is not established. Previous work has focused on algorithms for converting certain classes of bond graphs to port-Hamiltonian systems, see e.g. Pfeifer et al. (2020).
In contrast to bond graph notation, the EPHS syntax only has 00-junctions and every interconnection pattern has an outer box. As demonstrated by various applications, a single type of junctions is sufficient and moreover the simplification entails crucial advantages: Interconnection patterns are easier to interpret, in particular due to the direct correspondence between junctions and energy domains. This correspondence arises from the inherent distinction between the roles of flows and efforts, a feature that is lacking in bond graphs. Further, the direction in which exchange of power is counted positive is predetermined in accordance with the convention used in thermodynamics, i.e. energy supplied to a system has a positive sign. Last but not least, interconnection patterns can be hierarchically composed and decomposed in a natural way.
1.5.4 Applied category theory
The compositional nature of EPHS can be made precise using category theory. Category theory was first applied in algebraic geometry and topology and now is used to organize many branches of mathematics Mac Lane (1998). In recent decades, category theory has also grown beyond pure math to organize scientific and engineering disciplines as well, giving birth to ‘applied category theory’ Fong and Spivak (2019). A specific approach to categorical systems theory Myers (2022) that has inspired this work is based on the following principles, which can essentially also be found in the work of Willems Willems (2007).
  1. 1.
    A shift of focus from individual systems to how systems relate to each other.
  2. 2.
    An isolated system is a system with empty interface.
  3. 3.
    The description of systems involves syntax and semantics. Syntax typically consists of combinatorial data that can be manipulated computationally, while semantics are given by geometric objects that exist only platonically. Additionally, the syntax can be displayed and manipulated diagrammatically as well as in text. An explicit understanding of the translation from syntax to semantics serves as a guide for computer implementation.
For instance, there is a line of research which formalizes graphical languages for systems and process such as block diagrams Baez and Erberle (2015), Petri nets Baez and Pollard (2017), junctions of bond graphs Coya (2017), and passive linear circuits Baez and Fong (2018). These works consider a graphical syntax as some sort of string diagrams, established via a generators and relations approach.
More closely related to our work, another line of research uses the theory of operads to define a graphical syntax which supports hierarchical (de-)composition, see Baez et al. (2020); Foley et al. (2021). Operads organize formal operations with finitely many inputs (subsystems) and one output (composite system), see Leinster (2004); Yau (2016). As proposed in Lohmayer and Leyendecker (2022), the EPHS syntax, is based on the operad of undirected wiring diagrams, which initially appeared in Spivak (2013), together with its relational semantics used to express database queries, see also Spivak (2014); Yau (2018); Breiner et al. (2020). In a similar lane, directed wiring diagrams are used to formalize nested systems of Moore machines and ODEs with outputs, see Vagner et al. (2015); Schultz et al. (2019). A combination of dynamical systems based on directed and undirected wiring diagrams is presented in Libkind (2020); Libkind et al. (2021). The composition of lossless port-Hamiltonian systems was recently studied from a categorical perspective in Lynch (2022).
1.6 Outline
Section 2 defines system interfaces. Based on this, Section 3 defines interconnection patterns. Section 4 shows how every interconnection pattern implies a relation between the port variables of its interfaces. Section 5 discusses how exergy quantifies the energy stored in a system that is available for doing work. Based on this, Section 6 introduces the primitive systems. Section 7 defines composite systems and how they are formed using the graphical syntax. Section 8 shows that EPHS models are thermodynamically consistent. Section 9 presents an example of a motor model, together with a comparison to a classical bond graph. Section 10 concludes with a discussion.
2 Interfaces
While the narrative of discovery, usually proceeds from systems to interfaces, logically speaking, the syntax needs to be introduced before the semantics and hence we start with interfaces. Graphically, the interface of a system is a box with a set of ports drawn around it. While we could treat such sets abstractly, as is often done in applied category theory, for practical reasons we instead treat them as sets of names called namespaces.
2.1 Namespaces
Definition 2.1.
A string is a non-empty list of characters and a name is a list of strings. ∎
We write names using a monospace font and with strings separated by dots. For instance, the name consisting of the list of strings “oscillator”, “spring”, “q” is written as $\mathtt{oscillator.spring.q}$. Further, we use [][][ ] to denote the empty name, i.e. the empty list of strings. Given two names $n$ and $n^{\prime}$, we denote their concatenation by $n\ast n^{\prime}$. For example, $\mathtt{a.b}\ast\mathtt{c}=\mathtt{a.b.c}$.
Definition 2.2.
Given two names $p$ and $n$, we say that $p$ is a prefix of $n$, and write $p\leq n$ if there exists any name $s$ such that $n=p\ast s$. In case $s\neq[]$, we say that $p$ is a strict prefix of $n$ and write $p<n$. ∎
Example 2.3.
It holds that $\mathtt{a.b}<\mathtt{a.b.c}$ and $\mathtt{a.b}\leq\mathtt{a.b}$. Also $[]<\mathtt{a}$. It does not hold that $\mathtt{a.a}\leq\mathtt{a}$. ∎
Definition 2.4.
A set of names $N$ is prefix-free if for all $n,n^{\prime}\in N$, $n$ is not a strict prefix of $n^{\prime}$. A namespace is a finite, prefix-free set of names. Let $\mathbf{Namespace}$ refer to the set of all namespaces. ∎
Example 2.5.
The set of names $\{\mathtt{a},\,\mathtt{b}\}$ is a namespace and so is $\{\mathtt{a.a},\,\mathtt{a.b}\}$. Since $\mathtt{a}<\mathtt{a.b}$, the set of names $\{\mathtt{a},\,\mathtt{a.b}\}$ is not prefix-free and hence not a namespace. ∎
Sets of names are equivalent to prefix trees, called tries in computer science Fredkin (1960).
Example 2.6.
The namespace $\{\mathtt{a.b.b},\,\mathtt{a.b.c},\,\mathtt{b.a},\,\mathtt{b.y},\,\mathtt{c}\}$ corresponds to the following trie:
[Uncaptioned image][Uncaptioned $\qed$
Note that for a namespace $N$, the leaf nodes of its trie correspond to the names in $N$, while for a set of names, which is not prefix-free, at least one name corresponds to an internal node. In the following, we think of namespaces both as sets of names and as tries.
Given several namespaces, we could take their union as sets. However, this runs the risk of name conflicts, as it is possible for a given name in the union to have come from either of the original sets. It is for this reason that we introduce the concept of a named sum, which can be thought of as a practical implementation of disjoint union.
Definition 2.7.
Given a namespace $N$, and a function $f\colon N\to\mathbf{Namespace}$, the named sum of $f$ is
$\sum f\>=\>\Bigl{\{}n\ast a\,\mid\,n\in N,\>a\in f(n)\Bigr{\}}\,.\qed$
We also sometimes write expressions like $\sum\left[\mathtt{a}\mapsto I_{1},\,\mathtt{b}\mapsto I_{2}\right]$, which refers to the named sum of the anonymous function defined by
$x\>\mapsto\>\begin{cases}I_{1}&\text{if }x=\mathtt{a}\\ I_{2}&\text{if }x=\mathtt{b}\,.\end{cases}$
Example 2.8.
Given $N=\{\mathtt{q},\,\mathtt{p.1},\,\mathtt{p.2}\}$, we have
$\sum\left[\mathtt{l}\mapsto N,\,\mathtt{s.t}\mapsto N\right]\>=\>\bigl{\{}\mathtt{l.q},\,\mathtt{l.p.1},\,\mathtt{l.p.2},\,\mathtt{s.t.q},\,\mathtt{s.t.p.1},\,\mathtt{s.t.p.2}\bigr{\}}\,.\qed$
Proposition 2.9.
Given a namespace $N$ and a function $f\colon N\to\mathbf{Namespace}$, the named sum $\sum f$ is again a namespace, that is to say it is prefix-free. ∎
Proof.
The trie $\sum f$ is simply the trie $N$ with, for each leaf node $n\in N$, the trie $f(n)$ grafted on top. ∎
Namespaces and named sums provide a simple and effective formalism for working with large hierarchically-structured systems.
2.2 Packages
An interface has a namespace, where we interpret each name as referring to a port. However, this is not the only ingredient to an interface, as each port has two associated attributes. This motivates the definition of packages:
Definition 2.10.
Given a set $A$, we define
$\mathbf{Pkg}(A)\>=\>\Bigl{\{}(N,\,\tau)\,\mid\,N\in\mathbf{Namespace},\,\tau\colon N\to A\Bigr{\}}\,.$
An element $(N,\,\tau)\in\mathbf{Pkg}(A)$ is called a package of elements of $A$. We also just write $\tau\in\mathbf{Pkg}(A)$ with $N=\operatorname{dom}(\tau)$. ∎
The name ‘package’ is drawn from software engineering; one might consider a software package to be a ‘package of types and procedures’.
A package $(N,\,\tau)\in\mathbf{Pkg}(A)$ can be seen as a trie $N$ with an element $\tau(n)\in A$ associated to each leaf node $n\in N$. Hence, the concept of named sum carries over from $\mathbf{Namespace}$ to $\mathbf{Pkg}(A)$, for a fixed set $A$. Named sum can then be seen as a flattening operation, analogous to flattening a list of lists. Specifically, for the set of all lists of elements of $A$
$\operatorname{\mathbf{List}}(A)\>=\>\bigl{\{}(n,\,f)\,\mid\,n\in\mathbb{N},\,f\colon\{1,\,\ldots,\,n\}\to A\bigr{\}}\,,$
flattening is a map $\operatorname{\mathbf{List}}(\operatorname{\mathbf{List}}(A))\to\operatorname{\mathbf{List}}(A)$, which for instance sends [[1, 2],[3, 4]]1234[[1,\,2],\,[3,\,4]][ [ 1 , 2 ] , [ 3 , 4 ] ] to [1, 2, 3, 4]1234[1,\,2,\,3,\,4][ 1 , 2 , 3 , 4 ]. Similarly, named sum is a map $\mathbf{Pkg}(\mathbf{Pkg}(A))$ to $\mathbf{Pkg}(A)$. An input to this is a trie $N$, which has associated to each leaf node $n\in N$ another trie $N_{n}$, whose leaves carry elements of $A$. The named sum operation then simply grafts onto each node $n\in N$ its associated trie $N_{n}$, carrying along the elements of $A$ at the leaves. In terms of category theory, the flattening operation makes $\operatorname{\mathbf{List}}$ and $\mathbf{Pkg}$ examples of monads.
2.3 Bundle of quantities
The first attribute associated to any port represents the physical quantity that can be exchanged via the port, e.g. entropy or momentum, along with a state space in which values of the quantity live. For instance a momentum variable can take values in $\mathbb{R}$ if motion along a single axis is considered or it can take values in the vector space $\mathfrak{se}^{*}(3)$ if rigid body motions in $3$-dimensional Euclidian space are considered, etc.
Definition 2.11.
Let $\mathbf{S}$ be a finite set of state spaces (i.e. smooth manifolds) and for each space $\mathcal{X}\in\mathbf{S}$, let $\mathbf{Q}_{\mathcal{X}}$ be a finite set of physical quantities taking values in $\mathcal{X}$. Further, let $\mathbf{Q}$ be the disjoint union of all quantities, i.e.
$\mathbf{Q}\>=\>\bigsqcup_{\mathcal{X}\in\mathbf{S}}\mathbf{Q}_{\mathcal{X}}\,.$
This forms a bundle of quantities (in the category of finite sets), whose projection
$\mathrm{space}:\mathbf{Q}\twoheadrightarrow\mathbf{S}$
returns the underlying space of a quantity. ∎
Example 2.12.
The bundle of quantities given by
$\begin{split}\mathbf{S}\>&=\>\{\mathbb{R},\,\mathrm{SE}(3),\,\mathfrak{se}^{*}(3)\}\\ \mathbf{Q}_{\mathbb{R}}\>&=\>\{\mathtt{displacement},\,\mathtt{momentum},\,\mathtt{entropy},\,\mathtt{volume}\}\\ \mathbf{Q}_{\mathrm{SE}(3)}\>&=\>\{\mathtt{pose}\}\\ \mathbf{Q}_{\mathfrak{se}^{*}(3)}\>&=\>\{\mathtt{momentum}\},\end{split}$
would be sufficient for some spatially-lumped models from mechanics and thermodynamics. A port with associated quantity $q=(\mathbb{R},\,\mathtt{entropy})\in\mathbf{Q}$ can exchange entropy and its associated state space is $\mathrm{space}(q)=\mathbb{R}\in\mathbf{S}$. ∎
To distinguish reversible and irreversible dynamics later on in Sections 6.3 and 6.4, we make use of the concept of parity, see Pavelka et al. (2014); Lohmayer et al. (2021) for details.
Definition 2.13.
Let the parity of each quantity with respect to time-reversal transformation be given by a function $\mathrm{P}:\mathbf{Q}\rightarrow\{-1,\,+1\}$. ∎
Example 2.14.
For the bundle of quantities from 2.12, we have
$\mathrm{P}(\mathbb{R},\,\mathtt{displacement})=+1$
$\mathrm{P}(\mathbb{R},\,\mathtt{momentum})=-1$
$\mathrm{P}(\mathbb{R},\,\mathtt{entropy})=+1$
$\mathrm{P}(\mathbb{R},\,\mathtt{volume})=+1$
because displacement, entropy, and volume do not change instantaneously when a recording is suddenly played backwards (positive parity), whereas momentum changes its sign (negative parity). ∎
A comprehensive collection of physical quantities with their associated parities would be a built-in feature of a software implementation.
2.4 Interfaces as packages
Each port has two attributes, namely the already discussed quantity and further a Boolean attribute that indicates whether the port is a power port or a state port. Only power ports allow for exchange of energy associated with the attributed quantity. A state port is required whenever the behavior of a system depends on the state of another system, although no energy is exchanged between the two.
Definition 2.15.
An interface is a package of elements of $\mathbf{Q}\times\{\mathsf{p},\mathsf{s}\}$, i.e. let
$\mathbf{Interface}\>=\>\mathbf{Pkg}\bigl{(}\mathbf{Q}\times\{\mathsf{p},\mathsf{s}\}\bigr{)}\,.\qed$
Example 2.16.
Let $I=(N,\,\tau)$ be an interface with namespace $N=\{\mathtt{p_{1}},\,\mathtt{p_{2}},\,\mathtt{q}\}$ and attributes given by $\tau(\mathtt{p_{1}})=\tau(\mathtt{p_{2}})=\left((\mathbb{R},\,\mathtt{momentum}),\,\mathsf{p}\right)$ and $\tau(\mathtt{q})=\left((\mathbb{R},\,\mathtt{displacement}),\,\mathsf{s}\right)$. Ports $\mathtt{p_{1}}$ and $\mathtt{p_{2}}$ are power ports that can exchange kinetic energy, along with information about the current momentum. Port $\mathtt{q}$ is a state port that can only exchange information about the current displacement. ∎
We note that the named sum of interfaces is a function $\mathbf{Pkg}(\mathbf{Interface})\to\mathbf{Interface}$.
2.5 Bundles of port variables
A state port has one port variable called state. A power port has two further port variables called flow and effort. Given an interface, the port variables of all its ports combined together take values in a vector bundle, called the bundle of port variables associated to the interface.
Definition 2.17.
Let $I=(N,\,\tau)$ be an interface and let $p\in I$ be a port, i.e. $p\in N$.
The state space associated to port $p$ is
$\mathcal{X}_{p}\>=\>(\mathrm{space}\circ\mathrm{pr}_{1}\circ\tau)(p)\,,$
where $\mathrm{pr}_{1}$ projects onto the first attribute, i.e. the quantity.
The state space associated to interface $I$ is the Cartesian product
$\mathcal{X}_{I}\>=\>\prod_{p\,\in\,I}\mathcal{X}_{p}\,.\qed$
Example 2.18.
Let $I_{\text{gas}}=(N_{\text{gas}},\,\tau_{\text{gas}})$ be an interface with namespace $N_{\text{gas}}=\{\mathtt{s},\,\mathtt{v}\}$ and attributes given by $\tau_{\text{gas}}(\mathtt{s})=\left((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p}\right)$, $\tau_{\text{gas}}(\mathtt{v})=\left((\mathbb{R},\,\mathtt{volume}),\,\mathsf{p}\right)$. Then, its associated state space is $\mathcal{X}_{I_{\text{gas}}}=\mathcal{X}_{\mathtt{s}}\times\mathcal{X}_{\mathtt{v}}=\mathbb{R}\times\mathbb{R}$. Let $I_{\text{mass}}=(N_{\text{mass}},\,\tau_{\text{mass}})$ be another interface with $N_{\text{mass}}=\{\mathtt{p}\}$ and $\tau_{\text{mass}}(\mathtt{p})=\left((\mathbb{R},\,\mathtt{momentum}),\,\mathsf{p}\right)$. Then, the named sum
$I\>=\>(N,\,\tau)\>=\>\sum[\mathtt{gas}\mapsto I_{\text{gas}},\,\mathtt{mass}\mapsto I_{\text{mass}}]$
has the namespace $N=\{\mathtt{gas.s},\,\mathtt{gas.v},\,\mathtt{mass.p}\}$ with attributes given by $\tau(\mathtt{gas.s})=\left((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p}\right)$, $\tau(\mathtt{gas.v})=\left((\mathbb{R},\,\mathtt{volume}),\,\mathsf{p}\right)$, $\tau(\mathtt{mass.p}))=\left((\mathbb{R},\,\mathtt{momentum}),\,\mathsf{p}\right)$. We note that
$\mathcal{X}_{I}\>=\>\mathcal{X}_{I_{\text{gas}}}\times\mathcal{X}_{I_{\text{mass}}}\,.\qed$
Given any interface $I$, each port $p\in I$ has an associated state variable $p\mathtt{.x}\in\mathcal{X}_{p}$. If $p$ is a power port, it additionally has a flow variable $p\mathtt{.f}$ and an effort variable $p\mathtt{.e}$. All port variables implicitly depend on time. As detailed in Section 6.1, the flow variable is the instantaneous rate of change of the state variable and the effort variable is the differential of a function of the state variable. Hence, we have $\left(p\mathtt{.x},\,p\mathtt{.f},\,p\mathtt{.e}\right)\in\mathbb{T}\mathcal{X}_{p}$, where $\mathbb{T}\mathcal{X}_{p}=\mathrm{T}\mathcal{X}_{p}\oplus\mathrm{T}^{*}\mathcal{X}_{p}$ denotes the direct sum (or Whitney sum) of the respective tangent and cotangent bundles (see e.g. Lee (2012)). The flow and effort variables are called power variables because their duality pairing $\langle p\mathtt{.e}\mid p\mathtt{.f}\rangle\in\mathbb{R}$ yields the instantaneous power that is exchanged through port $p$.
Now we can define the bundle of port variables associated to an interface:
Definition 2.19.
Let $I$ be an interface. Based on the Boolean attribute of its ports, we can write its associated state space as the Cartesian product $\mathcal{X}_{I}\cong\mathcal{X}_{I,p}\times\mathcal{X}_{I,s}$, where $\mathcal{X}_{I,p}$ is the state space of all power ports and $\mathcal{X}_{I,s}$ is the state space of all state ports. Accordingly, we let $x=(x_{p},\,x_{s})\in\mathcal{X}_{I,p}\times\mathcal{X}_{I,s}$ denote all state variables and we let $(f,\,e)\in\mathbb{T}_{x_{p}}\mathcal{X}_{I,p}$ denote all flow and effort variables. The bundle of port variables $\mathrm{pr}_{I}:\mathcal{P}_{I}\twoheadrightarrow\mathcal{X}_{I}$ associated to interface $I$ is defined as the pullback bundle of $\mathbb{T}\mathcal{X}_{I,p}$ along the projection $\mathrm{pr}_{p}:\mathcal{X}_{I,p}\times\mathcal{X}_{I,s}\twoheadrightarrow\mathcal{X}_{I,p}$. Hence, $\mathcal{P}_{I}\cong\mathbb{T}\mathcal{X}_{I,p}\times\mathcal{X}_{I,s}$.
{$\begin{tikzcd}\qed$
Example 2.20.
Consider again the interface from 2.16. Its bundle of port variables $\mathcal{P}_{I}\twoheadrightarrow\mathcal{X}_{I}$ has total space (or port space)
$\mathcal{P}_{I}\>=\>\mathbb{T}\mathbb{R}\times\mathbb{T}\mathbb{R}\times\mathbb{R}\>\ni\>\bigl{(}(\mathtt{p_{1}.x},\,\mathtt{p_{1}.f},\,\mathtt{p_{1}.e}),\,(\mathtt{p_{2}.x},\,\mathtt{p_{2}.f},\,\mathtt{p_{2}.e}),\,\mathtt{q.x}\bigl{)}$
and base space (or state space)
$\mathcal{X}_{I}\>=\>\mathbb{R}\times\mathbb{R}\times\mathbb{R}\>\ni\>\left(\mathtt{p_{1}.x},\,\mathtt{p_{2}.x},\,\mathtt{q.x}\right).\qed$
3 Interconnection patterns
As introduced in Section 1.2, an interconnection pattern expresses how a composite system can be formed from a finite number of subsystems. The syntax defines the entirety of such expressions and their composition, i.e. their hierarchical nesting.
3.1 Interconnection patterns
Figure 4: Interconnection pattern of an isothermal oscillator model. The box $\mathtt{pkc}$ represents the coupling of the potential energy domain represented by the junction on its left and the kinetic energy domain represented by the junction on its right. The box $\mathtt{env}$ represents the isothermal environment which absorbs the heat that is generated by the damping. The port $\mathtt{p}$ of the outer box exposes the kinetic energy domain, allowing for an external forcing of the oscillator model.
Figure 4 shows an interconnection pattern. The mathematical content of such a diagram is captured by the following definition:
Definition 3.1.
An interconnection pattern $(\beta,\,I,\,J)$ is defined by the following data:
  1. 1.
    A package of interfaces $\beta\in\mathbf{Pkg}(\mathbf{Interface})$. Names in $\operatorname{dom}(\beta)$ refer to inner boxes, $\sum\beta$ is called inner interface and names in $\operatorname{dom}(\sum\beta)$ refer to inner ports.
  2. 2.
    An interface $I$ called outer interface with ports called outer ports.
  3. 3.
    A partition $J$ of the combined interface $P=\sum\left[\mathtt{inner}\mapsto\sum\beta,\,\mathtt{outer}\mapsto I\right]$ such that each subset or junction $j\in J$ contains
    • (a)
      at least one inner port,
    • (b)
      at most one outer port,
    • (c)
      and all ports have the same associated quantity. ∎
Although the names of inner ports and outer ports are distinguished by a prefix, usually the namespaces of $\sum\beta$ and $I$ are already disjoint. We then abuse notation and simply write for instance $\mathtt{p}$ rather than $\mathtt{outer.p}$.
We graphically depict interconnection patterns as follows. Each element $b\in\operatorname{dom}(\beta)$ is the name of an inner box, which we draw as a circle. The ports of its interface $\beta(b)$ are drawn as lines emanating from the box. The ports of the outer interface $I$ are drawn as lines to the surrounding outer box. Each subset $j\in J$ corresponds to a junction drawn as a black dot. The elements of $j\subseteq P$ are precisely the connected ports. To avoid confusion, we recall that the identifier isothermal_osc written on the top left of the outer box refers to the entire pattern (akin to a function name), rather than just to the outer box (which similar to the return value of a function does not need a name).
The pattern from Figure 4 is redrawn in Figure 5, firstly to emphasize that the color of the inner boxes is merely an annotation hinting at the nature of the considered subsystems and secondly to introduce an abbreviated way of writing port names.
Figure 5: Interconnection pattern of an isothermal oscillator model. Whenever all ports connected to a certain junction have the same name we write the name only once at the junction.
Example 3.2.
For the pattern $(\beta,\,I,\,J)$ depicted in Figure 5, the namespace of inner boxes is
$\operatorname{dom}(\beta)\>=\>\{\mathtt{spring},\,\mathtt{pkc},\,\mathtt{mass},\,\mathtt{damping},\,\mathtt{env}\}\,.$
The function $\beta:\operatorname{dom}(\beta)\rightarrow\mathbf{Interface}$ is given by
$\begin{split}\mathtt{spring}\>&\mapsto\>(\{\mathtt{q}\},\,\tau_{\text{spring}})\\ \mathtt{pkc}\>&\mapsto\>(\{\mathtt{q},\,\mathtt{p}\},\,\tau_{\text{pkc}})\\ \mathtt{mass}\>&\mapsto\>(\{\mathtt{p}\},\,\tau_{\text{mass}})\\ \mathtt{damping}\>&\mapsto\>(\{\mathtt{p},\,\mathtt{s}\},\,\tau_{\text{damping}})\\ \mathtt{env}\>&\mapsto\>(\{\mathtt{s}\},\,\tau_{\text{env}})\end{split}$
with $\tau_{\text{spring}}(\mathtt{q})=((\mathbb{R},\,\mathtt{displacement}),\,\mathtt{p})$, etc. Hence, the namespace of the inner interface $\sum\beta$ is
$\operatorname{dom}(\textstyle\sum\beta)\>=\>\{\mathtt{spring.q},\,\mathtt{pkc.q},\,\mathtt{pkc.p},\,\mathtt{mass.p},\,\mathtt{damping.p},\,\mathtt{damping.s},\,\mathtt{env.s}\}\,.$
The namespace of the outer interface $I$ is $\{\mathtt{p}\}$. The partition $J$ is given by
$\bigl{\{}\{\mathtt{spring.q},\,\mathtt{pkc.q}\},\,\{\mathtt{mass.p},\,\mathtt{pkc.p},\,\mathtt{damping.p},\,\mathtt{p}\},\,\{\mathtt{env.s},\,\mathtt{damping.s}\}\bigr{\}}\,.\qed$
We may also denote an interconnection pattern using the condensed notation
[…]
In this we implicitly let $\operatorname{dom}(\beta)=\{\mathtt{I_{1}},\,\ldots,\,\mathtt{I_{n}}\}$ and $\beta(\mathtt{I_{i}})=I_{i}$. The left arrow is a surjection because each junction has at least one connected inner port and the right arrow is an injection because each junction has at most one connected outer port. Interconnection patterns are a special case of undirected wiring diagrams Spivak (2013), which do not have the injectivity and surjectivity conditions.
As shown in Figure 1, interconnection patterns may have another feature that is not formalized by 3.1, namely port combinators for combining ports into multiports. Although namespaces seem to provide an adequate basis to formalize multiports and port combinators, we defer this to later work. At this point, we simply regard them as ad-hoc syntactic sugar for expressing parallel connections in a tidy and efficient manner.
3.2 Composition of interconnection patterns
Whenever the outer interface of one pattern and the interface corresponding to an inner box of another pattern are equivalent up to renaming of the ports, the former pattern can be substituted into the latter, with no further data required. This is called composition. Having a compositional syntax facilitates dealing with complex systems, as it allows to encapsulate subsystems, which can be easily reused and replaced, as long as interfaces match. As an example, the isothermal oscillator model can be refined into a nonisothermal oscillator model by substituting a more detailed damper model.
Figure 6: Interconnection pattern of a nonisothermal damper model. The box $\mathtt{mf}$ represents a mechanical friction model connected to the kinetic energy domain represented by the outer port $\mathtt{p}$. The box $\mathtt{tc}$ represents a thermal capacity, which can store the heat generated by friction. The box $\mathtt{ht}$ represents a heat transfer model, which allows the thermal capacity to reach equilibrium with the thermal energy domain represented by the outer port $\mathtt{s}$.
Figure 7: Interconnection pattern of a nonisothermal oscillator model.
Substituting the pattern in Figure 6 into the pattern in Figure 5, yields the pattern in Figure 7. Graphically, the former pattern is inserted into the inner box called $\mathtt{damping}$ of the latter pattern. In this step, the named sum for interfaces is used. Then, the common interface is deleted and for each of its ports, the two assigned junctions are identified.
Category theory provides a language to talk about composition and it includes well-behavedness conditions, which ensure that composition works in the way we expect, just as we need not worry about the difference between $(1+2)+1$ and $1+(2+1)$. In a following paper, we will formalize this using the language of operads.
4 Semantics of interconnection patterns
In this section, we discuss how to interpret an interconnection pattern as imposing a relation between the port variables of its inner and outer interface. We do this in two parts. In the first part, we talk about the relation as a collection of equations and discuss its physical interpretation. In the second part, we reify it as a geometric object, namely as a relation between bundles of port variables.
4.1 Interconnection patterns as equations
Informally, the semantics of an interconnection pattern $(\beta,\,I,\,J)$ can be seen as a collection of equations. Specifically, for each junction $j\in J$, we have the following:
Let $j_{p}\subseteq j$ be the subset of connected power ports.
  • 1.
    Equality of state. For all ports $p,p^{\prime}\in j$,
    $p\mathtt{.x}\>=\>p^{\prime}\mathtt{.x}$
  • 2.
    Equality of effort. For all power ports $p,p^{\prime}\in j_{p}$,
    $p\mathtt{.e}\>=\>p^{\prime}\mathtt{.e}$
  • 3.
    Equality of net flow.
    $\sum_{p\,\in\,j_{p}\cap\sum\beta}p\mathtt{.f}\>=\>\sum_{p\,\in\,j_{p}\cap I}p\mathtt{.f}$
As detailed in Section 7, combining these equations with equations that determine the semantics of each of the subsystems filling the inner boxes yields a collection of equations that determines the semantics of the resulting composite system.
Since all ports connected to a junction have the same quantity, their associated state variables take values in the same space, making equality of state a well-typed equation. As a consequence of equality of state, all effort variables take values in the same fiber of the same cotangent bundle, making equality of effort also well-typed. Similarly, all flow variables take values in the same fiber of the same tangent bundle, making equality of net flow well-typed.
Also note that in the equality of net flow, the inner ports are on the left hand side and the outer ports are on the right hand side. As a consequence, at every junction $j\in J$ the following power balance equation is satisfied:
$\sum_{p\,\in\,j_{p}\cap\sum\beta}\langle p\mathtt{.e}\mid p\mathtt{.f}\rangle\>=\>\sum_{p\,\in\,j_{p}\cap I}\langle p\mathtt{.e}\mid p\mathtt{.f}\rangle$
Summing over all junctions on both sides, we get that the net power supplied to all subsystems is equal to the net power supplied to the composite system. Taking into account also the choice of a sign made in Sections 6.1 and 6.2, we see that this agrees with the sign convention used in thermodynamics, stating that power supplied to any system, be it a subsystem or the composite system, is counted with a positive sign.
Junctions correspond to energy domains. Hence, an outer port exposes an energy domain of the respective composite system. Since this interpretation also applies to the primitive systems discussed in Section 6, we can say that ports also represent energy domains and whenever two ports are connected to the same junction the energy domains are identified. In other words, interconnection means that systems share common energy domains. The state of a shared energy domain is given by the shared state variable, according to the equality of state. The connected systems may exchange the respective physical quantity according to the equality of net flow. Moreover, such an exchange also implies an exchange of energy, according to the equality of effort.
Example 4.1.
Let’s again consider the pattern in Figure 7. At the junction representing the kinetic energy domain, the momentum variable (state) is shared and forces (flows) balance at equal velocity (effort):
[…]
Table 1 summarizes the physical interpretation of junctions representing a number of different energy domains.
energy domain state variable flows balance at equal efforts
kinetic momentum forces balance at equal velocity
magnetic flux linkage voltages balance at equal current
potential displacement velocities balance at equal force
hydraulic volume volume rates balance at equal pressure
electric charge currents balance at equal voltage
thermal entropy entropy rates balance at equal temperature
Table 1: Physical interpretation of junctions as energy domains.
We can now briefly reconsider the surjectivity and injectivity conditions which distinguish interconnection patterns form undirected wiring diagrams. The surjectivity condition says that each junction must have at least one connected inner port. In other words, an energy domain of the composite system must belong to at least one of its subsystems. Of course, an energy domain that is not part of the system cannot be exposed. The injectivity condition says that each junction may have at most one connected outer port. In other words, an energy domain can be exposed at most once. Exposing an energy domain twice and identifying it again at a higher level of nesting by connecting the two respective ports would lead to ill-defined flow variables.
4.2 Interconnection patterns as relations
[…]
Similarly to this prototypical example, the relation assigned to an interconnection pattern is given by a subbundle of the Cartesian product of the two bundles of port variables associated to the inner and the outer interface.
Definition 4.2.
Let $P=(\beta,\,I,\,J)$ be an interconnection pattern and let $I_{\text{in}}=\sum\beta$ denote its inner interface. Further, let $\mathcal{P}_{I_{\text{in}}}\twoheadrightarrow\mathcal{X}_{I_{\text{in}}}$ be the bundle of port variables associated to the inner interface and let $\mathcal{P}_{I}\twoheadrightarrow\mathcal{X}_{I}$ be the bundle of port variables associated to the outer interface. Based on the partition $J$, let $\mathcal{P}_{P}\subset\mathcal{P}_{I_{\text{in}}}\times\mathcal{P}_{I}$ be the subspace where equality of state, equality of effort, and equality of net flow are satisfied. Similarly, let $\mathcal{X}_{P}\subset\mathcal{X}_{I_{i}}\times\mathcal{X}_{I}$ be the subspace where equality of state is satisfied. Then, the semantics assigned to interconnection pattern $P$ is the relation $\mathcal{P}_{P}:\mathcal{P}_{I_{\text{in}}}\nrightarrow\mathcal{P}_{I}$ between the port variables of the inner interface and the port variables of the outer interface that is given by the vector subbundle $\mathcal{P}_{P}\twoheadrightarrow\mathcal{X}_{P}$. ∎
It can be shown that this assignment of semantics is functorial, which means that it makes no difference whether one first composes some interconnection patterns and then asks for the relation assigned to the composite or one first asks for the relations assigned to the original patterns and then composes these relations.
Figure 8: Interconnection pattern of a DC shunt motor model. The stator and rotor are connected to a common electric energy domain that is exposed via the outer port $\mathtt{q}$. The kinetic energy domain of the rotor is exposed via the outer port $\mathtt{p}$. The conversion between electric and kinetic energy in the rotor depends on the magnetic energy domain of the stator, which is represented by the junction with connected state ports named $\mathtt{b_{k}}$. The full model is discussed in Section 9.
Example 4.3.
Consider the interconnection pattern $P=(\beta,\,I,\,J)$ depicted in Figure 8 with names of inner boxes given by $\operatorname{dom}(\beta)=\{\mathtt{stator},\,\mathtt{rotor}\}$ and interfaces given by
$\beta(\mathtt{stator})\>=\>\left(\{\mathtt{q},\,\mathtt{b_{s}}\},\,\tau_{\text{s}}\right)$
$\beta(\mathtt{rotor})\>=\>\left(\{\mathtt{q},\,\mathtt{p},\,\mathtt{b_{s}}\},\,\tau_{\text{r}}\right)$
$I\>=\>\left(\{\mathtt{q},\,\mathtt{p}\},\,\tau\right)\,,$
where
$\tau_{\text{s}}(\mathtt{q})\>=\>\tau_{\text{r}}(\mathtt{q})\>=\>\tau(\mathtt{q})\>=\>((\mathbb{R},\,\mathtt{charge}),\,\mathsf{p})$
$\tau_{\text{s}}(\mathtt{b_{s}})\>=\>\tau_{\text{r}}(\mathtt{b_{s}})\>=\>((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{s})$
$\tau_{\text{r}}(\mathtt{p})\>=\>\tau(\mathtt{p})\>=\>((\mathbb{R},\,\mathtt{angular\_momentum}),\,\mathsf{p})\,.$
Furthermore, the interconnection is determined by
$J\>=\>\bigl{\{}\{\mathtt{stator.q},\,\mathtt{rotor.q},\,\mathtt{q}\},\,\{\mathtt{stator.b_{s}},\,\mathtt{rotor.b_{s}}\},\,\{\mathtt{rotor.p},\,\mathtt{p}\}\bigr{\}}\,.$
The port space of the inner interface $I_{\text{in}}=\sum\beta$ is
$\begin{split}\mathcal{P}_{I_{\text{in}}}\>=\>&\mathbb{T}\mathbb{R}\times\mathbb{R}\times\mathbb{T}\mathbb{R}\times\mathbb{T}\mathbb{R}\times\mathbb{R}\\ \>\ni\>&\bigl{(}(\mathtt{stator.q.x},\,\mathtt{stator.q.f},\,\mathtt{stator.q.e}),\,\mathtt{stator.b_{s}.x},\,\\ &\>\>(\mathtt{rotor.q.x},\,\mathtt{rotor.q.f},\,\mathtt{rotor.q.e}),\,\\ &\>\>(\mathtt{rotor.p.x},\,\mathtt{rotor.p.f},\,\mathtt{rotor.p.e}),\,\mathtt{rotor.b_{s}.x}\bigr{)}\end{split}$
and the port space of the outer interface $I$ is
$\mathcal{P}_{I}\>=\>\mathbb{T}\mathbb{R}\times\mathbb{T}\mathbb{R}\>\ni\>\bigl{(}(\mathtt{q.x},\,\mathtt{q.f},\,\mathtt{q.e}),\,(\mathtt{p.x},\,\mathtt{p.f},\,\mathtt{p.e})\bigr{)}\,.$
Finally, the subspace $\mathcal{P}_{P}\subset\mathcal{P}_{I_{\text{in}}}\times\mathcal{P}_{I}$ defining the relation is
$\begin{split}\mathcal{P}_{P}\>=\>\bigl{\{}\bigl{(}&(\mathtt{stator.q.x},\,\mathtt{stator.q.f},\,\mathtt{stator.q.e})\,,\,\mathtt{stator.b_{s}.x}\,,\,\\ &(\mathtt{rotor.q.x},\,\mathtt{rotor.q.f},\,\mathtt{rotor.q.e})\,,\,\\ &(\mathtt{rotor.p.x},\,\mathtt{rotor.p.f},\,\mathtt{rotor.p.e})\,,\,\mathtt{rotor.b_{s}.x}\,,\,\\ &(\mathtt{q.x},\,\mathtt{q.f},\,\mathtt{q.e})\,,\,(\mathtt{p.x},\,\mathtt{p.f},\,\mathtt{p.e})\bigr{)}\in\mathcal{P}_{I_{i}}\times\mathcal{P}_{I}\,\mid\,\\ &\mathtt{stator.q.x}=\mathtt{rotor.q.x}=\mathtt{q.x}\,,\,\\ &\mathtt{stator.q.e}=\mathtt{rotor.q.e}=\mathtt{q.e}\,,\,\\ &\mathtt{stator.q.f}+\mathtt{rotor.q.f}=\mathtt{q.f}\,,\,\\ &\mathtt{stator.b_{s}.x}=\mathtt{rotor.b_{s}.x}\,,\,\\ &\mathtt{rotor.p.x}=\mathtt{p.x}\,,\,\\ &\mathtt{rotor.p.e}=\mathtt{p.e}\,,\,\\ &\mathtt{rotor.p.f}=\mathtt{p.f}\,\bigr{\}}\,.\qed\end{split}$
5 Exergy
Every system is ultimately determined by a power-preserving interconnection of primitive subsystems, also called components, which represent elementary physical behaviors such as storage and reversible as well as irreversible exchange of energy. To achieve a thermodynamically consistent combination of reversible and irreversible dynamics, components are defined with respect to a fixed exergy reference environment that is discussed in this section.
5.1 Intuition
As briefly discussed in Section 1.4, thermodynamic models do not have all their physically-relevant degrees of freedom determined by their state variables, which leads to irreversibility. According to the first and second law of thermodynamics, internal energy, which is given by an imprecise phenomenological description, cannot be fully converted into mechanical or electromagnetic energy, which is precise, in the sense that the relevant degrees of freedom are determined by the state variables.
The irreversible degradation of ‘mechanically precise’ forms of energy into ‘thermodynamically imprecise’ forms of energy can be quantified using the concept of exergy. A hypothetical conversion device, whose mode of operation is only restricted by the first and second law of thermodynamics is used to determine how much mechanical (or electromagnetic) energy can be extracted from a system before it reaches thermodynamic equilibrium with a reference environment, whose intensive variables, such as temperature and pressure, are constant (in time and space), see Figure 9.
Figure 9: To assess the exergy content of a system, it is combined with a reference environment. The exergy content is then defined as the maximum amount of work that can be extracted from the combined system via a hypothetical conversion device that operates completely reversibly.
Crucially, the ideal conversion device is not restricted to interacting with the system through its interface. Instead, the device operates in a counterfactual where all energy domains of both the system and the environment can be accessed. This means that the exergy content of a (hierarchically flattened) composite system can be obtained by summing the exergy stored in each of its storage components. So, in particular, the exergy content of a system does not depend on how its subsystems are interconnected. The exergy content of a component storing internal energy is assessed based on an ideal conversion device and the reference environment, while the exergy content of a component storing only mechanical or electromagnetic forms of energy is equal to its energy content.
As also remark that a similar concept of available energy could be defined intrinsically, meaning without relying on a reference environment. The reversible conversion device would then interact only with the system of interest until it reaches equilibrium with itself. However, the use of a reference environment with fixed intensive quantities has the important advantage that the exergy content of a system is given explicitly by its exergy storage function, which follows systematically from its energy storage function and the definition of the reference environment, as detailed in Section 6.1. In contrast, the intrinsic concept in general would lead to an optimization problem without a simple closed-form solution.
More information and references about exergy, its history, and its links to port-Hamiltonian as well as GENERIC systems can be found in Lohmayer et al. (2021).
5.2 Reference environment
Definition 5.1.
A reference environment $(I_{R},\,\lambda_{R})$ is given by an interface $I_{R}=(N_{R},\,\tau_{R})$ such that each port $p\in I_{R}$ is a power port with an unique extensive quantity in $\mathbf{Q}_{\mathbb{R}}$ and an associated value $\lambda_{R}(p)\in\mathbb{R}^{*}$ of the energy-conjugate intensive quantity. ∎
Uniqueness of quantities, i.e. injectivity of $\tau_{R}:N_{R}\rightarrow\mathbf{Q}_{\mathbb{R}}\times\{\mathsf{p}\}$, is required to determine the exergy storage function of a system from its energy storage function. Regarding the physical interpretation, each port corresponds to an energy domain that is present in the environment. Allowing multiple energy domains of the same kind would conflict with the environment always being in equilibrium with itself.
Example 5.2.
The minimal definition of a reference environment represents an isothermal heat bath. This amounts to an interface with a single entropy port and an associated reference temperature, henceforth denoted by ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}$:
$N_{R}\>$ =$=\>\{\mathtt{s}\}$
$\tau_{R}(\mathtt{s})\>$ $=\>\left((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p}\right)$
[…]
To determine the exergy content of pressurized fluids with respect to an isobaric atmosphere, the definition of the environment can be extended to $N_{R}=\{\mathtt{s},\,\mathtt{v}\}$ with
$\tau_{R}(\mathtt{v})\>$ $=\>\left((\mathbb{R},\,\mathtt{volume}),\,\mathsf{p}\right)$
[…]
The minus sign in front of the reference pressure ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\pi_{0}}$ stems from the fact that more volume means less energy. To model for instance a mass diffusion process, a port $\mathtt{m}$ representing the mass of a chemical species with a fixed chemical potential ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\mu_{0}}$ should be included, i.e. let $N_{R}=\{\mathtt{s},\,\mathtt{v},\,\mathtt{m}\}$ with
[…]
A software implementation should include a comprehensive definition of a reference environment and allow the user to redefine the values of the intensive quantities.
6 Primitive systems
The lowest level of any hierarchical model specification is defined by primitive systems, also called components. There are four classes of components:
  • 1.
    Storage components represent energy storage.
  • 2.
    Reversible components represent reversible energy exchange
    (or a transformation between different forms of describing an energy domain).
  • 3.
    Irreversible components represent irreversible energy exchange.
  • 4.
    Environment components are used to model energy exchange with
    the reference environment.
Components are in general defined with respect to a reference environment $(I_{R},\,\lambda_{R})$. Further, every component has an interface $I=(N,\,\tau)$ with port variables $(x,\,f,\,e)\in\mathcal{P}_{I}$. To define the exergy storage function of storage components and to state fundamental thermodynamic properties of reversible as well as irreversible components, we define the following notation: For any port $p\in I$, let
$\lambda(p)\>=\>\begin{cases}\lambda_{R}(p_{R})&\exists\,p_{R}\in I_{R}\,\colon\,\tau_{R}(p_{R})\>=\>\tau(p)\\ 0\in\mathrm{T}^{*}\mathcal{X}_{p}&\text{else.}\end{cases}$
For instance, if $p$ represents a thermal energy domain then $\lambda(p)={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}$, whereas if $p$ represents a kinetic energy domain then $\lambda(p)=0$, since the environment does not have a kinetic energy domain. This can also be applied to the entire interface. E.g. for $N=\{\mathtt{p_{1}},\,\ldots,\,\mathtt{p_{n}}\}$, let
$\lambda(I)\>=\>\bigl{(}\lambda(\mathtt{p_{1}}),\,\ldots,\,\lambda(\mathtt{p_{n}})\bigr{)}\,.$
We also define a similar notation that references a specific environment port: For any port $p\in I$ and $p_{R}\in I_{R}$, let
$\lambda(p,\,p_{R})\>=\>\begin{cases}\lambda_{R}(p_{R})&\tau_{R}(p_{R})\>=\>\tau(p)\\ 0\in\mathrm{T}^{*}\mathcal{X}_{p}&\text{else.}\end{cases}$
Again, given e.g. $N=\{\mathtt{p_{1}},\,\ldots,\,\mathtt{p_{n}}\}$, let
$\lambda(I,\,p_{R})\>=\>\bigl{(}\lambda(\mathtt{p_{1}},\,p_{R}),\,\ldots,\,\lambda(\mathtt{p_{n}},\,p_{R})\bigr{)}\,.$
6.1 Storage components
Storage components are primitive systems representing energy storage. Examples include elastic springs, movable masses, capacitors, inductors, and thermal capacities.
Definition 6.1.
A storage component $(I,\,E)$ is defined by an interface $I$ containing only power ports and an energy storage function $E:\mathcal{X}_{I}\rightarrow\mathbb{R}$. ∎
The function $E$ simply yields the stored energy for any state $x\in\mathcal{X}_{I}$. Based on the reference environment $I_{R}$, the exergy content is determined as follows.
Definition 6.2.
Let $(I,\,E)$ be a storage component. Its exergy storage function $H:\mathcal{X}_{I}\rightarrow\mathbb{R}$ is given by
$H(x)\>=\>E(x)-\sum_{p\,\in\,I}\sum_{\begin{subarray}{c}p_{R}\,\in\,I_{R}\\ \tau_{R}(p_{R})\,=\,\tau(p)\end{subarray}}\langle\lambda_{R}(p_{R})\mid p\mathtt{.x}\rangle\,.\qed$
[…]
$\forall\,p\in I\>:\>p\mathtt{.e}\>=\>\left.\partialderivative{H}{(p\mathtt{.x})}\right|_{x}$
In case the constant offset should matter for thermodynamic analysis, it can be added to the total exergy function once the equilibrium state of the system under study has been determined. Finally, the trajectory of the state variables is obtained by integrating the flow variables:
[…]
Example 6.3.
Consider a storage component $(I_{\text{gas}},\,E_{\text{gas}})$ representing a gas-filled compartment whose entropy and volume can change. Its interface $I_{\text{gas}}$ is defined as in 2.18 and we assume knowledge of its energy storage function $E_{\text{gas}}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$. Writing $x=(s,\,v)$, the exergy storage function is then given by
$H_{\text{gas}}(s,\,v)\>=\>E_{\text{gas}}(s,\,v)\,-\,{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}\,s\,+\,{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\pi_{0}}\,v\,.$
With the temperature being defined as $\theta=\partialderivative{E_{\text{gas}}}{s}$ and the negative pressure being defined as $-\pi=\partialderivative{E_{\text{gas}}}{v}$, the effort variables are given by
$\begin{split}\mathtt{s.e}\>&=\>\theta-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}\\ \mathtt{v.e}\>&=\>-(\pi-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\pi_{0}})\,.\end{split}$
Since the entropy rate $\dot{s}=\mathtt{s.f}$ is related to the heat flow rate $\Phi$ via $\dot{s}=\frac{1}{\theta}\,\Phi$, the exergetic power at the entropy port is $\langle\mathtt{s.e}\mid\mathtt{s.f}\rangle=\frac{\theta-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}}{\theta}\,\Phi=\eta_{\text{Carnot}}\,\Phi$, where $\eta_{\text{Carnot}}$ is the efficiency of an ideal heat engine operating between a hot source with temperature $\theta$ and a cold sink with temperature ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}$. Considering a piston with the pressurized gas on one side and the atmosphere on the other, $\langle\mathtt{v.e}\mid\mathtt{v.f}\rangle=-(\pi-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\pi_{0}})\,\dot{v}$ is the power that can be extracted from the pressurized gas when its volume grows at rate $\dot{v}$. ∎
Next, we define the semantics of a storage component with interface $I$ in terms of a relation $\mathcal{S}:\mathrm{T}\mathcal{X}_{I}\nrightarrow\mathcal{P}_{I}$. This is used in Section 7, where we define the semantics of a composite system in terms of the semantics of its primitive subsystems and the semantics of its (hierarchy of nested) interconnection pattern(s).
Definition 6.4.
The semantics assigned to storage component $(I,\,E)$ is the relation $\mathcal{S}:\mathrm{T}\mathcal{X}_{I}\nrightarrow\mathcal{P}_{I}$ that is given by the subbundle
[…]
where $H$ is induced from $E$ according to 6.2. ∎
6.2 Environment components
While the primary role of the reference environment is to enable the thermodynamically consistent combination of reversible and irreversible dynamics, it may also be represented by a component to model energy exchange with the environment. For instance, an environment component can represent an isothermal heat bath, which may absorb heat generated by irreversible components, or an isobaric atmosphere, which may exchange pressure-volume work with other components.
Definition 6.5.
An environment component is defined by a subinterface $I\subseteq I_{R}$ containing the exposed energy domains of the reference environment. ∎
The state variables keep track of the exchanged entropy, volume, etc. As for storage components, their rate of change is given by the flow variables:
[…]
Since the exergy content of the environment is zero by definition, the effort variables are also zero:
[…]
Consequently, the semantics of an environment component is assigned as follows:
Definition 6.6.
The semantics assigned to environment component $I\subseteq I_{R}$ is the relation $\mathcal{E}:\mathrm{T}\mathcal{X}_{I}\nrightarrow\mathcal{P}_{I}$ that is given by the subbundle
[…]
6.3 Reversible components
In a generalized sense, reversible components represent gyrators, transformers, and/or kinematic constraints. Generalized gyrators describe the reversible coupling of different kinds of energy domains, such as a coupling of kinetic and potential energy domains in mechanics or an electro-magnetic coupling. Examples where one domain reversibly interacts with itself are gyroscopic forces in rigid body dynamics and self-advection of kinetic energy in fluid dynamics. The Lorentz force is another example where a kinetic energy domain interacts with itself in a way that depends on the state of a magnetic energy domain. Generalized transformers couple energy domains of the same kind. They may for instance represent mechanical levers or ideal electrical transformers. Frame transformations in multibody systems present an example where a transformation between different representations of the same energy domain is realized. Examples of kinematic constraints include joint constraints in multibody dynamics and the incompressibility constraint in fluid dynamics. Further, the parallel connection of capacitors, the series connection of inductors or springs, as well as rigid links between point masses are described by constraints.
The geometric structure behind reversible components is called Dirac structure. For our purposes, an (almost) Dirac structure $\mathcal{D}$ is a power-preserving, locally linear relation of port variables. More precisely, we could define $\mathcal{D}\subset\mathcal{P}_{I}$ as a maximally isotropic vector subbundle of the bundle of port variables $\mathcal{P}_{I}\twoheadrightarrow\mathcal{X}_{I}$, see Courant (1990); Bursztyn (2013); van der Schaft and Jeltsema (2014). However, to state fundamental properties of reversible components, we give a more practical definition of Dirac structures in terms of a particular representation, which may be called constrained hybrid input-output representation (cf. van der Schaft and Jeltsema (2014)):
Definition 6.7.
Let $I$ be an interface and let $\mathcal{X}_{I,p}\cong\mathcal{X}_{1}\times\mathcal{X}_{2}\ni x_{p}=(x_{1},\,x_{2})$ be a factorization of the state space of its power ports with flow and effort variables correspondingly denoted by $f=(f_{1},\,f_{2})$ and $e=(e_{1},\,e_{2})$. A Dirac structure $\mathcal{D}\subset\mathcal{P}_{I}$ admits the following representation:
$\mathcal{D}\,=\,\left\{(x,\,f,\,e)\in\mathcal{P}_{I}\,\mid\,\exists\,\lambda_{c}\in V^{*}\,\colon\left[\begin{array}[]{c}f_{1}\\ \hline\cr e_{2}\\ \hline\cr 0\end{array}\right]\,=\,\left[\begin{array}[]{c | c | c}L(x)&-g(x)&C^{*}(x)\\ \hline\cr g^{*}(x)&0&0\\ \hline\cr-C(x)&0&0\end{array}\right]\,\left[\begin{array}[]{c}e_{1}\\ \hline\cr f_{2}\\ \hline\cr\lambda_{c}\end{array}\right]\right\}$
Here, $L:\mathrm{T}^{*}\mathcal{X}_{1}\rightarrow\mathrm{T}\mathcal{X}_{1}$, $g:\mathrm{T}\mathcal{X}_{2}\rightarrow\mathrm{T}\mathcal{X}_{1}$, its linear dual $g^{*}:\mathrm{T}^{*}\mathcal{X}_{1}\rightarrow\mathrm{T}^{*}\mathcal{X}_{2}$, $C:\mathrm{T}^{*}\mathcal{X}_{1}\rightarrow\mathcal{X}_{I}\times V$, and its dual $C^{*}:\mathcal{X}_{I}\times V^{*}\rightarrow\mathrm{T}\mathcal{X}_{1}$ are vector bundle maps over the identity on $\mathcal{X}_{I}$. Further, $V$ is a vector space and $L(x)$ is skew-symmetric for all $x\in\mathcal{X}_{I}$. ∎
The maps $L$, $g$ and $C$ define a gyrator, a transformer and a constraint, respectively. The constraint is enforced via the Lagrange multiplier $\lambda_{c}$. Due to the skew-symmetry of the matrix in 6.7, exergetic power is conserved, i.e.
[…]
Example 6.8.
Let the interface $I_{\text{hk}}=(N_{\text{hk}},\,\tau_{\text{hk}})$ be given by
$\begin{split}N_{\text{hk}}\>&=\>\{\mathtt{v_{1}},\,\mathtt{v_{2}},\,\mathtt{p}\}\\ \tau_{\text{hk}}(\mathtt{v_{1}})\>&=\>\tau_{\text{hk}}(\mathtt{v_{2}})\>=\>((\mathbb{R},\,\mathtt{volume}),\,\mathsf{p})\\ \tau_{\text{hk}}(\mathtt{p})\>&=\>((\mathbb{R},\,\mathtt{momentum}),\,\mathsf{p})\end{split}$
and let $a\in\mathbb{R}$ be a constant surface area of a piston. Then, the following equation defines a Dirac structure $\mathcal{D}_{\text{hk}}\subset\mathcal{P}_{I_{\text{hk}}}$:
$\left[\begin{array}[]{c}\mathtt{v_{1}.f}\\ \mathtt{v_{2}.f}\\ \mathtt{p.f}\end{array}\right]\>=\>\left[\begin{array}[]{rrr}0&0&a\\ 0&0&-a\\ -a&a&0\end{array}\right]\,\left[\begin{array}[]{c}\mathtt{v_{1}.e}\\ \mathtt{v_{2}.e}\\ \mathtt{p.e}\end{array}\right]$
This describes a generalized gyrator that couples the two hydraulic energy domains at the front and the back side of a piston with a kinetic energy domain corresponding to the piston’s moving mass. ∎
A reversible component is defined by a Dirac structure which satisfies two extra properties ensuring thermodynamic consistency.
Definition 6.9.
A reversible component $(I,\,\mathcal{D})$ is given by an interface $I$ and a Dirac structure $\mathcal{D}\subset\mathcal{P}_{I}$ such that all quantities of the reference environment are conserved and the time-reversal invariance property holds. The first condition requires that
$\forall\,p_{R}\in I_{R}\>:\>\forall\,x\in\mathcal{X}_{I}\>:\>\Bigl{(}0\>=\>L(x)\,\lambda(I_{1},\,p_{R})\quad\text{and}\quad 0\>=\>g^{*}(x)\,\lambda(I_{1},\,p_{R})\Bigr{)}\,.$
The time-reversal invariance property holds if $\forall\,x\in\mathcal{X}_{I}\,\colon$
${(L(x))}^{ij}\neq 0$ $\implies$-⟹ - $\mathrm{P}(x_{1}^{i})\>$ $=\>\mathrm{P}({(L(x))}^{ij})\,\mathrm{P}(x_{1}^{j})$
${(g(x))}^{ik}\neq 0$$\implies$ $\mathrm{P}(x_{1}^{i})\>$ $=\>\mathrm{P}({(g(x))}^{ik})\,\mathrm{P}(x_{2}^{k})\,.$
Here, we use abstract index notation without implied summation. Indices $i$ and $j$ range over the power ports associated with $\mathcal{X}_{1}$ and $k$ ranges over the power ports associated with $\mathcal{X}_{2}$. The state variable is accordingly split as $x=(x_{p},\,x_{s})=((x_{1},\,x_{2}),\,x_{s})$ and $\mathrm{P}$ yields the parity of the respective quantity with $\mathrm{P}(\mathrm{const.})=+1$. ∎
Example 6.10.
The reversible component $(I_{\text{hk}},\,\mathcal{D}_{\text{hk}})$, as already defined in 6.8, describes the coupling between the kinetic energy domain of a piston with surface area $a$ and the two hydraulic energy domains on either side of the piston. The first condition here ensures that volume is conserved. This is satisfied since
$\left[\begin{array}[]{c}0\\ 0\\ 0\end{array}\right]\>=\>\left[\begin{array}[]{rrr}0&0&a\\ 0&0&-a\\ -a&a&0\end{array}\right]\,\left[\begin{array}[]{c}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}-\pi_{0}}\\ {\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}-\pi_{0}}\\ 0\end{array}\right]\,.$
The time-reversal invariance property holds since
$\begin{split}-\mathrm{P}(\mathtt{(\mathbb{R},\,volume}))\>&=\>\mathrm{P}(a)\,\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\\ -\mathrm{P}((\mathbb{R},\,\mathtt{volume}))\>&=\>\mathrm{P}(-a)\,\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\,.\end{split}$
Checks for the other two non-zero entries of $L$ must pass due to its skew-symmetry. ∎
The semantics of a reversible component is simply given by its Dirac structure interpreted as a relation:
Definition 6.11.
The semantics assigned to reversible component $(I,\,\mathcal{D})$ is the relation $\mathcal{D}:1\nrightarrow\mathcal{P}_{I}$ that is given by its Dirac structure $\mathcal{D}$. ∎
The symbol $1$ in $1\nrightarrow\mathcal{P}_{I}$ refers to the Cartesian unit for vector bundles, i.e. the vector bundle with a single point as its base space and the zero-dimensional vector space above it.
6.4 Irreversible components
Irreversible components are phenomenological models of irreversible processes with local thermodynamic equilibrium. Examples include mechanical friction, fluid viscosity, electrical conduction / resistance, thermal conduction, and mass diffusion. Such models can be defined in terms of a function that maps thermodynamic forces, which drive the relaxation process, to thermodynamic fluxes, which seek to restore equilibrium. Based on the assumption that the microscopic dynamics are reversible, Onsager found that such maps must satisfy a symmetry property Onsager (1931). This motivates our definition of an Onsager structure:
Definition 6.12.
Let $I$ be an interface. An Onsager structure $\mathcal{O}\subset\mathcal{P}_{I}$ is a fiber subbundle that admits the following representation:
$\mathcal{O}\,=\,\left\{(x,\,f,\,e)\in\mathcal{P}_{I}\,\mid\,f\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}\,M(e)\,e\right\}\,,$
where $M$ is a smooth function that yields a symmetric non-negative definite linear operator $M(e)\colon\mathrm{T}^{*}\mathcal{X}_{I,p}\to\mathrm{T}\mathcal{X}_{I,p}$ for all effort variables $e\in\mathrm{T}^{*}\mathcal{X}_{I,p}$. ∎
In contrast to a Dirac structure, an Onsager structure is neither power-preserving, nor is it a locally linear relation of the port variables, since destruction of exergy, or equivalently production of entropy, is described in general by a non-linear relation. Hence, $\mathcal{O}\subset\mathcal{P}_{I}$ is not a vector subbundle of $\mathcal{P}_{I}$, but merely a fiber subbundle. The symmetry of $M(e)$ corresponds to Onsager’s reciprocal relations and its non-negative definiteness implies a non-negative exergy destruction rate.
Example 6.13.
Let the interface $I_{\text{mf}}=(N_{\text{mf}},\,\tau_{\text{mf}})$ be given by
$\begin{split}N_{\text{mf}}\>&=\>\{\mathtt{p},\,\mathtt{s}\}\\ \tau_{\text{mf}}(\mathtt{p})\>&=\>((\mathbb{R},\,\mathtt{momentum}),\,\mathsf{p})\\ \tau_{\text{mf}}(\mathtt{s})\>&=\>((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p})\end{split}$
and let $d\in\mathbb{R}_{>0}$ be a friction coefficient. Then, the following defines an Onsager structure $\mathcal{O}_{\text{mf}}\subset\mathcal{P}_{I_{\text{mf}}}$ representing mechanical friction:
$\begin{bmatrix}\mathtt{p.f}\\ \mathtt{s.f}\end{bmatrix}\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}\,d\,\begin{bmatrix}\theta&-\upsilon\\ -\upsilon&\frac{\upsilon^{2}}{\theta}\end{bmatrix}\,\begin{bmatrix}\mathtt{p.e}\\ \mathtt{s.e}\end{bmatrix}$
Here, $\upsilon=\mathtt{p.e}$ is the velocity and $\theta={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}+\mathtt{s.e}$ is the absolute temperature at which kinetic energy is dissipated into the thermal energy domain. The phenomenological parameter $d$ could also be a function of the effort variables $e$, i.e. it could depend on the velocity and the temperature, or theoretically also the base point $x$. The above can be simplified to $\mathtt{p.f}=d\,\upsilon$ and $\mathtt{s.f}=-\frac{1}{\theta}\,d\,\upsilon^{2}$. Hence, the environment temperature ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}$ cancels out in the flow variables, which here are the friction force and the entropy production rate (with a minus sign since thermal energy is leaving the component). ∎
An irreversible component is defined by an Onsager structure which satisfies two extra properties ensuring thermodynamic consistency:
Definition 6.14.
An irreversible component $(I,\,\mathcal{O})$ is given by an interface $I$ and an Onsager structure $\mathcal{O}\subset\mathcal{P}_{I}$ such that energy and all quantities of the reference environment, except for entropy, are conserved and the time-reversal invariance property does strictly not hold. The condition for conservation of energy requires that
[…]
The condition that all quantities of the reference environment, except for entropy, are conserved requires that
[…]
The time-reversal invariance property does strictly not hold if $\forall\,x\in\mathcal{X}_{I}\,\colon$
${(M(x))}^{ij}\neq 0\implies\mathrm{P}(x^{i})\>=\>\mathrm{P}({(M(x))}^{ij})\,\mathrm{P}(x^{j})\,.$
Here, we use the same notation as in 6.9 with indices $i$ and $j$ ranging over all power ports of $I$. ∎
Example 6.15.
The irreversible component $(I_{\text{mf}},\,\mathcal{O}_{\text{mf}})$, as already defined in 6.13, models mechanical friction. The condition for conservation of energy is satisfied since
$\begin{bmatrix}0\\ 0\end{bmatrix}\>=\>d\,\begin{bmatrix}\theta&-\upsilon\\ -\upsilon&\frac{\upsilon^{2}}{\theta}\end{bmatrix}\,\begin{bmatrix}\upsilon\\ \theta\end{bmatrix}\,.$
The condition that all quantities of the reference environment, except for entropy, are conserved is trivially satisfied. The time-reversal invariance property does strictly not hold since
$\begin{split}\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\>&=\>\mathrm{P}(d)\,\mathrm{P}(\theta)\,\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\\ \mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\>&=\>\mathrm{P}(d)\,\mathrm{P}(-\upsilon)\,\mathrm{P}((\mathbb{R},\,\mathtt{entropy}))\\ \mathrm{P}((\mathbb{R},\,\mathtt{entropy}))\>&=\>\mathrm{P}(d)\,\mathrm{P}(-\upsilon)\,\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\\ \mathrm{P}((\mathbb{R},\,\mathtt{entropy}))\>&=\>\mathrm{P}(d)\,{(\mathrm{P}(\upsilon))}^{2}\,/\,\mathrm{P}(\theta)\,\mathrm{P}((\mathbb{R},\,\mathtt{entropy}))\,,\end{split}$
where we used that temperature has positive parity and velocity has negative parity. ∎
Example 6.16.
Let the interface $I_{\text{ht}}=(N_{\text{ht}},\,\tau_{\text{ht}})$ be given by
$\begin{split}N_{\text{ht}}\>&=\>\{\mathtt{s_{1}},\,\mathtt{s_{2}}\}\\ \tau_{\text{ht}}(\mathtt{s_{1}})\>&=\>\tau_{\text{ht}}(\mathtt{s_{2}})\>=\>((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p})\,,\end{split}$
let $\alpha\in\mathbb{R}_{>0}$ be a heat transfer coefficient, and let the Onsager structure $\mathcal{O}_{\text{ht}}\subset\mathcal{P}_{I_{\text{ht}}}$ be given by
$\begin{bmatrix}\mathtt{s_{1}.f}\\ \mathtt{s_{2}.f}\end{bmatrix}\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}\,\alpha\,\begin{bmatrix}\frac{\theta_{2}}{\theta_{1}}&-1\\ -1&\frac{\theta_{1}}{\theta_{2}}\end{bmatrix}\,\begin{bmatrix}\mathtt{s_{1}.e}\\ \mathtt{s_{2}.e}\end{bmatrix}\,,$
where $\theta_{1}={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}+\mathtt{s_{1}.e}$ and $\theta_{2}={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}+\mathtt{s_{2}.e}$ are the absolute temperatures of the two thermal domains. Then, we can check that $(I_{\text{mf}},\,\mathcal{O}_{\text{mf}})$, defines an irreversible component. Again, the phenomenological parameter $\alpha$ could for instance be a function of the temperatures. The above can be simplified to $\mathtt{s_{1}.f}=-\frac{1}{\theta_{1}}\,\alpha\,(\theta_{2}-\theta_{1})$ and $\mathtt{s_{2}.f}=-\frac{1}{\theta_{2}}\,\alpha\,(\theta_{1}-\theta_{2})$. Hence, the environment temperature ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}$ cancels out in the flow variables, which are rates of entropy change due to heat transfer. ∎
The semantics of an irreversible component is simply given by its Onsager structure interpreted as a relation:
Definition 6.17.
The semantics assigned to irreversible component $(I,\,\mathcal{O})$ is the relation $\mathcal{O}:1\nrightarrow\mathcal{P}_{I}$ that is given by its Onsager structure $\mathcal{O}$. ∎
7 Composite systems
A composite system is formed by interconnecting finitely many systems, be they primitive systems or other composite systems, according to an interconnection pattern. The data defining a composite system is therefore given by an interconnection pattern and for each of its inner boxes a system with the corresponding interface.
Definition 7.1.
Let $\mathbf{System}$ refer to the set of all systems defined with respect to a fixed reference environment. A composite system $(I,\,\gamma,\,J)$ is defined by the following data:
  1. 1.
    An interface $I$ of the composite system.
  2. 2.
    A package of systems $\gamma\in\mathbf{Pkg}(\mathbf{System})$, which can be reduced to a package of interfaces $\beta\in\mathbf{Pkg}(\mathbf{Interface})$ by forgetting further data.
  3. 3.
    A partition $J$ of the combined interface $P=\sum\left[\mathtt{inner}\mapsto\sum\beta,\,\mathtt{outer}\mapsto I\right]$ such that $(\beta,\,I,\,J)$ defines an interconnection pattern. ∎
class of system interface relation definition
[…]
Table 2: Overview of systems and their semantics given by a relation.
Table 2 recalls the different classes of components defined in Section 6. In order to write the relations that determine their semantics in a consistent form, the Cartesian unit is inserted a few times. As already shown in the last row of the table, we can state the semantics of a composite system in the same form. The idea is that we just need to compose the relation assigned to the interconnection pattern with the Cartesian product of the relations assigned to the subsystems filling its inner boxes.
Definition 7.2.
Let $(I,\,\gamma,\,J)$ be a composite system and let $P=(\beta,\,I,\,J)$ be its underlying interconnection pattern. Further, let $\mathcal{P}_{P}$ be the relation assigned to pattern $P$ according to 4.2. For any inner box $b\in\operatorname{dom}(\gamma)$, let $\left\llbracket{s}\right\rrbracket$ denote the relation assigned to the subsystem $s=\gamma(b)$. In case system $s$ is itself a composite system, this definition applies recursively. Then, the semantics assigned to composite system $(I,\,\gamma,\,J)$ is the relation $\mathcal{C}:\mathrm{T}\mathcal{X}_{I_{s}}\times\mathrm{T}\mathcal{X}_{I_{e}}\nrightarrow\mathcal{P}_{I}$ given by
$\mathcal{C}\>=\>\mathcal{P}_{P}\circ\left(\prod_{b\,\in\,\operatorname{dom}(\gamma)}\left\llbracket{\gamma(b)}\right\rrbracket\right)\,.$
Here, $\mathcal{X}_{I_{s}}$ denotes the state space of all storage components and $\mathcal{X}_{I_{e}}$ denotes the state space of all environment components, which are subsystems of the composite system in a recursive sense. ∎
Example 7.3.
Consider the composite system given by filling the inner boxes of the interconnection pattern for the nonisothermal damper shown in Figure 6 with three suitable systems. The system filling the box $\mathtt{tc}$ represents a thermal capacity and it is defined analogous to 6.3, but without a port for volume exchange. It gives a relation
$\mathcal{S}_{\text{tc}}:\mathrm{T}\mathcal{X}_{\text{tc}}\times 1\nrightarrow\mathbb{T}\mathcal{X}_{\text{tc}}$
with $\mathcal{X}_{\text{tc}}=\mathbb{R}$. The system filling the box $\mathtt{mf}$ is a mechanical friction model as defined in 6.15 and it gives a relation
$\mathcal{O}_{\text{mf}}:1\times 1\nrightarrow\mathbb{T}\mathcal{X}_{\text{mf}}$
with $\mathcal{X}_{\text{mf}}=\mathbb{R}\times\mathbb{R}$. The system filling the box $\mathtt{ht}$ is a heat transfer model as defined in 6.16 and it gives a relation
$\mathcal{O}_{\text{ht}}:1\times 1\nrightarrow\mathbb{T}\mathcal{X}_{\text{ht}}$
with $\mathcal{X}_{\text{ht}}=\mathbb{R}\times\mathbb{R}$. Let the Cartesian product of the three relations be denoted by
$\mathcal{S}_{\text{tc}}\times\mathcal{O}_{\text{mf}}\times\mathcal{O}_{\text{ht}}\colon\mathrm{T}\mathcal{X}_{\text{tc}}\times 1\nrightarrow\mathbb{T}\mathcal{X}_{\text{tc}}\times\mathbb{T}\mathcal{X}_{\text{mf}}\times\mathbb{T}\mathcal{X}_{\text{ht}}\,.$
Further, let the semantics of the interconnection pattern be given by the relation
$\mathcal{P}_{P}\colon\mathbb{T}\mathcal{X}_{\text{tc}}\times\mathbb{T}\mathcal{X}_{\text{mf}}\times\mathbb{T}\mathcal{X}_{\text{ht}}\nrightarrow\mathcal{P}_{I}\,,$
where $\mathcal{P}_{I}=\mathbb{T}\mathcal{X}_{\mathtt{p}}\times\mathbb{T}\mathcal{X}_{\mathtt{s}}=\mathbb{T}\mathbb{R}\times\mathbb{T}\mathbb{R}$ is the bundle of port variables of the outer interface. Then, the semantics of the composite system is given by the relation
$\mathcal{C}\>=\>\mathcal{P}_{P}\circ\left(\mathcal{S}_{\text{tc}}\times\mathcal{O}_{\text{mf}}\times\mathcal{O}_{\text{ht}}\right)$
from $\mathrm{T}\mathcal{X}_{\text{tc}}\times 1$ to $\mathcal{P}_{I}$. ∎
For an isolated composite system the assigned relation $\mathcal{C}\colon\mathrm{T}\mathcal{X}_{I_{s}}\times\mathrm{T}\mathcal{X}_{I_{e}}\nrightarrow 1$ simply amounts to a subbundle of $\mathrm{T}\mathcal{X}_{I_{s}}\times\mathrm{T}\mathcal{X}_{I_{e}}$, which should implicitly define a vector field on $\mathcal{X}_{I_{s}}\times\mathcal{X}_{I_{e}}$. Given an initial condition, one can integrate this vector field to determine the evolution of the system. We simply say that $(x,\,\dot{x})\in\mathcal{C}$, where $x=(x_{s},\,x_{e})$, gives the dynamics of the isolated system. This is generalized by the following definition:
Definition 7.4.
Let $s\in\mathbf{System}$ be a system with interface $I$ and let the relation $\mathcal{R}\colon\mathrm{T}\mathcal{X}_{I_{s}}\times\mathrm{T}\mathcal{X}_{I_{e}}\nrightarrow\mathcal{P}_{I}$ define its semantics, according to one of the definitions mentioned in Table 2. Further, let $(x_{s},\,x_{e})\in\mathcal{X}_{I_{s}}\times\mathcal{X}_{I_{e}}$ denote the state variables associated to all (nested) storage and environment components. If $s$ is a primitive system this may be vacuous. Moreover, let $(x,f,e)\in\mathcal{P}_{I}$ denote the port variables of its (outer) interface. Again, if the system is isolated this is vacuous. Then, the dynamics of system $s$ is determined by the implicit differential-algebraic equation
$\bigl{(}(x_{s},\,\dot{x}_{s}),\,(x_{e},\,\dot{x}_{e}),\,(x,\,f,\,e)\bigr{)}\in\mathcal{R}\,.\qed$
Example 7.5.
The dynamics of the composite system described in 7.3 is given by
$\bigl{(}(s,\dot{s}),\,(\mathtt{p.x},\,\mathtt{p.f},\,\mathtt{p.e}),\,(\mathtt{s.x},\,\mathtt{s.f},\,\mathtt{s.e})\bigr{)}\in\mathcal{C}\,.$
This boils down to the following equations:
$\dot{s}\>$ $=\>\frac{1}{\theta_{1}}\left(d\,\upsilon^{2}\,+\,\alpha\,(\theta_{2}-\theta_{1})\right)$
$\mathtt{p.f}\>$ $=\>d\,\upsilon$
$\mathtt{s.f}\>$ $=\>-\frac{1}{\theta_{2}}\,\alpha\,(\theta_{1}-\theta_{2})$
[…]
For programming languages it usually holds that not everything that is syntactically correct is also semantically meaningful. There are also cases where a compiler for the EPHS language should throw an error, despite of syntactic correctness. Every junction must have either zero or one connected port belonging to a storage or environment component. Also, one cannot connect more than two power ports with a quantity such as $(\mathbb{R},\,\mathtt{displacement})$ to the same junction, since velocities don’t balance. As of now, we don’t have a comprehensive set of restrictions that should be imposed on the interconnection of systems in order to get something meaningful, but we hope that this can be worked out in the future.
8 Thermodynamic consistency
As already touched upon in the introduction, the GENERIC establishes a general structure, also called metriplectic structure, for combining reversible and irreversible dynamics in a way that respects fundamental principles from thermodynamics, namely the first law, the second law, and Onsager reciprocal relations.
The vector field which determines the dynamics of a metriplectic system is given as the sum of a reversible Hamiltonian vector field that conserves energy due to the skew-symmetry of the Poisson bivector and a vector field of an irreversible gradient flow that has non-negative entropy production due to the non-negative definiteness of the metric tensor. Further, so-called non-interacting or degeneracy conditions ensure that the reversible dynamics conserves entropy and the irreversible dynamics conserves energy. The first law then holds since both parts conserve energy and the second law holds because the irreversible part has non-negative entropy production, while the reversible part conserves entropy, see e.g. Morrison (1984); Grmela and Öttinger (1997). Onsager reciprocal relations hold, as the symmetry of the metric tensor implies equal cross-effect coefficients, see e.g. Mielke et al. (2016).
Usually, the reversible part and the irreversible part both satisfy further degeneracy conditions that ensure the conservation of further quantities such as volume or mass. Given that (some of) these quantities, together with entropy, are present in an exergy reference environment that is adjoined to the metriplectic system, it follows that the reversible dynamics conserve exergy and the irreversible dynamics are characterized by a non-negative exergy destruction. This immediately follows from the definition of the exergy storage function analogous to 6.2.
The argument for EPHS is very similar in nature, although it is in a sense reversed. To keep it short and simple, we focus on isolated systems.
Proposition 8.1.
The dynamics of any isolated EPHS respects the first law, the second law, and Onsager reciprocal relations. ∎
Proof.
We argue in terms of components and their interconnection.
  1. 1.
    According to 6.9, all reversible components conserve exergy as well as all quantities of the reference environment.
    Considering the definition of exergy in 6.2, it follows immediately that all reversible components also conserve energy.
  2. 2.
    According to 6.14, all irreversible components are defined by a non-negative exergy destruction rate and they conserve energy as well as all quantities of the reference environment, except for entropy.
    Considering the definition of exergy, it follows that all irreversible components also have a non-negative entropy production rate.
  3. 3.
    According to 4.2, any interconnection of the components conserves exergy and all quantities (of the reference environment), since flows balance at equal efforts.
    Considering the definition of exergy, it follows that any interconnection of the components also conserves energy.
The first law holds since all three aspects conserve energy. The second law holds because the irreversible components have a non-negative entropy production rate, while the reversible components and the interconnection conserve entropy. The reciprocal relations hold because, according to 6.12, thermodynamic fluxes (flows) depend on thermodynamic forces (efforts) via a symmetric locally-linear operator. ∎
We note that the argument is reversed because, according to Definitions 6.4 and 7.2, the semantics of EPHS is defined in terms of a storage function that is interpreted as the exergy of the system, as opposed to two separate functions representing the energy and entropy of a system. The concept of a storage function originates in Willems’ theory of dissipative systems Willems (1972). In this context, reversible components as well as all EPHS without any irreversible components are said to be lossless, while (all systems with) irreversible components are said to be dissipative.
9 Example
This section compares a bond graph and an EPHS model of a DC shunt motor. A direct current (DC) motor has a fixed outer part, called stator, and a rotating inner part, called rotor. Both parts contain a coil of wire through which current flows. In the case of a shunt motor, these two inductors are connected in parallel (aka in shunt) to a source of electric energy. The working principle of such motors is based on a coupling of the electro-magnetic and the kinetic energy domains via the Lorentz force and the commutation of the rotor coil after every half turn. However, both aspects are not directly taken into account by the two models, which instead consider the coupling as being lumped into a gyrator element.
Figure 10: Bond graph model of a DC shunt motor.
We start with the bond graph model that is taken from Borutzky (2010) and redrawn in Figure 10. The half arrows are called bonds and each bond has two associated variables, called flow and effort. The name of the flow variable is written on the side of the arrow head, while the name of the effort variable is written on the opposite side. However, not all variables are explicitly named in the example. At any 00-junction the effort variables of all connected bonds are equal, while the flow variables balance, with their sign given by the bond direction. Dually, at $1$-junctions flows are equal and efforts balance. Specifically, at the 00-junction on the left the current from the $\mathrm{Se}$-element (effort source) with effort/voltage $E$ splits into the current $i_{s}$ through the stator coil and the current $i_{r}$ through the rotor coil. The inductance of the rotor coil is represented by the $\mathrm{I}$-element (generalized inductor) with inductance $L_{r}$, while an $\mathrm{I}$-element representing the stator coil inductance is omitted. The (steady-state) stator coil current $i_{s}$ is instead entirely determined by the $\mathrm{R}$-element (generalized resistor) at the top representing the coil’s electrical resistance $R_{s}$. The magnetic flux of the stator coil $b_{s}$ is instead obtained as the signal that results from multiplying the signal $i_{s}$, which is the shared effort variable at the upper $1$-junction, with the constant gain $L_{s}$. The bond graph hence combines energy-based modeling with signal-based modeling known from block diagrams. At the $1$-junction on the left, the source voltage $E$ is equal to the sum of the voltage over the rotor coil inductance $L_{r}$, the voltage $u_{R}$ over the rotor coil resistance $R_{r}$, and the induced (back EMF) voltage $u_{r}$. The latter is determined by the $\mathrm{MGY}$-element (modulated gyrator), which describes the coupling of the electrical and the mechanical parts. The gyrator ratio is the signal $b_{s}$ coming from the gain block. At the gyrator, the effort variable of one port is equal to the product of the gyrator ratio and the flow variable of the other port. On the mechanical side, the load torque $M_{\text{load}}\leq 0$ is modeled by the effort source on the right, while the rotor’s angular mass $J_{r}$ is represented by an $\mathrm{I}$-element. At the $1$-junction the sum of the generated torque $M$ and the load torque $M_{\text{load}}$ is equal to the sum of the torque accelerating the angular mass and the friction torque $M_{R}$. The latter is the product of the friction coefficient $d_{r}$ and the angular velocity $\omega$. Putting all this together, the bond graph gives the system of ordinary differential equations
[…]
where
$i_{s}\>$ $=\>\frac{1}{R_{s}}\,E$
$b_{s}\>$ $=\>L_{s}\,i_{s}$
$u_{r}\>$ $=\>b_{s}\,\omega$
$M\>$ $=\>b_{s}\,i_{r}$
$M_{R}\>$ $=\>d_{r}\,\omega$
$u_{R}\>$ $=\>R_{r}\,i_{r}\,.$
As already shown in 4.3, the EPHS model is specified as an interconnection of two subsystems named $\mathtt{stator}$ and $\mathtt{rotor}$. The two systems share an electric energy domain represented by the junction with connected ports named $\mathtt{q}$. The kinetic energy domain of the rotor is represented by the junction with connected ports named $\mathtt{p}$. Both energy domains are exposed such that the motor model can be integrated into another system. The conversion between electric and kinetic energy in the rotor depends on the state of the magnetic energy domain of the stator, which is represented by the junction with connected state ports named $\mathtt{b_{s}}$.
Figure 11: Interconnection pattern of the stator model.
The interconnection pattern of the stator model is shown in Figure 11. The box $\mathtt{em}$ is filled by a reversible component that represents the coupling of the electric energy domain with port $\mathtt{q}$ and the magnetic energy domain with port $\mathtt{b_{s}}$. The box $\mathtt{coil}$ is filled by a storage component that represents the inductance of the stator coil, while the box $\mathtt{emloss}$ is filled by an irreversible component that represents its electric resistance. The box $\mathtt{env}$ is filled by an environment component whose thermal energy domain with port $\mathtt{s}$ absorbs the dissipated electromagnetic energy.
The reversible component $(I_{\text{em}},\,\mathcal{D}_{\text{em}})$ is defined by its interface $I_{\text{em}}=(\{\mathtt{q},\,\mathtt{b_{s}}\},\,\tau_{\text{em}})$ with $\tau_{\text{em}}(\mathtt{q})=((\mathbb{R},\,\mathtt{charge}),\,\mathsf{p})$, $\tau_{\text{em}}(\mathtt{b_{s}})=((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{p})$ and its Dirac structure $\mathcal{D}_{\text{em}}\subset\mathcal{P}_{I_{\text{em}}}$ given by
$\begin{bmatrix}\mathtt{q.f}\\ \mathtt{b_{s}.f}\end{bmatrix}\>=\>\begin{bmatrix}0&1\\ -1&0\end{bmatrix}\,\begin{bmatrix}\mathtt{q.e}\\ \mathtt{b_{s}.e}\end{bmatrix}\,.$
The storage component $(I_{\text{coil}},\,E_{\text{coil}})$ is defined by its interface $I_{\text{coil}}=(\{\mathtt{b_{s}}\},\,\tau_{\text{coil}})$ with $\tau_{\text{coil}}(\mathtt{b_{s}})=((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{p})$ and its energy function $E_{\text{coil}}:\mathcal{X}_{I_{\text{coil}}}\rightarrow\mathbb{R}$ with $\mathcal{X}_{I_{\mathtt{coil}}}=\mathbb{R}\ni\mathtt{b_{s}.x}=b_{s}$ given by
$E_{\mathtt{coil}}(b_{s})\>=\>\frac{1}{2\,L_{s}}\,b_{s}^{2}\,,$
where the parameter $L_{s}$ is the inductance. Since $H_{\mathtt{coil}}=E_{\mathtt{coil}}$, the effort variable $\mathtt{b_{s}.e}=\frac{b_{s}}{L_{s}}$ is the current through the coil.
The irreversible component $(I_{\text{emloss}},\,\mathcal{O}_{\text{emloss}})$ is defined by $I_{\text{emloss}}=(\{\mathtt{b_{s}},\,\mathtt{s}\},\,\tau_{\text{emloss}})$ with $\tau_{\text{emloss}}(\mathtt{b_{s}})=((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{p})$, $\tau_{\text{emloss}}(\mathtt{s})=((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p})$ and its Onsager structure $\mathcal{O}_{\text{emloss}}\subset\mathcal{P}_{\text{emloss}}$ given by
$\begin{bmatrix}\mathtt{b_{s}.f}\\ \mathtt{s.f}\end{bmatrix}\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}\,R_{s}\,\begin{bmatrix}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}+}\mathtt{s.e}&-\mathtt{b_{s}.e}\\ -\mathtt{b_{s}.e}&\frac{{(\mathtt{b_{s}.e})}^{2}}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}+}\mathtt{s.e}}\end{bmatrix}\,\begin{bmatrix}\mathtt{b_{s}.e}\\ \mathtt{s.e}\end{bmatrix}\>=\>\begin{bmatrix}R_{s}\,\mathtt{b_{s}.e}\\ -\frac{R_{s}\,{(\mathtt{b_{s}.e})}^{2}}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}+}\mathtt{s.e}}\end{bmatrix}\,,$
where the parameter $R_{s}$ is the electric resistance of the coil.
The environment component is defined by the subinterface $I_{\text{env}}=\{\mathtt{s}\}\subseteq I_{R}$.
Figure 12: Interconnection pattern of the rotor model.
The interconnection pattern of the rotor model is shown in Figure 12. The components filling the boxes $\mathtt{coil}$, $\mathtt{em}$, $\mathtt{emloss}$, and $\mathtt{env}$ are defined as for the stator model, except for a different port name, namely $\mathtt{b}$ instead of $\mathtt{b_{s}}$, and different parameters, namely $L_{r}$ instead of $L_{s}$ and $R_{r}$ instead of $R_{s}$. The box $\mathtt{mk}$ is filled by a reversible component that describes the coupling of the magnetic energy domain with port $\mathtt{b}$ and the (rotational) kinetic energy domain with port $\mathtt{p}$. This coupling depends on the magnetic flux $\mathtt{b_{s}.x}$ that is generated by the stator coil. The box $\mathtt{mass}$ is filled by a storage component that represents the rotor’s angular mass and the box $\mathtt{mloss}$ is filled by an irreversible component that represents its mechanical friction.
The reversible component $(I_{\text{mk}},\mathcal{D}_{\text{mk}})$ is defined by $I_{\text{mk}}=(\{\mathtt{b},\,\mathtt{p},\,\mathtt{b_{s}}\},\,\tau_{\text{mk}})$ with $\tau_{\text{mk}}(\mathtt{b})=((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{p})$, $\tau_{\text{mk}}(\mathtt{p})=((\mathbb{R},\,\mathtt{angular\_momentum}),\,\mathsf{p})$, $\tau_{\text{mk}}(\mathtt{b_{s}})=((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{s})$ and its Dirac structure $\mathcal{D}_{\mathtt{mk}}\subset\mathcal{P}_{I_{\text{mk}}}$ given by
$\begin{bmatrix}\mathtt{b.f}\\ \mathtt{p.f}\end{bmatrix}\>=\>\begin{bmatrix}0&\mathtt{b_{s}.x}\\ -\mathtt{b_{s}.x}&0\end{bmatrix}\,\begin{bmatrix}\mathtt{b.e}\\ \mathtt{p.e}\end{bmatrix}\,.$
The storage component $(I_{\text{mass}},\,E_{\text{mass}})$ is defined by $I_{\text{mass}}=(\{\mathtt{p}\},\,\tau_{\text{mass}})$ with $\tau_{\text{mass}}(\mathtt{p})=((\mathbb{R},\,\mathtt{angular\_momentum}),\,\mathsf{p})$ and $E_{\text{mass}}:\mathcal{X}_{I_{\text{mass}}}\rightarrow\mathbb{R}$ with $\mathcal{X}_{I_{\text{mass}}}=\mathbb{R}\ni\mathtt{p.x}=p$ given by
$E_{\text{mass}}(p)\>=\>\frac{1}{2\,J_{r}}\,p^{2}\,,$
where the parameter $J_{r}$ is the angular mass.
The irreversible component $(I_{\text{mloss}},\mathcal{O}_{\text{mloss}})$ is defined by $I_{\text{mloss}}=(\{\mathtt{p},\,\mathtt{s}\},\,\tau_{\text{mloss}})$ with $\tau_{\text{mloss}}(\mathtt{p})=((\mathbb{R},\,\mathtt{angular\_momentum}),\,\mathsf{p})$, $\tau_{\text{mloss}}(\mathtt{s})=((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p})$ and its Onsager structure $\mathcal{O}_{\text{mloss}}\subset\mathcal{P}_{I_{\text{mloss}}}$ given by
$\begin{bmatrix}\mathtt{p.f}\\ \mathtt{s.f}\end{bmatrix}\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}\,d_{r}\,\begin{bmatrix}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}+}\mathtt{s.e}&-\mathtt{p.e}\\ -\mathtt{p.e}&\frac{{(\mathtt{p.e})}^{2}}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}+}\mathtt{s.e}}\end{bmatrix}\,\begin{bmatrix}\mathtt{p.e}\\ \mathtt{s.e}\end{bmatrix}\>=\>\begin{bmatrix}d_{r}\,\mathtt{p.e}\\ -\frac{d_{r}\,{(\mathtt{p.e})}^{2}}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}+}\mathtt{s.e}}\end{bmatrix}\,,$
where the parameter $d_{r}$ is the mechanical friction coefficient.
We write $(b_{s},\,b,\,p,\,s)\in\mathbb{R}^{4}$ for the state of the interconnected motor model, where we identify the entropy state variables $s$ of both environment components. In future work, the identification of environment components can be formalized by composing with a suitable relation. The EPHS model then yields the equations
$\dot{b}_{s}\>$ $=\>\mathtt{q.e}-R_{s}\,i_{s}$
$\dot{b}\>$ $=\>\mathtt{q.e}-b_{s}\,\omega-R_{r}\,i_{r}$
$\dot{p}\>$ $=\>b_{s}\,i_{r}-d_{r}\,\omega+\mathtt{p.f}$
$\dot{s}\>$ $=\>\frac{1}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}}\left(R_{s}\,i_{s}^{2}+R_{r}\,i_{r}^{2}+d_{r}\,\omega^{2}\right)$
$\mathtt{q.f}\>$ $=\>i_{s}+i_{r}\,,$
where
$i_{s}\>$ $=\>\frac{1}{L_{s}}\,b_{s}$
$i_{r}\>$ $=\>\frac{1}{L_{r}}\,b$
$\omega\>$ $=\>\frac{1}{J_{r}}\,p\,.$
10 Discussion
While it still seems reasonable to package a motor model as a single bond graph, integrating the model as part of a more complex system is not supported by the bond graph notation. Hence, bond graph users have to reproduce models as parts of more complex models. EPHS improves upon this, making it nearly trivial to encapsulate, reuse, and swap out submodels. Another main feature of EPHS is their built-in thermodynamic consistency. Restricting the space of models in this way not only provides important guarantees, it also makes models compatible with each other, as a close link between their mathematical structure and physical interpretation is established. This is reflected by the correspondence of junctions and energy domains, which in our opinion contributes substantially to the improved readability of the EPHS syntax. The reason why EPHS users can easily integrate existing models without first needing to adapt them to fit into the larger system is that EPHS simply compose by sharing energy domains. In this way, we hope to improve upon port-Hamiltonian systems, which lack an intuitive graphical syntax and require that users define the ports of a system in a somewhat arbitrary manner.
Author contribution statement
Markus Lohmayer: Conceptualization, Investigation, Writing – Original Draft, Writing – Review & Editing, Visualization; Owen Lynch: Investigation, Writing – Original Draft; Sigrid Leyendecker: Supervision
Acknowledgements
We thank David Spivak for help when working out the concept of namespaces.

### v2

@@
 Mathematical modeling of real-world physical systems requires the consistent combination of a multitude of physical laws and phenomenological models.
-This challenging task can be greatly simplified by hierarchically decomposing systems.
-Moreover, the use of diagrams for expressing such decompositions helps make the process more intuitive and facilitates communication, even with non-experts.
+This challenging task can be greatly simplified by hierarchically decomposing systems into ultimately simple components.
+Moreover, the use of diagrams for expressing the decomposition helps make the process more intuitive and facilitates communication, even with non-experts.
 As an important requirement, models have to respect fundamental physical laws such as the first and the second law of thermodynamics.
-While some existing modeling frameworks can make such guarantees based on structural properties of their models, they lack a formal graphical syntax.
+While some existing modeling frameworks make such guarantees based on structural properties of their models, they lack a formal graphical syntax.
 We present a compositional and thermodynamically consistent modeling language with a graphical syntax.
-As its semantics, port-Hamiltonian systems are endowed with further structural properties and a fixed physical interpretation such that thermodynamic consistency is ensured in a way that is closely related to the GENERIC framework.
+In terms of its semantics, we essentially endow port-Hamiltonian systems with additional structural properties and a fixed physical interpretation, ensuring thermodynamic consistency in a manner closely related to the metriplectic or GENERIC formalism.
 While port-Hamiltonian systems are inspired by graphical modeling with bond graphs, neither the link between the two, nor bond graphs themselves, can be easily formalized.
 In contrast, our syntax is based on a refinement of the well-studied operad of undirected wiring diagrams.
-The language effectively decouples the construction of complex models via the graphical syntax from physical concerns, which are dealt with only at the level of primitive subsystems that represent elementary physical behaviors.
-As a consequence, reuse of models and substitution of their parts becomes possible.
-Finally, by construction, systems interact by exchanging exergy, i.e.
-energy that is available for doing work, so the language is particularly well suited for thermodynamic analysis and optimization.
+By combining a compositional, graphical syntax with an energy-based, thermodynamic approach, the presented modeling language simplifies the understanding, reuse, and modification of complex physical models.
 keywords:
-applied category theory, bond graphs, classical electromagnetism, classical mechanics, compositionality, exergy, GENERIC, graphical syntax, local thermodynamic equilibrium, multiphysics, nonequilibrium thermodynamics, port-Hamiltonian systems
+bond graphs , compositionality , GENERIC , metriplectic structure , nonequilibrium thermodynamics , undirected wiring diagrams
 journal: arXiv
 1 Introduction
-In this introduction we want to answer the following questions:
+In this introduction we mainly want to answer the following questions:
   1. 1.
@@
     Why is this practically relevant?
-In addition, we briefly summarize key ideas of nonequilibrium thermodynamics and we discuss related work that inspired our developments.
+In addition, we briefly summarize key ideas of nonequilibrium thermodynamics, and we discuss related work that inspired our developments.
 The introduction finishes with an outline of the main part of the paper.
 1.1 What?
-EPHS provide a compositional and thermodynamically consistent language for expressing mathematical models of multiphysical systems at macroscopic scales. Such systems may combine
+Exergetic Port-Hamiltonian Systems (EPHS) provide a compositional and thermodynamically consistent language for expressing mathematical models of multiphysical systems at macroscopic scales.
+Such systems may combine
   1. 1.
@@
     (e.g. friction/viscosity, electrical resistance/conduction, thermal conduction).
-The EPHS language is termed compositional since it enables a completely modular and hierarchical specification of models based on its simple graphical syntax, which is discussed in more detail in the next subsection.
-If one flattens any hierarchical nesting of systems, one finds that every system is determined by a power-preserving interconnection of primitive subsystems.
-There are three main classes of primitive systems:
+The EPHS language is termed compositional because it enables a fully modular and hierarchical approach to model specification through its simple, graphical syntax, as elaborated in the next subsection.
+Flattening any hierarchical nesting reveals that every EPHS model is fundamentally defined by a power-preserving interconnection of primitive subsystems.
+There are three main classes of primitive EPHS:
   1. 1.
-    Systems that represent energy storage are characterized by a function which yields the stored energy for each possible state of the primitive system.
+    Systems that represent energy storage are characterized by a function which yields the stored energy for each state of the primitive system.
   2. 2.
-    Systems that represent reversible energy exchange are characterized by a so-called Dirac structure, as known e.g. from (constrained) mechanical systems Courant (1990).
+    Systems that represent reversible energy exchange are characterized by a so-called Dirac structure [1], as known e.g. from (constrained) mechanical systems.
   3. 3.
@@
 corresponds to (c).
-Structural properties of the primitive systems guarantee that all models respect the first and the second law of thermodynamics, Onsager’s reciprocal relations, as well as further conservation laws (e.g.
-for mass).
-Section 1.4 aims for a very brief introduction to nonequilibrium thermodynamics.
+Structural properties of the primitive systems and their interconnection guarantee that all models respect the first and the second law of thermodynamics, as well as Onsager’s reciprocal relations.
+Inspired by the contribution in [2], previous work [3] has explored EPHS as a systematic approach to thermodynamic modeling based on port-Hamiltonian systems (see, e.g., [4]).
+By interpreting the ‘Hamiltonian’ as an exergy storage function, a stricter version of port-Hamiltonian structure can be formulated, ensuring a thermodynamically consistent reversible-irreversible splitting of the dynamics, as known from the metriplectic or GENERIC formalism (see, e.g., [5, 6]).
+The exergy storage function represents the theoretically ‘useful’ amount of energy, i.e., the maximum amount of work that can be extracted from a thermodynamic system using an ideal device, such as a Carnot engine.
+This ideal device has access not only to all forms of energy stored in the system but also to a (reference) environment, primarily characterized by a heat bath with constant temperature.
+The present work formalizes EPHS as a compositional modeling language with a simple, graphical syntax inspired by bond graphs (see, e.g., [7, 8]).
+This approach enables the construction of complex models by composing them from simpler subsystems, which can themselves be constructed from more basic subsystems, and so on.
+We draw on and extend ideas form (applied) category theory (see, e.g., [9, 10]) to propose a simple mathematical framework suitable for computer implementation.
+In principle, the graphical syntax and the associated compositional framework could also be adopted without the parts that ensure thermodynamic consistency.
+While this would relax certain conditions imposed on the primitive systems, it would also weaken the correspondence between the remaining structure and its physical interpretation.
 1.2 How?
-Figure 1: Graphical representation of an interconnection pattern with different syntactic elements shown in a different color.
-The outer box shown in orange represents the interface of the resulting composite system.
-The two protruding lines correspond to its ports called $\mathtt{f}$ and $\mathtt{b_{k}}$.
+Figure 1: Graphical display of an interconnection pattern with different syntactic elements shown in different colors.
+The outer box shown in purple represents the interface of the resulting composite system.
+The three protruding lines represent its ports called $\mathtt{f.p}$, $\mathtt{f.m}$ and $\mathtt{b_{k}}$.
 The two inner boxes shown in red represent (the interface of) the subsystems called $\mathtt{ke}$ (storage of kinetic energy) and $\mathtt{sa}$ (self-advection of kinetic energy).
-Both have ports called $\mathtt{p}$ (momentum) and $\mathtt{m}$ (mass), while $\mathtt{sa}$ has another port called $\mathtt{b_{k}}$ (advection of kinetic energy across the boundary of the spatial domain).
-The port combinator shown in purple combines the ports $\mathtt{p}$ (fluid momentum) and $\mathtt{m}$ (fluid mass) into the multiport $\mathtt{f}$ (fluid).
+Both have ports called $\mathtt{p}$ (momentum) and $\mathtt{m}$ (mass).
+Box $\mathtt{sa}$ has another port called $\mathtt{b_{k}}$ (advection of kinetic energy across the boundary of the spatial domain).
+The common prefix $\mathtt{f}$ (fluid) of the outer ports $\mathtt{f.p}$ (fluid momentum) and $\mathtt{f.m}$ (fluid mass) combines them into a multiport.
 The black dots represent junctions where power is exchanged among the connected ports.
-The dashed line indicates that port $\mathtt{m}$ of the port combinator does not partake in power exchange, but merely shares information about the state (fluid mass).
-Expressions in the graphical syntax are called interconnection patterns because a pattern specifies how a composite system can be formed through a power-preserving interconnection of subsystems.
-Figure 1 shows an example from fluid dynamics.
+The dashed line indicates that only information about the state (fluid mass) is exchanged via the outer port $\mathtt{f.m}$, with no energy transfer occurring.
+Expressions in the graphical syntax are called interconnection patterns, since such a pattern specifies how a composite system is formed through a power-preserving interconnection of subsystems.
+Figure 1 shows an example from fluid dynamics, see [11].
 The syntax deals with systems only in terms of their interfaces, i.e.
@@
 The second attribute is of Boolean type and indicates whether power can be exchanged over the port or merely information about system state, as indicated graphically by a solid or a dashed line, respectively.
-An interconnection pattern should be thought of as a function which, for every round inner box, takes as input a system with matching interface and returns a system whose interface matches the outer box.
-For instance, the pattern shown in Figure 1 represents a function which takes two subsystems whose interfaces are determined by the ports of the boxes $\mathtt{ke}$ and $\mathtt{sa}$, and it returns a composite system whose interface is determined by the ports of the outer box.
-In the graphical representation, we often draw inner boxes filled with a color that indicates the nature of the subsystems that we consider as inputs.
-In Figure 1, $\mathtt{ke}$ has a blue filling as it represents energy storage, i.e.
-the fluid’s kinetic energy, while $\mathtt{sa}$ has a green filling as it represents reversible energy exchange, i.e.
-the advection of momentum and mass by the fluid flow.
-Written on top of the outer box, kin is an identifier akin to a function name and the text in parenthesis is just a short description.
-The port combinator and the corresponding multiport $\mathtt{f}$ are syntactic sugar that allows the composite system to be connected to other systems without drawing parallel connections.
+We think of an interconnection pattern as a function.
+For each inner box, it takes as input a system with matching interface and it outputs a system whose interface matches the outer box.
+For instance, the pattern shown in Figure 1 represents a function that takes two subsystems, whose interfaces are determined by the ports of the boxes $\mathtt{ke}$ and $\mathtt{sa}$, and it combines them into a composite system, whose interface is given by the ports of the outer box.
+We draw inner boxes filled with a color to indicate the nature of the subsystems that we consider as inputs.
+In Figure 1, box $\mathtt{ke}$ has a blue filling, as it represents energy storage (the fluid’s kinetic energy).
+Box $\mathtt{sa}$ has a green filling, as it represents reversible energy exchange (advection of momentum and mass).
+Written on top of the outer box, kin is an identifier akin to a function name and the text in parentheses is just a short description.
+The multiport $\mathtt{f}$ facilitates integrating the system into a more complex model without drawing parallel connections.
 (a) Kinetic energy system
@@
 (c) EMHD system
-Figure 2: Hierarchical specification of an electro-magneto-hydrodynamics model: The fluid model f shown in (b) comprises the kinetic energy system kin shown in (a) and four further models which are not shown, namely an internal energy system int, a thermal conduction model th, and volume / shear viscosity models vol / shr.
-The electro-magneto-hydrodynamics model emhd shown in (c) comprises the fluid model f and three further models which are not shown, namely an electro-magnetic system em, an electric conduction model el, and a model for the electro-kinetic coupling ek.
-Figure 2 demonstrates the hierarchical specification of a model for electro-magneto-hydrodynamics.
-For simplicity, a spatial domain with empty or isolated boundary is assumed such that only the in-domain interconnection of the different models needs to be dealt with.
-The yellow filling of the inner boxes in Figures 1(b) and 1(c) indicates that another composite system, which in turn is specified with its own interconnection pattern, is supplied as input.
-Recursively substituting into all yellow inner boxes the respective interconnection patterns would yield an equivalent model specification without hierarchical decomposition.
-The composition of interconnection patterns through substitution is uniquely defined as long as interfaces match.
-Similarly, one can add a level of nesting by refactoring a number of subsystems into a reusable composite subsystem.
-In the fluid model, one could for instance factor out the kinetic energy system filling the box $\mathtt{kin}$ and the internal energy subsystem filling the box $\mathtt{int}$ into a reusable ideal compressible fluid model.
+Figure 2: Hierarchical specification of an electro-magneto hydrodynamics (EMHD) model [11]: The fluid model f shown in (b) includes the kinetic energy system kin depicted in (a), as well as four additional models not detailed here: an internal energy system int, a thermal conduction model th, and volume and shear viscosity models vol and shr.
+The electro-magneto hydrodynamics model emhd shown in (c) includes the fluid model f depicted in (b), as well as three additional models not detailed here: an electromagnetic system em, an electric conduction model el, and a model for the electro-kinetic coupling ekc.
+Figure 2 shows the hierarchical specification of a plasma model.
+For simplicity, a spatial domain with isolated boundary is assumed such that only the in-domain interconnection of the different models needs to be dealt with.
+A red filling indicates that the respective input is a primitive system modeling an irreversible process.
+A yellow filling signifies that the input is a nested composite system, specified in terms of its own pattern.
+Recursively substituting into all yellow boxes the respective patterns would yield an equivalent model specification without the hierarchical decomposition.
+The composition of patterns through substitution is uniquely defined whenever interfaces match (possibly up to a given renaming of the ports).
+No additional data or handling of ‘edge cases’ is required.
+One can hence easily add levels of abstraction by refactoring subsystems into reusable parts.
+In the fluid model, one could for instance factor out the kinetic energy system filling box $\mathtt{kin}$ and the internal energy system filling box $\mathtt{int}$ into a reusable model of an ideal compressible fluid, see [11].
 1.3 Why?
 EPHS provide a formal and intuitive modeling language suited for humans and software tools alike.
-As shown above, increasingly complex models can be assembled from simpler and ultimately primitive parts.
+As shown above, increasingly complex models are assembled from simpler and ultimately primitive parts.
 This fosters the reusability of models and encourages the conceptual separation of different levels of detail.
-Beyond the modeling process, this provides value in interdisciplinary communication and education.
-Some modeling errors can be completely avoided due to typed interfaces (given by ports with fixed physical quantities) and the built-in thermodynamic consistency.
-Since interconnection patterns can be used as exergy flow diagrams, EPHS models are particularly useful for thermodynamic analysis and optimization.
-The compositional nature of EPHS could enable a scalable approach to model transformations, such as discretizations, and also to the integration of scientific machine learning techniques in first-principles-based physical modeling.
-1.4 Nonequilibrium thermodynamics in a tiny nutshell
+Beyond the modeling process, this can be beneficial when the language is used in interdisciplinary communication and education.
+Some modeling errors can be completely avoided due to the built-in thermodynamic consistency and the typed interfaces (given by ports with fixed physical quantities).
+Since interconnection patterns are essentially energy/exergy flow diagrams, EPHS models are particularly useful for thermodynamic analysis and optimization.
+The compositional nature of EPHS could enable a scalable approach to model transformations, such as discretizations.
+It could also enable a practical integration of first-principles-based thermodynamic modeling and scientific machine learning.
+1.4 Nonequilibrium thermodynamics in a nutshell
 Most physical systems are characterized by a huge number of degrees of freedom but keeping track of them is usually neither possible nor useful.
-Nonequilibrium thermodynamics aims to provide a framework for coarse-grained modeling, meaning that fast dynamics at small scales are not resolved and a quantity called entropy is introduced to account for the missing information, see e.g.
-Pavelka et al.
-(2018).
-Coarse-graining is possible because, despite of fluctuations at the microscopic scale, one can eventually observe non-fluctuating properties at a more macroscopic scale; at least this is the case for isolated systems.
-For instance, temperature is a property that does not appear at the microscopic scale.
+Nonequilibrium thermodynamics aims to provide a framework for coarse-grained modeling, meaning that fast dynamics at small scales are not resolved, and a quantity called entropy is introduced to account for the missing information, see, e.g., [6].
+Coarse-graining is possible because, despite fluctuations at the microscopic scale, one can eventually observe non-fluctuating properties at a more macroscopic scale; at least this is the case for isolated systems.
+For instance, temperature is a property that only appears at more macroscopic scales.
 Thermodynamic properties are classified as extensive when they are proportional to the size of the system (e.g.
 entropy, volume, mass) or as intensive when they do not depend on the size of the system (e.g.
-temperature, pressure, chemical potential), see e.g.
-Callen (1985)
-A system is said to be in local thermodynamic equilibrium if for every material point there exists a neighborhood large enough to observe thermodynamic properties, yet small enough for the properties to be non-varying inside the neighborhood, see e.g.
-Öttinger (2005).
-The size of these neighborhoods presents a macroscopic scale at which the dynamics of the thermodynamic properties can be observed and also modeled.
-The first law of thermodynamics just restates a well-known fact from mechanics, namely that energy is a conserved quantity.
-However, instead of considering directly the mechanical and electromagnetic energy associated to microscopic degrees of freedom, a thermodynamic model with local equilibrium resorts to a phenomenological description called internal energy, which is considered to depend on a number of extensive properties, including entropy, and possibly further mechanical or electromagnetic state variables which are observable on the considered macroscopic scale.
+temperature, pressure, chemical potential), see, e.g., [12]
+A system is said to be in local thermodynamic equilibrium if for every material point there exists a neighborhood large enough to observe thermodynamic properties, yet small enough for the properties to be essentially non-varying inside the neighborhood, see, e.g., [13].
+The size of these neighborhoods corresponds to a macroscopic scale at which the dynamics of the thermodynamic properties can be observed and also modeled.
+The first law of thermodynamics restates a fact known from mechanics, namely that energy is a conserved quantity.
+However, instead of considering directly the mechanical and electromagnetic energy associated to microscopic degrees of freedom, a thermodynamic model with local equilibrium resorts to a phenomenological description.
+This so-called internal energy is considered to depend on a number of extensive properties, including entropy, and possibly further mechanical or electromagnetic state variables which are observable on the considered macroscopic scale.
 These quantities then serve as the state variables of a coarse-grained description.
-Moreover, temperature is the derivative of the internal energy with respect to entropy.
-Similarly, pressure is given by the derivative with respect to volume and the chemical potential is the derivative with respect to mass, see e.g.
-Callen (1985).
-Classical mechanics and electromagnetism exhibit reversible dynamics, meaning that one could in principle arrive again at the initial condition by reversing the direction of time (although for many systems this would be impossible due to chaos).
-Incomplete knowledge about the microscopic state immediately leads to irreversible dynamics and an irreversible dynamics in turn leads to a growth of entropy, as information (e.g.
-about the initial state of the system) is lost in the process.
-This is reflected by the second law of thermodynamics which states that the entropy of an isolated system never decreases and that the system evolves towards a state of maximum entropy, see e.g.
-Öttinger (2005).
+Moreover, temperature is defined as the derivative of the internal energy with respect to entropy.
+Similarly, pressure is given by the derivative with respect to volume and the chemical potential is the derivative with respect to mass, see, e.g., [12].
+Models in classical mechanics and electromagnetism exhibit reversible dynamics, meaning that one could in principle arrive again at the initial condition by formally reversing the direction of time (although for many systems this would be practically impossible due to chaos).
+Incomplete knowledge about the microscopic state immediately leads to an irreversible dynamics and an irreversible dynamics in turn leads to a growth of entropy, as information (e.g.
+about the initial state of the system) is lost as the system evolves.
+This is reflected by the second law of thermodynamics which states that the entropy of an isolated system never decreases and that the system evolves towards a state of maximum entropy, see, e.g., [13].
 In other words, the uncertainty about the microscopic state can only grow, not decrease.
 For models with local equilibrium, the irreversible dynamics can be expressed in terms of so-called thermodynamic fluxes that determine instantaneous changes of coarse-grained state variables.
-The fluxes are expressed as a function of thermodynamic forces, which are given by differences in intensive quantities, see e.g.
-Öttinger (2005); de Groot and Mazur (1984).
+The fluxes are expressed as a function of so-called thermodynamic forces, which are given by local differences in intensive quantities, see, e.g., [13, 14].
 For instance, temperature differences cause heat flux, according to Fourier’s law, and differences in chemical potential cause a diffusive mass flux, according to Fick’s law.
 Further, one may observe cross effects such as a diffusive mass flux caused by temperature differences (Soret effect) and likewise a heat flux caused by differences in chemical potential (Dufour effect).
-Onsager’s reciprocal relations state that the relation between fluxes and forces must possess a certain symmetry Onsager (1931).
-For instance, the ratio of the mass flux caused by the Soret effect and the respective temperature difference is equal to the ratio of the heat flux caused by the Dufour effect and the respective difference in chemical potential.
+Onsager’s reciprocal relations state that the relation between fluxes and forces must possess a certain symmetry [15].
+Specifically, the ratio of the mass flux caused by the Soret effect and the respective temperature difference is equal to the ratio of the heat flux caused by the Dufour effect and the respective difference in chemical potential.
 1.5 Related work
 Figure 3 gives a high-level view of this subsection.
-Figure 3: Informal Venn diagram illustrating that EPHS integrates ideas from graphical and energy-based modeling of physical systems with bond graphs, the GENERIC framework for nonequilibrium thermodynamics, the port-Hamiltonian theory of open dissipative systems (PHS), and applied category theory (ACT) research on the formalization of graphical languages and compositional dynamical systems.
+Figure 3: Informal Venn diagram illustrating that EPHS integrates ideas from graphical and energy-based modeling of physical systems with bond graphs (see, e.g., [7, 8]), the metriplectic/GENERIC framework for nonequilibrium thermodynamics (see, e.g., [5, 13]), port-Hamiltonian theory for open, dissipative systems (PHS) (see, e.g., [4]), and applied category theory (ACT) research on the formalization of graphical languages and compositional dynamical systems (see, e.g., [9, 16, 17, 18, 19]).
+1.5.1 Bond graphs
+Bond graphs provide a graphical notation for representing multi-domain models, particularly those involving mechanical and electrical systems, see, e.g., [7, 8].
+The main idea is that elements interact via so-called bonds that represent links for energy exchange.
+Each bond is associated with two dynamic variables, called flow and effort, whose product gives the exchanged power.
+Bonds can be connected through two types of junctions: 00-junctions, where the flow variables of all connected ports balance while all effort variables are equal, and $1$-junctions, where the effort variables balance while all flow variables are equal.
+In both cases, the net power at a junction is zero.
+The elements are commonly generalized capacitors, inductors, transformers, resistors, etc.
+The junctions are then seen as generalized Kirchhoff’s circuit laws.
 1.5.2 Port-Hamiltonian systems
-Apparent from their name, EPHS have a lot in common with Port-Hamiltonian Systems (PHS).
-PHS were originally inspired by relating the dynamical systems defined by generalized bond graphs Breedveld (1982) (i.e.
-bond graphs using a single kind of energy storage elements together with gyrator elements) to Hamiltonian mechanics Maschke (1991).
-While many authors make use of explicit input-output port-Hamiltonian systems, PHS constitute an acausal systems theory, inspired also by the behavioral approach of Willems Willems (1974, 2007).
-The dynamics of networks comprising energy storing elements and reversible coupling elements between them (such as generalized transformers and gyrators) can be formalized, from a geometric viewpoint, as a Hamiltonian system whose Poisson structure reflects the network topology Maschke et al.
-(1992).
-The generalization to networks with algebraic constraints and externally controlled ports is achieved by using Dirac structures, which unify Poisson structures and presymplectic structures van der Schaft and Maschke (1995).
-By specifying a resistive relation on external port variables, the implicit Hamiltonian dynamics can be combined with a dissipative gradient dynamics, allowing for network elements that represent e.g.
-dampers and resistors.
-The key properties of PHS are firstly a power-balance equation which states that the sum of the stored power and the dissipated power is equal to the supplied power (passivity) and secondly the possibility to combine port-Hamiltonian systems by composition of their Dirac structures with an interconnecting Dirac structure, see for instance van der Schaft and Jeltsema (2014).
-While the dynamical systems defined by EPHS models are essentially PHS, they have further structural properties and a fixed physical interpretation.
-PHS in general have no uniform thermodynamic interpretation and they lack a formal graphical syntax to express their interconnection.
-In contrast to EPHS, PHS also lack a built-in mechanism for referring to subsystems and ports.
-1.5.2 GENERIC framework
-The general equation for reversible-irreversible coupling (GENERIC) provides a framework for combining reversible and irreversible dynamics in a way that respects fundamental thermodynamic principles, namely the first law, the second law and the reciprocal relations discovered by Onsager Grmela and Öttinger (1997).
-The reversible part is defined by an entropy-conserving Hamiltonian dynamics, while the irreversible part is given by an energy-conserving gradient flow, see e.g.
-Öttinger (2005).
-We note that the same holds for closely-related earlier developments in this direction, including metriplectic systems, see Kaufman (1984); Grmela (1984a, b); Morrison (1984).
-While a large class of well-known dynamical systems from physics can be identified as instances of the GENERIC, see e.g.
-Pavelka et al.
-(2018), the framework is especially helpful for the development of new thermodynamic models.
-The combination of Hamiltonian and gradient dynamics in both PHS and GENERIC systems led to the initial observation of a relationship between the two frameworks in Badlyan et al.
-(2018).
-GENERIC systems with local thermodynamic equilibrium can be seen as port-Hamiltonian systems, where the storage function corresponds to the system’s exergy.
-This function yields the maximum amount of work that can be extracted by an ideal (i.e.
-purely reversible) process which may arbitrarily interact with the system’s storage components and a given reference environment Lohmayer et al.
-(2021).
-The dynamical systems defined by (isolated) EPHS models are basically GENERIC systems, since their built-in thermodynamic consistency and associated physical interpretation follow from the same ideas.
-Beyond the aspect of compositionality, there are two main differences: 1.
-GENERIC systems are more general in the sense of including systems beyond local thermodynamics equilibrium (where the concept of temperature does not apply).
-2.
-The reversible part of EPHS is defined using Dirac structures, rather than less general Poisson structures (which cannot express kinematic constraints).
-This is crucial e.g.
-for modeling multibody systems and electrical circuits.
-1.5.3 Bond graph modeling
-Bond graph modeling is based on a graphical notation for representing dynamical systems and in particular multi-domain models involving mechanical and electrical elements Paynter (1961); Borutzky (2010).
-The main idea is that elements interact via bonds that represent links for energy exchange.
-Associated to each bond, there are two dynamic variables called flow and effort, whose product (or duality pairing) gives the exchanged power.
-Bonds can be connected via two types of junctions, namely 00-junctions, where the flow variables of all connected ports balance, with all effort variables being equal, and dually defined $1$-junctions, where the effort variables balance, with all flow variables being equal.
-In both cases, the net power at a junction is zero.
-While PHS can be seen as a geometric formalization of the dynamical systems defined by generalized bond graphs, the algebraic formalization of the bond graph notation itself is not straightforward, when compared to e.g.
-standard multigraphs, see Birkett and Roe (1989); Coya (2017).
-The inherent complexity of the notation is evident also in practice, since the direction in which exchange of power is counted positive and sometimes also the question which variable is flow and which is effort have to be decided arbitrarily, bond by bond.
-While bond graphs are frequently used in the literature to illustrate PHS, their use as a formal syntax for model specification is not established.
-Previous work has focused on algorithms for converting certain classes of bond graphs to port-Hamiltonian systems, see e.g.
-Pfeifer et al.
-(2020).
-In contrast to bond graph notation, the EPHS syntax only has 00-junctions and every interconnection pattern has an outer box.
-As demonstrated by various applications, a single type of junctions is sufficient and moreover the simplification entails crucial advantages: Interconnection patterns are easier to interpret, in particular due to the direct correspondence between junctions and energy domains.
-This correspondence arises from the inherent distinction between the roles of flows and efforts, a feature that is lacking in bond graphs.
-Further, the direction in which exchange of power is counted positive is predetermined in accordance with the convention used in thermodynamics, i.e.
-energy supplied to a system has a positive sign.
-Last but not least, interconnection patterns can be hierarchically composed and decomposed in a natural way.
+Port-Hamiltonian Systems (PHS) were originally inspired by relating the dynamical systems obtained from generalized bond graphs111 While bond graphs commonly use generalized capacitors and inductors as storage elements, so-called generalized bond graphs use only generalized capacitors together with gyrator elements [21].
+to Hamiltonian mechanics [20].
+While many authors make use of PHS in input-state-output form, PHS generally constitute an acausal systems theory, also inspired by Willems’ behavioral approach [22, 23].
+The dynamics of networks composed of energy storing elements and reversible coupling elements (generalized transformers and gyrators) can, from a geometric perspective, be formalized as Hamiltonian systems whose Poisson structure reflects the network topology [24].
+The generalization to systems with algebraic constraints and/or external port variables is achieved using Dirac structures [25].
+These offer a relational approach that unifies Poisson and presymplectic structures (see, e.g., [1, 26]).
+By modeling resistive elements as relations imposed on external port variables, Hamiltonian dynamics can be combined with (dissipative) gradient dynamics.
+A key feature of PHS is their composability: interconnecting simpler systems yields a more complex one.
+This is achieved by composing the Dirac structures of the individual systems with an ‘interconnecting Dirac structure’ (see, e.g., [4]).
+Another key feature is that every PHS satisfies a power-balance equation, stating that at every instant the sum of the stored energy and the dissipated energy equals the supplied energy.
+As the dissipated energy cannot be negative, PHS are inherently passive.
+In this context, the term ‘energy’ stems from the ‘Hamiltonian’ storage function.
+For EPHS, this function is interpreted as the stored exergy, making the power-balance equation a statement about rates of exergy transfer.
+This interpretation fully reconciles the inherent passivity of PHS with the first and second law of thermodynamics, as exergy is indeed dissipated (or destroyed) by irreversible dynamics.
+1.5.3 Metriplectic and GENERIC framework
+Metriplectic systems as well as the general equation for reversible-irreversible coupling (GENERIC) refer to a framework for combining reversible and irreversible dynamics in a thermodynamically consistent manner, see, e.g.
+[5, 27, 28, 29, 30].
+The reversible part is modeled as a Hamiltonian dynamics, which inherently conserves energy, while the irreversible part is modeled as a (generalized) gradient flow, which inherently drives the system toward maximum entropy.
+Two additional conditions ensure that the reversible dynamics conserves entropy, and the irreversible dynamics conserves energy.
+As a result, the combined dynamics respects both the first law and the second law (see, e.g., [13, 6]).
+While many well-known physical models can be identified as metriplectic systems or instances of the GENERIC, the framework is particularly useful as a guide for developing new thermodynamic models.
 1.5.4 Applied category theory
-The compositional nature of EPHS can be made precise using category theory.
-Category theory was first applied in algebraic geometry and topology and now is used to organize many branches of mathematics Mac Lane (1998).
-In recent decades, category theory has also grown beyond pure math to organize scientific and engineering disciplines as well, giving birth to ‘applied category theory’ Fong and Spivak (2019).
-A specific approach to categorical systems theory Myers (2022) that has inspired this work is based on the following principles, which can essentially also be found in the work of Willems Willems (2007).
+Category theory was first applied in algebraic geometry and topology and now is used to organize many branches of mathematics (see, e.g., [10]).
+In recent decades, category theory has grown beyond pure math to organize scientific and engineering disciplines, giving birth to applied category theory (see, e.g.
+[9]).
+The following ideas from categorical systems theory (see, e.g., [19]) have inspired our work.
   1. 1.
-    A shift of focus from individual systems to how systems relate to each other.
+    The shift of focus from individual systems to how systems relate to each other.
   2. 2.
@@
     The description of systems involves syntax and semantics.
-Syntax typically consists of combinatorial data that can be manipulated computationally, while semantics are given by geometric objects that exist only platonically.
-Additionally, the syntax can be displayed and manipulated diagrammatically as well as in text.
+Syntax consists of combinatorial data that can be manipulated computationally, while semantics are given by geometric objects that exist only platonically.
 An explicit understanding of the translation from syntax to semantics serves as a guide for computer implementation.
-For instance, there is a line of research which formalizes graphical languages for systems and process such as block diagrams Baez and Erberle (2015), Petri nets Baez and Pollard (2017), junctions of bond graphs Coya (2017), and passive linear circuits Baez and Fong (2018).
-These works consider a graphical syntax as some sort of string diagrams, established via a generators and relations approach.
-More closely related to our work, another line of research uses the theory of operads to define a graphical syntax which supports hierarchical (de-)composition, see Baez et al.
-(2020); Foley et al.
-(2021).
-Operads organize formal operations with finitely many inputs (subsystems) and one output (composite system), see Leinster (2004); Yau (2016).
-As proposed in Lohmayer and Leyendecker (2022), the EPHS syntax, is based on the operad of undirected wiring diagrams, which initially appeared in Spivak (2013), together with its relational semantics used to express database queries, see also Spivak (2014); Yau (2018); Breiner et al.
-(2020).
-In a similar lane, directed wiring diagrams are used to formalize nested systems of Moore machines and ODEs with outputs, see Vagner et al.
-(2015); Schultz et al.
-(2019).
-A combination of dynamical systems based on directed and undirected wiring diagrams is presented in Libkind (2020); Libkind et al.
-(2021).
-The composition of lossless port-Hamiltonian systems was recently studied from a categorical perspective in Lynch (2022).
+A prominent line of research formalizes graphical languages such as block diagrams [17], Petri nets [31], junctions of bond graphs [32], and passive linear circuits [33].
+In each case, the graphical syntax is defined as some sort of string diagrams, using a generators and relations approach.
+Another line of research uses the theory of operads to define a graphical syntax that supports hierarchical (de-)composition (see, e.g., [34, 35]).
+Operads organize formal operations with finitely many inputs (subsystems) and one output (composite system) (see, e.g., [36, 37]).
+As proposed in [38], the EPHS syntax, is based on the operad of undirected wiring diagrams, which initially appeared in [16], together with a relational semantics used to express database queries, see also [39, 40].
+Directed wiring diagrams have been used to formalize nested systems of Moore machines and ODEs with inputs and outputs [41, 42].
+A framework for discrete and continuous dynamical systems with directed and undirected notions of composition is presented in [43].
+The composition of lossless PHS is studied from a categorical perspective in [44].
 1.6 Outline
@@
 Based on this, Section 3 defines interconnection patterns.
-Section 4 shows how every interconnection pattern implies a relation between the port variables of its interfaces.
+Section 4 shows how an interconnection pattern implies a relation between the port variables of its interfaces.
 Section 5 discusses how exergy quantifies the energy stored in a system that is available for doing work.
@@
 Section 8 shows that EPHS models are thermodynamically consistent.
-Section 9 presents an example of a motor model, together with a comparison to a classical bond graph.
+Section 9 presents an example of a direct current (DC) motor model, along with a comparison to a classical bond graph.
 Section 10 concludes with a discussion.
@@
 Definition 2.2.
-Given two names $p$ and $n$, we say that $p$ is a prefix of $n$, and write $p\leq n$ if there exists any name $s$ such that $n=p\ast s$.
+Given two names $p$ and $n$, we say that $p$ is a prefix of $n$ and write $p\leq n$ if there exists any name $s$ such that $n=p\ast s$.
 In case $s\neq[]$, we say that $p$ is a strict prefix of $n$ and write $p<n$.
@@
 Example 2.3.
-It holds that $\mathtt{a.b}<\mathtt{a.b.c}$ and $\mathtt{a.b}\leq\mathtt{a.b}$. Also $[]<\mathtt{a}$. It does not hold that $\mathtt{a.a}\leq\mathtt{a}$. ∎
+It holds that $\mathtt{a.b}<\mathtt{a.b.c}$ and $\mathtt{a.b}\leq\mathtt{a.b}$. Also, $[]<\mathtt{a}$. It does not hold that $\mathtt{a.a}\leq\mathtt{a}$. ∎
 Definition 2.4.
@@
 A namespace is a finite, prefix-free set of names.
-Let $\mathbf{Namespace}$ refer to the set of all namespaces.
+Let $\mathsf{Namespace}$ refer to the set of all namespaces.
 ∎
@@
 ∎
-Sets of names are equivalent to prefix trees, called tries in computer science Fredkin (1960).
+Sets of names are equivalent to prefix trees, called tries in computer science [45].
 Example 2.6.
@@
 [Uncaptioned image][Uncaptioned $\qed$
-Note that for a namespace $N$, the leaf nodes of its trie correspond to the names in $N$, while for a set of names, which is not prefix-free, at least one name corresponds to an internal node.
+Note that for a namespace $N$, the leaf nodes of its trie correspond to the names in $N$, while for a set of names that is not prefix-free, at least one name corresponds to an internal node.
 In the following, we think of namespaces both as sets of names and as tries.
@@
 Definition 2.7.
-Given a namespace $N$, and a function $f\colon N\to\mathbf{Namespace}$, the named sum of $f$ is
+Given a namespace $N$, and a function $f\colon N\to\mathsf{Namespace}$, the named sum of $f$ is
 $\sum f\>=\>\Bigl{\{}\,n\ast a\,\mid\,n\in N,\>a\in f(n)\,\Bigr{\}}\,.\qed$
@@
 Example 2.8.
-Given $N=\{\mathtt{q},\,\mathtt{p.1},\,\mathtt{p.2}\}$, we have
-$\sum\left[\mathtt{l}\mapsto N,\,\mathtt{s.t}\mapsto N\right]\>=\>\bigl{\{}\mathtt{l.q},\,\mathtt{l.p.1},\,\mathtt{l.p.2},\,\mathtt{s.t.q},\,\mathtt{s.t.p.1},\,\mathtt{s.t.p.2}\bigr{\}}\,.\qed$
+Given $N=\{\mathtt{p},\,\mathtt{q}\}$, we have
+$\sum\left[\mathtt{l}\mapsto N,\,\mathtt{r}\mapsto N\right]\>=\>\bigl{\{}\mathtt{l.p},\,\mathtt{l.q},\,\mathtt{r.p},\,\mathtt{r.q}\bigr{\}}\,.\qed$
 Proposition 2.9.
-Given a namespace $N$ and a function $f\colon N\to\mathbf{Namespace}$, the named sum $\sum f$ is again a namespace, that is to say it is prefix-free. ∎
+Given a namespace $N$ and a function $f\colon N\to\mathsf{Namespace}$, the named sum $\sum f$ is again a namespace, that is to say, it is prefix-free. ∎
 Proof.
-The trie $\sum f$ is simply the trie $N$ with, for each leaf node $n\in N$, the trie $f(n)$ grafted on top. ∎
-Namespaces and named sums provide a simple and effective formalism for working with large hierarchically-structured systems.
-2.2 Packages
-An interface has a namespace, where we interpret each name as referring to a port.
-However, this is not the only ingredient to an interface, as each port has two associated attributes.
-This motivates the definition of packages:
+The trie $\sum f$ is simply the trie $N$ with, for each leaf node $n\in N$, the trie $f(n)$ grafted onto it. ∎
+As shown next, we can use namespaces and named sums to develop a simple and effective formalism for working with hierarchically-structured data.
+2.2 Directories
+The interface of a system is given by a namespace, where each name refers to a port. However, this is not all, as each port has two associated attributes. This motivates the definition of directories:
 Definition 2.10.
-Given a set $A$, we define
-$\mathbf{Pkg}(A)\>=\>\Bigl{\{}(N,\,\tau)\,\mid\,N\in\mathbf{Namespace},\,\tau\colon N\to A\Bigr{\}}\,.$
-An element $(N,\,\tau)\in\mathbf{Pkg}(A)$ is called a package of elements of $A$. We also just write $\tau\in\mathbf{Pkg}(A)$ with $N=\operatorname{dom}(\tau)$. ∎
-The name ‘package’ is drawn from software engineering; one might consider a software package to be a ‘package of types and procedures’.
-A package $(N,\,\tau)\in\mathbf{Pkg}(A)$ can be seen as a trie $N$ with an element $\tau(n)\in A$ associated to each leaf node $n\in N$.
-Hence, the concept of named sum carries over from $\mathbf{Namespace}$ to $\mathbf{Pkg}(A)$, for a fixed set $A$.
+Given a set $A$ (of possible attributes), we define
+$\mathsf{Dtry}(A)\>=\>\Bigl{\{}\,(N,\,\tau)\,\mid\,N\in\mathsf{Namespace},\,\tau\colon N\to A\,\Bigr{\}}\,.$
+An element $(N,\,\tau)\in\mathsf{Dtry}(A)$ is called a directory of elements of $A$. We also just write $\tau\in\mathsf{Dtry}(A)$ with $N=\operatorname{dom}(\tau)$. ∎
+The name ‘directory’ points to the similarity with hierarchical file systems, where complete paths (names) are associated to file contents (attributes).
+A directory $(N,\,\tau)\in\mathsf{Dtry}(A)$ can be seen as a trie $N$, where each leaf node $n\in N$ has an associated attribute $\tau(n)\in A$.
+Hence, the concept of named sum carries over from $\mathsf{Namespace}$ to $\mathsf{Dtry}(A)$, for any fixed set $A$.
 Named sum can then be seen as a flattening operation, analogous to flattening a list of lists.
 Specifically, for the set of all lists of elements of $A$
-$\operatorname{\mathbf{List}}(A)\>=\>\bigl{\{}(n,\,f)\,\mid\,n\in\mathbb{N},\,f\colon\{1,\,\ldots,\,n\}\to A\bigr{\}}\,,$
-flattening is a map $\operatorname{\mathbf{List}}(\operatorname{\mathbf{List}}(A))\to\operatorname{\mathbf{List}}(A)$, which for instance sends [[1, 2],[3, 4]]1234[[1,\,2],\,[3,\,4]][ [ 1 , 2 ] , [ 3 , 4 ] ] to [1, 2, 3, 4]1234[1,\,2,\,3,\,4][ 1 , 2 , 3 , 4 ].
-Similarly, named sum is a map $\mathbf{Pkg}(\mathbf{Pkg}(A))$ to $\mathbf{Pkg}(A)$.
+$\mathsf{List}(A)\>=\>\Bigl{\{}\,(n,\,f)\,\mid\,n\in\mathbb{N},\,f\colon\{1,\,\ldots,\,n\}\to A\,\Bigr{\}}\,,$
+flattening is a map $\mathsf{List}(\mathsf{List}(A))\to\mathsf{List}(A)$, which for instance sends [[1, 2],[3, 4]]1234[[1,\,2],\,[3,\,4]][ [ 1 , 2 ] , [ 3 , 4 ] ] to [1, 2, 3, 4]1234[1,\,2,\,3,\,4][ 1 , 2 , 3 , 4 ].
+Similarly, named sum is a map $\mathsf{Dtry}(\mathsf{Dtry}(A))$ to $\mathsf{Dtry}(A)$.
 An input to this is a trie $N$, which has associated to each leaf node $n\in N$ another trie $N_{n}$, whose leaves carry elements of $A$.
 The named sum operation then simply grafts onto each node $n\in N$ its associated trie $N_{n}$, carrying along the elements of $A$ at the leaves.
-In terms of category theory, the flattening operation makes $\operatorname{\mathbf{List}}$ and $\mathbf{Pkg}$ examples of monads.
+In terms of category theory, the flattening operation, together with a unit operation, makes $\mathsf{List}$ and $\mathsf{Dtry}$ examples of monads.
 2.3 Bundle of quantities
-The first attribute associated to any port represents the physical quantity that can be exchanged via the port, e.g.
-entropy or momentum, along with a state space in which values of the quantity live.
-For instance a momentum variable can take values in $\mathbb{R}$ if motion along a single axis is considered or it can take values in the vector space $\mathfrak{se}^{*}(3)$ if rigid body motions in $3$-dimensional Euclidian space are considered, etc.
+The first attribute associated to a port represents the physical quantity that can be exchanged via the port, e.g.
+momentum or entropy, along with a state space in which values of the quantity live.
+For instance a momentum variable can take values in $\mathbb{R}$ if motion along a single axis is considered, or it can take values in the vector space $\mathfrak{se}^{*}(3)$ if rigid body motions in $3$-dimensional Euclidean space are considered, etc.
 Definition 2.11.
@@
 ∎
-To distinguish reversible and irreversible dynamics later on in Sections 6.3 and 6.4, we make use of the concept of parity, see Pavelka et al. (2014); Lohmayer et al. (2021) for details.
+We note that $(\mathbb{R},\,\mathtt{energy})\notin\mathbf{Q}$, because in the context of EPHS, internal energy is a function of the state variable entropy, and not vice versa, see also [3].
+To distinguish reversible and irreversible dynamics later on in Sections 6.3 and 6.4, we make use of the concept of parity, see [46, 3] for details.
 Definition 2.13.
-Let the parity of each quantity with respect to time-reversal transformation be given by a function $\mathrm{P}:\mathbf{Q}\rightarrow\{-1,\,+1\}$. ∎
+The parity of each quantity with respect to time-reversal transformation is given by a function $\mathrm{P}:\mathbf{Q}\rightarrow\{-1,\,+1\}$. ∎
 Example 2.14.
@@
 because displacement, entropy, and volume do not change instantaneously when a recording is suddenly played backwards (positive parity), whereas momentum changes its sign (negative parity). ∎
-A comprehensive collection of physical quantities with their associated parities would be a built-in feature of a software implementation.
-2.4 Interfaces as packages
-Each port has two attributes, namely the already discussed quantity and further a Boolean attribute that indicates whether the port is a power port or a state port.
-Only power ports allow for exchange of energy associated with the attributed quantity.
-A state port is required whenever the behavior of a system depends on the state of another system, although no energy is exchanged between the two.
+A comprehensive collection of physical quantities with their associated parities would be part of a software implementation.
+2.4 Interfaces as directories
+Each port has two attributes: the previously discussed quantity, and a Boolean attribute that indicates whether the port is a power port or a state port.
+Indeed, only power ports allow for an exchange of energy, based on an exchange of the respective quantity.
+State ports are required whenever the behavior of one system depends on the state of another system, although no energy is exchanged between the two.
 Definition 2.15.
-An interface is a package of elements of $\mathbf{Q}\times\{\mathsf{p},\mathsf{s}\}$, i.e. let
-$\mathbf{Interface}\>=\>\mathbf{Pkg}\bigl{(}\mathbf{Q}\times\{\mathsf{p},\mathsf{s}\}\bigr{)}\,.\qed$
+An interface is a directory of elements of $\mathbf{Q}\times\{\mathsf{p},\mathsf{s}\}$, i.e. let
+$\mathsf{Interface}\>=\>\mathsf{Dtry}\bigl{(}\mathbf{Q}\times\{\mathsf{p},\mathsf{s}\}\bigr{)}\,.$
+Here, element $\mathsf{p}$ stands for power port and $\mathsf{s}$ stands for state port. ∎
 Example 2.16.
 Let $I=(N,\,\tau)$ be an interface with namespace $N=\{\mathtt{p_{1}},\,\mathtt{p_{2}},\,\mathtt{q}\}$ and attributes given by $\tau(\mathtt{p_{1}})=\tau(\mathtt{p_{2}})=\left((\mathbb{R},\,\mathtt{momentum}),\,\mathsf{p}\right)$ and $\tau(\mathtt{q})=\left((\mathbb{R},\,\mathtt{displacement}),\,\mathsf{s}\right)$.
-Ports $\mathtt{p_{1}}$ and $\mathtt{p_{2}}$ are power ports that can exchange kinetic energy, along with information about the current momentum.
+Ports $\mathtt{p_{1}}$ and $\mathtt{p_{2}}$ are power ports that can exchange kinetic energy by exchanging momentum, along with information about the current amount of momentum.
 Port $\mathtt{q}$ is a state port that can only exchange information about the current displacement.
 ∎
-We note that the named sum of interfaces is a function $\mathbf{Pkg}(\mathbf{Interface})\to\mathbf{Interface}$.
 2.5 Bundles of port variables
 A state port has one port variable called state.
-A power port has two further port variables called flow and effort.
-Given an interface, the port variables of all its ports combined together take values in a vector bundle, called the bundle of port variables associated to the interface.
+A power port has two additional port variables called flow and effort.
+Given an interface, the port variables of all its ports combined take values in a vector bundle, called the bundle of port variables associated to the interface.
 Definition 2.17.
@@
 where $\mathrm{pr}_{1}$ projects onto the first attribute, i.e. the quantity.
-The state space associated to interface $I$ is the Cartesian product
-$\mathcal{X}_{I}\>=\>\prod_{p\,\in\,I}\mathcal{X}_{p}\,.\qed$
+The state space associated to interface $I$ is given, up to isomorphism, by the Cartesian product
+$\mathcal{X}_{I}\>=\>\prod_{p\,\in\,I}\mathcal{X}_{p}\,.$
+To define $\mathcal{X}_{I}$ strictly, we assume that the elements of $I$ are in lexicographic order and that the Cartesian product preserves this order. ∎
 Example 2.18.
@@
 As detailed in Section 6.1, the flow variable is the instantaneous rate of change of the state variable and the effort variable is the differential of a function of the state variable.
-Hence, we have $\left(p\mathtt{.x},\,p\mathtt{.f},\,p\mathtt{.e}\right)\in\mathbb{T}\mathcal{X}_{p}$, where $\mathbb{T}\mathcal{X}_{p}=\mathrm{T}\mathcal{X}_{p}\oplus\mathrm{T}^{*}\mathcal{X}_{p}$ denotes the direct sum (or Whitney sum) of the respective tangent and cotangent bundles (see e.g.
-Lee (2012)).
-The flow and effort variables are called power variables because their duality pairing $\langle p\mathtt{.e}\mid p\mathtt{.f}\rangle\in\mathbb{R}$ yields the instantaneous power that is exchanged through port $p$.
-Now we can define the bundle of port variables associated to an interface:
+Hence, we have $\left(p\mathtt{.x},\,p\mathtt{.f},\,p\mathtt{.e}\right)\in\mathbb{T}\mathcal{X}_{p}$, where $\mathbb{T}\mathcal{X}_{p}=\mathrm{T}\mathcal{X}_{p}\oplus\mathrm{T}^{*}\mathcal{X}_{p}$ denotes the direct sum (or Whitney sum) of the respective tangent and cotangent bundles (see, e.g., [47]).
+The flow and effort variables are called power variables because their duality pairing $\langle p\mathtt{.e}\mid p\mathtt{.f}\rangle$ yields the instantaneous power that is exchanged through port $p$.
+Now we can formally define the bundle of port variables associated to an interface:
 Definition 2.19.
 Let $I$ be an interface.
-Based on the Boolean attribute of its ports, we can write its associated state space as the Cartesian product $\mathcal{X}_{I}\cong\mathcal{X}_{I,p}\times\mathcal{X}_{I,s}$, where $\mathcal{X}_{I,p}$ is the state space of all power ports and $\mathcal{X}_{I,s}$ is the state space of all state ports.
-Accordingly, we let $x=(x_{p},\,x_{s})\in\mathcal{X}_{I,p}\times\mathcal{X}_{I,s}$ denote all state variables and we let $(f,\,e)\in\mathbb{T}_{x_{p}}\mathcal{X}_{I,p}$ denote all flow and effort variables.
-The bundle of port variables $\mathrm{pr}_{I}:\mathcal{P}_{I}\twoheadrightarrow\mathcal{X}_{I}$ associated to interface $I$ is defined as the pullback bundle of $\mathbb{T}\mathcal{X}_{I,p}$ along the projection $\mathrm{pr}_{p}:\mathcal{X}_{I,p}\times\mathcal{X}_{I,s}\twoheadrightarrow\mathcal{X}_{I,p}$.
+Based on the Boolean attribute of its ports, we can think of its associated state space $\mathcal{X}_{I}$ as the Cartesian product $\mathcal{X}_{I,p}\times\mathcal{X}_{I,s}$, where $\mathcal{X}_{I,p}$ is the state space of all power ports and $\mathcal{X}_{I,s}$ is the state space of all state ports.
+Accordingly, we let $x=(x_{p},\,x_{s})\in\mathcal{X}_{I,p}\times\mathcal{X}_{I,s}$ denote all state variables, and we let $(f,\,e)\in\mathbb{T}_{x_{p}}\mathcal{X}_{I,p}$ denote the flow and effort variables.
+The bundle of port variables $\mathrm{pr}_{I}:\mathcal{P}_{I}\twoheadrightarrow\mathcal{X}_{I}$ associated to interface $I$ is defined, up to isomorphism, as the pullback bundle of $\mathbb{T}\mathcal{X}_{I,p}$ along the projection $\mathrm{pr}_{p}:\mathcal{X}_{I,p}\times\mathcal{X}_{I,s}\twoheadrightarrow\mathcal{X}_{I,p}$.
 Hence, $\mathcal{P}_{I}\cong\mathbb{T}\mathcal{X}_{I,p}\times\mathcal{X}_{I,s}$.
-{$\begin{tikzcd}\qed$
+[…]
+Again, to define the bundle of port variables strictly, we assume that the elements of $I$ are in lexicographic order and that the port variables are written in the same order as the ports. ∎
 Example 2.20.
-Consider again the interface from 2.16. Its bundle of port variables $\mathcal{P}_{I}\twoheadrightarrow\mathcal{X}_{I}$ has total space (or port space)
+Consider again the interface from 2.16. Its bundle of port variables $\mathcal{P}_{I}\twoheadrightarrow\mathcal{X}_{I}$ has the total space (or port space)
 $\mathcal{P}_{I}\>=\>\mathbb{T}\mathbb{R}\times\mathbb{T}\mathbb{R}\times\mathbb{R}\>\ni\>\bigl{(}(\mathtt{p_{1}.x},\,\mathtt{p_{1}.f},\,\mathtt{p_{1}.e}),\,(\mathtt{p_{2}.x},\,\mathtt{p_{2}.f},\,\mathtt{p_{2}.e}),\,\mathtt{q.x}\bigl{)}$
-and base space (or state space)
+and the base space (or state space)
 $\mathcal{X}_{I}\>=\>\mathbb{R}\times\mathbb{R}\times\mathbb{R}\>\ni\>\left(\mathtt{p_{1}.x},\,\mathtt{p_{2}.x},\,\mathtt{q.x}\right).\qed$
-3 Interconnection patterns
-As introduced in Section 1.2, an interconnection pattern expresses how a composite system can be formed from a finite number of subsystems.
-The syntax defines the entirety of such expressions and their composition, i.e.
+3 Syntax
+As introduced in Section 1.2, an interconnection pattern expresses how a composite system is formed from a finite number of subsystems.
+Syntax refers to the entirety of such expressions and their composition, i.e.
 their hierarchical nesting.
@@
 Figure 4: Interconnection pattern of an isothermal oscillator model.
-The box $\mathtt{pkc}$ represents the coupling of the potential energy domain represented by the junction on its left and the kinetic energy domain represented by the junction on its right.
-The box $\mathtt{env}$ represents the isothermal environment which absorbs the heat that is generated by the damping.
-The port $\mathtt{p}$ of the outer box exposes the kinetic energy domain, allowing for an external forcing of the oscillator model.
-Figure 4 shows an interconnection pattern. The mathematical content of such a diagram is captured by the following definition:
+Box $\mathtt{pkc}$ represents the reversible coupling of the potential energy domain represented by the junction on its left and the kinetic energy domain represented by the junction on its right.
+Box $\mathtt{env}$ represents the isothermal environment, which absorbs the heat that is generated by the damping.
+The outer box represents the interface of the composite system.
+Its port $\mathtt{p}$ exposes the kinetic energy domain, allowing for an external forcing of the oscillator model.
+Figure 4 shows an interconnection pattern. The mathematical content of such a diagram is captured by the following definition.
 Definition 3.1.
@@
   1. 1.
-    A package of interfaces $\beta\in\mathbf{Pkg}(\mathbf{Interface})$.
+    A directory of interfaces $\beta\in\mathsf{Dtry}(\mathsf{Interface})$.
 Names in $\operatorname{dom}(\beta)$ refer to inner boxes, $\sum\beta$ is called inner interface and names in $\operatorname{dom}(\sum\beta)$ refer to inner ports.
   2. 2.
-    An interface $I$ called outer interface with ports called outer ports.
+    An interface $I$ called outer interface, whose ports are called outer ports.
   3. 3.
@@
     • (c)
-      and all ports have the same associated quantity. ∎
-Although the names of inner ports and outer ports are distinguished by a prefix, usually the namespaces of $\sum\beta$ and $I$ are already disjoint.
-We then abuse notation and simply write for instance $\mathtt{p}$ rather than $\mathtt{outer.p}$.
+      and all ports in $j$ have the same associated quantity. ∎
+To prevent name conflicts, the definition distinguishes inner ports and outer ports with the prefixes $\mathtt{inner}$ and $\mathtt{outer}$.
+While preventing name conflicts is crucial for implementation of the modeling language, we omit these prefixes for conciseness in what follows.
 We graphically depict interconnection patterns as follows.
@@
 The elements of $j\subseteq P$ are precisely the connected ports.
-To avoid confusion, we recall that the identifier isothermal_osc written on the top left of the outer box refers to the entire pattern (akin to a function name), rather than just to the outer box (which similar to the return value of a function does not need a name).
+To avoid confusion, we recall that the identifier isothermal_osc written on the top left of the outer box refers to the entire pattern (akin to a function name in programming), rather than just to the outer box (which similar to the return value of a function does not need a name).
 The pattern from Figure 4 is redrawn in Figure 5, firstly to emphasize that the color of the inner boxes is merely an annotation hinting at the nature of the considered subsystems and secondly to introduce an abbreviated way of writing port names.
-Figure 5: Interconnection pattern of an isothermal oscillator model. Whenever all ports connected to a certain junction have the same name we write the name only once at the junction.
+Figure 5: Interconnection pattern of an isothermal oscillator model. Whenever all ports connected to a certain junction have the same name, we write the name only once at the junction.
 Example 3.2.
 For the pattern $(\beta,\,I,\,J)$ depicted in Figure 5, the namespace of inner boxes is
-$\operatorname{dom}(\beta)\>=\>\{\mathtt{spring},\,\mathtt{pkc},\,\mathtt{mass},\,\mathtt{damping},\,\mathtt{env}\}\,.$
-The function $\beta:\operatorname{dom}(\beta)\rightarrow\mathbf{Interface}$ is given by
-$\begin{split}\mathtt{spring}\>&\mapsto\>(\{\mathtt{q}\},\,\tau_{\text{spring}})\\ \mathtt{pkc}\>&\mapsto\>(\{\mathtt{q},\,\mathtt{p}\},\,\tau_{\text{pkc}})\\ \mathtt{mass}\>&\mapsto\>(\{\mathtt{p}\},\,\tau_{\text{mass}})\\ \mathtt{damping}\>&\mapsto\>(\{\mathtt{p},\,\mathtt{s}\},\,\tau_{\text{damping}})\\ \mathtt{env}\>&\mapsto\>(\{\mathtt{s}\},\,\tau_{\text{env}})\end{split}$
-with $\tau_{\text{spring}}(\mathtt{q})=((\mathbb{R},\,\mathtt{displacement}),\,\mathtt{p})$, etc. Hence, the namespace of the inner interface $\sum\beta$ is
-$\operatorname{dom}(\textstyle\sum\beta)\>=\>\{\mathtt{spring.q},\,\mathtt{pkc.q},\,\mathtt{pkc.p},\,\mathtt{mass.p},\,\mathtt{damping.p},\,\mathtt{damping.s},\,\mathtt{env.s}\}\,.$
+$\operatorname{dom}(\beta)\>=\>\{\mathtt{damping},\,\mathtt{env},\,\mathtt{mass},\,\mathtt{pkc},\,\mathtt{spring}\}\,.$
+The function $\beta:\operatorname{dom}(\beta)\rightarrow\mathsf{Interface}$ is given by
+$\begin{split}\mathtt{damping}\>&\mapsto\>(\{\mathtt{p},\,\mathtt{s}\},\,\tau_{\text{damping}})\\ \mathtt{env}\>&\mapsto\>(\{\mathtt{s}\},\,\tau_{\text{env}})\\ \mathtt{mass}\>&\mapsto\>(\{\mathtt{p}\},\,\tau_{\text{mass}})\\ \mathtt{pkc}\>&\mapsto\>(\{\mathtt{p},\,\mathtt{q}\},\,\tau_{\text{pkc}})\\ \mathtt{spring}\>&\mapsto\>(\{\mathtt{q}\},\,\tau_{\text{spring}})\end{split}$
+with $\tau_{\text{damping}}(\mathtt{p})=((\mathbb{R},\,\mathtt{momentum}),\,\mathsf{p})$, etc. Hence, the namespace of the inner interface $\sum\beta$ is
+$\operatorname{dom}(\textstyle\sum\beta)\>=\>\{\mathtt{damping.p},\,\mathtt{damping.s},\,\mathtt{env.s},\,\mathtt{mass.p},\,\mathtt{pkc.p},\,\mathtt{pkc.q},\,\mathtt{spring.q}\}\,.$
 The namespace of the outer interface $I$ is $\{\mathtt{p}\}$. The partition $J$ is given by
-$\bigl{\{}\{\mathtt{spring.q},\,\mathtt{pkc.q}\},\,\{\mathtt{mass.p},\,\mathtt{pkc.p},\,\mathtt{damping.p},\,\mathtt{p}\},\,\{\mathtt{env.s},\,\mathtt{damping.s}\}\bigr{\}}\,.\qed$
+$\bigl{\{}\{\mathtt{damping.p},\,\mathtt{mass.p},\,\mathtt{pkc.p},\,\mathtt{p}\},\,\{\mathtt{pkc.q},\,\mathtt{spring.q}\},\,\{\mathtt{damping.s},\,\mathtt{env.s}\}\bigr{\}}\,.\qed$
 We may also denote an interconnection pattern using the condensed notation
-[…]
+${\sum\left[\mathtt{I_{1}}\mapsto I_{1},\,\ldots,\,\mathtt{I_{n}}\mapsto I_{n}\right]}$${J}$${I.}$
 In this we implicitly let $\operatorname{dom}(\beta)=\{\mathtt{I_{1}},\,\ldots,\,\mathtt{I_{n}}\}$ and $\beta(\mathtt{I_{i}})=I_{i}$.
 The left arrow is a surjection because each junction has at least one connected inner port and the right arrow is an injection because each junction has at most one connected outer port.
-Interconnection patterns are a special case of undirected wiring diagrams Spivak (2013), which do not have the injectivity and surjectivity conditions.
-As shown in Figure 1, interconnection patterns may have another feature that is not formalized by 3.1, namely port combinators for combining ports into multiports.
-Although namespaces seem to provide an adequate basis to formalize multiports and port combinators, we defer this to later work.
-At this point, we simply regard them as ad-hoc syntactic sugar for expressing parallel connections in a tidy and efficient manner.
+Interconnection patterns are a special case of undirected wiring diagrams [16], which do not have the injectivity and surjectivity conditions.
+Multiports are just multiple ports with a common prefix.
+As demonstrated in Figure 2, this enables a syntactic sugar for expressing parallel connections in a tidy and efficient manner.
+We defer the details to later work, as the examples considered here make no use of this.
 3.2 Composition of interconnection patterns
-Whenever the outer interface of one pattern and the interface corresponding to an inner box of another pattern are equivalent up to renaming of the ports, the former pattern can be substituted into the latter, with no further data required.
+Whenever the outer interface of one pattern and the interface corresponding to an inner box of another pattern are equal (up to a given renaming of the ports), the former pattern can be substituted into the latter, with no further data required.
 This is called composition.
-Having a compositional syntax facilitates dealing with complex systems, as it allows to encapsulate subsystems, which can be easily reused and replaced, as long as interfaces match.
-As an example, the isothermal oscillator model can be refined into a nonisothermal oscillator model by substituting a more detailed damper model.
+Having a compositional syntax facilitates dealing with complex systems, as it enables the encapsulation of subsystems.
+These can be easily reused and replaced, as long as interfaces match.
+As an example, the isothermal oscillator model can be refined into a nonisothermal oscillator model based on the damper model in Figure 6.
 Figure 6: Interconnection pattern of a nonisothermal damper model.
-The box $\mathtt{mf}$ represents a mechanical friction model connected to the kinetic energy domain represented by the outer port $\mathtt{p}$.
-The box $\mathtt{tc}$ represents a thermal capacity, which can store the heat generated by friction.
-The box $\mathtt{ht}$ represents a heat transfer model, which allows the thermal capacity to reach equilibrium with the thermal energy domain represented by the outer port $\mathtt{s}$.
-Figure 7: Interconnection pattern of a nonisothermal oscillator model.
+Box $\mathtt{mf}$ represents a mechanical friction model connected to the kinetic energy domain represented by the outer port $\mathtt{p}$.
+Box $\mathtt{tc}$ represents a thermal capacity, which stores the heat generated by friction.
+Box $\mathtt{ht}$ represents a heat transfer model, which allows the thermal capacity to reach equilibrium with the thermal energy domain represented by the outer port $\mathtt{s}$.
+Figure 7: Interconnection pattern of a nonisothermal oscillator model. Except for the graphical layout, the pattern is uniquely defined as the composite of the patterns in Figures 5 and 6.
 Substituting the pattern in Figure 6 into the pattern in Figure 5, yields the pattern in Figure 7.
 Graphically, the former pattern is inserted into the inner box called $\mathtt{damping}$ of the latter pattern.
-In this step, the named sum for interfaces is used.
-Then, the common interface is deleted and for each of its ports, the two assigned junctions are identified.
-Category theory provides a language to talk about composition and it includes well-behavedness conditions, which ensure that composition works in the way we expect, just as we need not worry about the difference between $(1+2)+1$ and $1+(2+1)$.
-In a following paper, we will formalize this using the language of operads.
+The common interface is deleted and for each of its ports, the two assigned junctions (one in each pattern) are identified.
+The named sum for interfaces is used to obtain the inner interface of the composed pattern.
+Category theory provides a language to talk about composition, and it includes well-behavedness conditions, which ensure that composition works in the way we expect, just as we need not worry about the difference between $(1+2)+3$ and $1+(2+3)$.
+A future paper shall formalize the composition of interconnection patterns based on directories and multicategories/operads.
+To limit the scope of this article, we just mention that it is not difficult to see why the procedure for composition outlined in the previous paragraph preserves the three conditions in 3.1.
 4 Semantics of interconnection patterns
-In this section, we discuss how to interpret an interconnection pattern as imposing a relation between the port variables of its inner and outer interface.
+In Section 2.5 we have defined bundles of port variables associated to interfaces.
+In Section 3.1 we have defined interconnection patterns, which have an inner interface (containing all ports of the subsystems) and an outer interface (containing the ports of the composite system).
+We now discuss how to interpret an interconnection pattern as imposing a relation between the port variables of its inner interface and the port variables of its outer interface.
 We do this in two parts.
-In the first part, we talk about the relation as a collection of equations and discuss its physical interpretation.
+In the first part, we talk about the relation as a collection of equations, and we discuss its physical interpretation.
 In the second part, we reify it as a geometric object, namely as a relation between bundles of port variables.
 4.1 Interconnection patterns as equations
-Informally, the semantics of an interconnection pattern $(\beta,\,I,\,J)$ can be seen as a collection of equations. Specifically, for each junction $j\in J$, we have the following:
+The semantics of an interconnection pattern $(\beta,\,I,\,J)$ can be seen as a collection of equations. Specifically, for each junction $j\in J$, we have the following.
 Let $j_{p}\subseteq j$ be the subset of connected power ports.
@@
     $\sum_{p\,\in\,j_{p}\cap\sum\beta}p\mathtt{.f}\>=\>\sum_{p\,\in\,j_{p}\cap I}p\mathtt{.f}$
-As detailed in Section 7, combining these equations with equations that determine the semantics of each of the subsystems filling the inner boxes yields a collection of equations that determines the semantics of the resulting composite system.
+We note that the sum on the left-hand side has at least one term (surjectivity condition) and the sum on the right-hand side has at most one term (injectivity condition).
+As detailed in Section 7, combining these equations with equations that determine the semantics of the subsystems filling the inner boxes yields a collection of equations that determines the semantics of the resulting composite system.
 Since all ports connected to a junction have the same quantity, their associated state variables take values in the same space, making equality of state a well-typed equation.
-As a consequence of equality of state, all effort variables take values in the same fiber of the same cotangent bundle, making equality of effort also well-typed.
+As a consequence of this, all effort variables take values in the same fiber of the same cotangent bundle, making equality of effort also well-typed.
 Similarly, all flow variables take values in the same fiber of the same tangent bundle, making equality of net flow well-typed.
-Also note that in the equality of net flow, the inner ports are on the left hand side and the outer ports are on the right hand side.
-As a consequence, at every junction $j\in J$ the following power balance equation is satisfied:
+Equality of state is required, since the behavior of one system may in general depend on the state of another system, even if the two systems do not exchange energy.
+We refer to Figures 8 and 9 for an example of this.
+As a consequence of equality of effort and equality of net flow, at every junction $j\in J$ the following power balance equation is satisfied:
 $\sum_{p\,\in\,j_{p}\cap\sum\beta}\langle p\mathtt{.e}\mid p\mathtt{.f}\rangle\>=\>\sum_{p\,\in\,j_{p}\cap I}\langle p\mathtt{.e}\mid p\mathtt{.f}\rangle$
-Summing over all junctions on both sides, we get that the net power supplied to all subsystems is equal to the net power supplied to the composite system.
-Taking into account also the choice of a sign made in Sections 6.1 and 6.2, we see that this agrees with the sign convention used in thermodynamics, stating that power supplied to any system, be it a subsystem or the composite system, is counted with a positive sign.
-Junctions correspond to energy domains.
-Hence, an outer port exposes an energy domain of the respective composite system.
-Since this interpretation also applies to the primitive systems discussed in Section 6, we can say that ports also represent energy domains and whenever two ports are connected to the same junction the energy domains are identified.
-In other words, interconnection means that systems share common energy domains.
+Summing over all junctions on both sides gives that the net power supplied to all subsystems is equal to the net power supplied to the composite system.
+Junctions correspond to energy domains and so do (outer) ports, since they expose energy domains.
+Whenever two ports are connected to the same junction the respective energy domains are identified.
+This is exactly how interconnection patterns compose.
+Interconnection means that systems share energy domains.
 The state of a shared energy domain is given by the shared state variable, according to the equality of state.
 The connected systems may exchange the respective physical quantity according to the equality of net flow.
-Moreover, such an exchange also implies an exchange of energy, according to the equality of effort.
+Moreover, based on the equality of effort, this exchange implies a transfer of (exergetic) power.
 Example 4.1.
-Let’s again consider the pattern in Figure 7. At the junction representing the kinetic energy domain, the momentum variable (state) is shared and forces (flows) balance at equal velocity (effort):
+Let’s again consider the pattern in Figure 7. At the junction which represents the kinetic energy domain, the momentum variable (state) is shared and forces (flows) balance at equal velocity (effort):
 […]
@@
 Table 1: Physical interpretation of junctions as energy domains.
-We can now briefly reconsider the surjectivity and injectivity conditions which distinguish interconnection patterns form undirected wiring diagrams.
+We can now make sense of the surjectivity and injectivity conditions distinguishing interconnection patterns form undirected wiring diagrams.
 The surjectivity condition says that each junction must have at least one connected inner port.
-In other words, an energy domain of the composite system must belong to at least one of its subsystems.
-Of course, an energy domain that is not part of the system cannot be exposed.
+In other words, an energy domain can only belong to a composite system if it belongs to at least one subsystem.
 The injectivity condition says that each junction may have at most one connected outer port.
 In other words, an energy domain can be exposed at most once.
-Exposing an energy domain twice and identifying it again at a higher level of nesting by connecting the two respective ports would lead to ill-defined flow variables.
+Exposing an energy domain twice and identifying it again at a higher level by connecting the two respective ports would lead to undefined flow variables.
 4.2 Interconnection patterns as relations
 […]
-Similarly to this prototypical example, the relation assigned to an interconnection pattern is given by a subbundle of the Cartesian product of the two bundles of port variables associated to the inner and the outer interface.
+Analogous to the prototypical example of a relation between two sets, the relation assigned to an interconnection pattern is defined by a vector subbundle of the Cartesian product of the bundle of port variables associated to the inner interface and the bundle of port variables associated to the outer interface.
 Definition 4.2.
 Let $P=(\beta,\,I,\,J)$ be an interconnection pattern and let $I_{\text{in}}=\sum\beta$ denote its inner interface.
-Further, let $\mathcal{P}_{I_{\text{in}}}\twoheadrightarrow\mathcal{X}_{I_{\text{in}}}$ be the bundle of port variables associated to the inner interface and let $\mathcal{P}_{I}\twoheadrightarrow\mathcal{X}_{I}$ be the bundle of port variables associated to the outer interface.
+Further, let $\mathcal{P}_{I_{\text{in}}}\twoheadrightarrow\mathcal{X}_{I_{\text{in}}}$ be the bundle of port variables associated to the inner interface, and let $\mathcal{P}_{I}\twoheadrightarrow\mathcal{X}_{I}$ be the bundle of port variables associated to the outer interface.
 Based on the partition $J$, let $\mathcal{P}_{P}\subset\mathcal{P}_{I_{\text{in}}}\times\mathcal{P}_{I}$ be the subspace where equality of state, equality of effort, and equality of net flow are satisfied.
 Similarly, let $\mathcal{X}_{P}\subset\mathcal{X}_{I_{i}}\times\mathcal{X}_{I}$ be the subspace where equality of state is satisfied.
-Then, the semantics assigned to interconnection pattern $P$ is the relation $\mathcal{P}_{P}:\mathcal{P}_{I_{\text{in}}}\nrightarrow\mathcal{P}_{I}$ between the port variables of the inner interface and the port variables of the outer interface that is given by the vector subbundle $\mathcal{P}_{P}\twoheadrightarrow\mathcal{X}_{P}$.
+Then, the semantics assigned to interconnection pattern $P$ is the relation $\mathcal{P}_{P}:\mathcal{P}_{I_{\text{in}}}\nrightarrow\mathcal{P}_{I}$ between the port variables of the inner interface and the port variables of the outer interface that is defined by the vector subbundle $\mathcal{P}_{P}\twoheadrightarrow\mathcal{X}_{P}$.
 ∎
-It can be shown that this assignment of semantics is functorial, which means that it makes no difference whether one first composes some interconnection patterns and then asks for the relation assigned to the composite or one first asks for the relations assigned to the original patterns and then composes these relations.
-Figure 8: Interconnection pattern of a DC shunt motor model.
+Figure 8: Interconnection pattern of the DC shunt motor model discussed in Section 9.
 The stator and rotor are connected to a common electric energy domain that is exposed via the outer port $\mathtt{q}$.
 The kinetic energy domain of the rotor is exposed via the outer port $\mathtt{p}$.
-The conversion between electric and kinetic energy in the rotor depends on the magnetic energy domain of the stator, which is represented by the junction with connected state ports named $\mathtt{b_{k}}$.
-The full model is discussed in Section 9.
+The conversion between electric and kinetic energy in the rotor depends on the magnetic energy domain of the stator, which is represented by the junction with two connected state ports named $\mathtt{b_{s}}$.
 Example 4.3.
-Consider the interconnection pattern $P=(\beta,\,I,\,J)$ depicted in Figure 8 with names of inner boxes given by $\operatorname{dom}(\beta)=\{\mathtt{stator},\,\mathtt{rotor}\}$ and interfaces given by
-$\beta(\mathtt{stator})\>=\>\left(\{\mathtt{q},\,\mathtt{b_{s}}\},\,\tau_{\text{s}}\right)$
-$\beta(\mathtt{rotor})\>=\>\left(\{\mathtt{q},\,\mathtt{p},\,\mathtt{b_{s}}\},\,\tau_{\text{r}}\right)$
-$I\>=\>\left(\{\mathtt{q},\,\mathtt{p}\},\,\tau\right)\,,$
+Consider the interconnection pattern $P=(\beta,\,I,\,J)$ depicted in Figure 8 with names of inner boxes given by $\operatorname{dom}(\beta)=\{\mathtt{rotor},\,\mathtt{stator}\}$ and interfaces given by
+$\beta(\mathtt{rotor})\>=\>\left(\{\mathtt{b_{s}},\,\mathtt{p},\,\mathtt{q}\},\,\tau_{\text{r}}\right)$
+$\beta(\mathtt{stator})\>=\>\left(\{\mathtt{b_{s}},\,\mathtt{q}\},\,\tau_{\text{s}}\right)$
+$I\>=\>\left(\{\mathtt{p},\,\mathtt{q}\},\,\tau\right)\,,$
 where
-$\tau_{\text{s}}(\mathtt{q})\>=\>\tau_{\text{r}}(\mathtt{q})\>=\>\tau(\mathtt{q})\>=\>((\mathbb{R},\,\mathtt{charge}),\,\mathsf{p})$
-$\tau_{\text{s}}(\mathtt{b_{s}})\>=\>\tau_{\text{r}}(\mathtt{b_{s}})\>=\>((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{s})$
-$\tau_{\text{r}}(\mathtt{p})\>=\>\tau(\mathtt{p})\>=\>((\mathbb{R},\,\mathtt{angular\_momentum}),\,\mathsf{p})\,.$
-Furthermore, the interconnection is determined by
-$J\>=\>\bigl{\{}\{\mathtt{stator.q},\,\mathtt{rotor.q},\,\mathtt{q}\},\,\{\mathtt{stator.b_{s}},\,\mathtt{rotor.b_{s}}\},\,\{\mathtt{rotor.p},\,\mathtt{p}\}\bigr{\}}\,.$
+$\tau_{\text{r}}(\mathtt{b_{s}})\>=\>\tau_{\text{s}}(\mathtt{b_{s}})\>=\>((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{s})$
+$\tau_{\text{r}}(\mathtt{p})\>=\>\tau(\mathtt{p})\>=\>((\mathbb{R},\,\mathtt{angular\_momentum}),\,\mathsf{p})$
+$\tau_{\text{r}}(\mathtt{q})\>=\>\tau_{\text{s}}(\mathtt{q})\>=\>\tau(\mathtt{q})\>=\>((\mathbb{R},\,\mathtt{charge}),\,\mathsf{p})\,.$
+Further, the interconnection is determined by
+$J\>=\>\bigl{\{}\{\mathtt{rotor.b_{s}},\,\mathtt{stator.b_{s}}\},\,\{\mathtt{rotor.p},\,\mathtt{p}\},\,\{\mathtt{rotor.q},\,\mathtt{stator.q},\,\mathtt{q}\}\bigr{\}}\,.$
 The port space of the inner interface $I_{\text{in}}=\sum\beta$ is
-$\begin{split}\mathcal{P}_{I_{\text{in}}}\>=\>&\mathbb{T}\mathbb{R}\times\mathbb{R}\times\mathbb{T}\mathbb{R}\times\mathbb{T}\mathbb{R}\times\mathbb{R}\\ \>\ni\>&\bigl{(}(\mathtt{stator.q.x},\,\mathtt{stator.q.f},\,\mathtt{stator.q.e}),\,\mathtt{stator.b_{s}.x},\,\\ &\>\>(\mathtt{rotor.q.x},\,\mathtt{rotor.q.f},\,\mathtt{rotor.q.e}),\,\\ &\>\>(\mathtt{rotor.p.x},\,\mathtt{rotor.p.f},\,\mathtt{rotor.p.e}),\,\mathtt{rotor.b_{s}.x}\bigr{)}\end{split}$
+$\begin{split}\mathcal{P}_{I_{\text{in}}}\>=\>&\mathbb{R}\times\mathbb{T}\mathbb{R}\times\mathbb{T}\mathbb{R}\times\mathbb{R}\times\mathbb{T}\mathbb{R}\\ \>\ni\>&\bigl{(}\mathtt{rotor.b_{s}.x},\,(\mathtt{rotor.p.x},\,\mathtt{rotor.p.f},\,\mathtt{rotor.p.e}),\,\\ &\>\>(\mathtt{rotor.q.x},\,\mathtt{rotor.q.f},\,\mathtt{rotor.q.e}),\,\\ &\>\>\mathtt{stator.b_{s}.x},\,(\mathtt{stator.q.x},\,\mathtt{stator.q.f},\,\mathtt{stator.q.e})\bigr{)}\end{split}$
 and the port space of the outer interface $I$ is
-$\mathcal{P}_{I}\>=\>\mathbb{T}\mathbb{R}\times\mathbb{T}\mathbb{R}\>\ni\>\bigl{(}(\mathtt{q.x},\,\mathtt{q.f},\,\mathtt{q.e}),\,(\mathtt{p.x},\,\mathtt{p.f},\,\mathtt{p.e})\bigr{)}\,.$
+$\mathcal{P}_{I}\>=\>\mathbb{T}\mathbb{R}\times\mathbb{T}\mathbb{R}\>\ni\>\bigl{(}(\mathtt{p.x},\,\mathtt{p.f},\,\mathtt{p.e}),\,(\mathtt{q.x},\,\mathtt{q.f},\,\mathtt{q.e})\bigr{)}\,.$
 Finally, the subspace $\mathcal{P}_{P}\subset\mathcal{P}_{I_{\text{in}}}\times\mathcal{P}_{I}$ defining the relation is
-$\begin{split}\mathcal{P}_{P}\>=\>\bigl{\{}\bigl{(}&(\mathtt{stator.q.x},\,\mathtt{stator.q.f},\,\mathtt{stator.q.e})\,,\,\mathtt{stator.b_{s}.x}\,,\,\\ &(\mathtt{rotor.q.x},\,\mathtt{rotor.q.f},\,\mathtt{rotor.q.e})\,,\,\\ &(\mathtt{rotor.p.x},\,\mathtt{rotor.p.f},\,\mathtt{rotor.p.e})\,,\,\mathtt{rotor.b_{s}.x}\,,\,\\ &(\mathtt{q.x},\,\mathtt{q.f},\,\mathtt{q.e})\,,\,(\mathtt{p.x},\,\mathtt{p.f},\,\mathtt{p.e})\bigr{)}\in\mathcal{P}_{I_{i}}\times\mathcal{P}_{I}\,\mid\,\\ &\mathtt{stator.q.x}=\mathtt{rotor.q.x}=\mathtt{q.x}\,,\,\\ &\mathtt{stator.q.e}=\mathtt{rotor.q.e}=\mathtt{q.e}\,,\,\\ &\mathtt{stator.q.f}+\mathtt{rotor.q.f}=\mathtt{q.f}\,,\,\\ &\mathtt{stator.b_{s}.x}=\mathtt{rotor.b_{s}.x}\,,\,\\ &\mathtt{rotor.p.x}=\mathtt{p.x}\,,\,\\ &\mathtt{rotor.p.e}=\mathtt{p.e}\,,\,\\ &\mathtt{rotor.p.f}=\mathtt{p.f}\,\bigr{\}}\,.\qed\end{split}$
+$\begin{split}\mathcal{P}_{P}\>=\>\bigl{\{}&\bigl{(}\mathtt{rotor.b_{s}.x}\,,\,(\mathtt{rotor.p.x},\,\mathtt{rotor.p.f},\,\mathtt{rotor.p.e})\,,\,\\ &(\mathtt{rotor.q.x},\,\mathtt{rotor.q.f},\,\mathtt{rotor.q.e})\,,\,\\ &\mathtt{stator.b_{s}.x}\,,\,(\mathtt{stator.q.x},\,\mathtt{stator.q.f},\,\mathtt{stator.q.e})\,,\,\\ &(\mathtt{p.x},\,\mathtt{p.f},\,\mathtt{p.e})\,,\,(\mathtt{q.x},\,\mathtt{q.f},\,\mathtt{q.e})\bigr{)}\in\mathcal{P}_{I_{i}}\times\mathcal{P}_{I}\,\mid\,\\ &\mathtt{rotor.b_{s}.x}=\mathtt{stator.b_{s}.x}\,,\,\\ &\mathtt{rotor.p.x}=\mathtt{p.x}\,,\,\\ &\mathtt{rotor.p.f}=\mathtt{p.f}\,,\,\\ &\mathtt{rotor.p.e}=\mathtt{p.e}\,,\,\\ &\mathtt{rotor.q.x}=\mathtt{stator.q.x}=\mathtt{q.x}\,,\,\\ &\mathtt{rotor.q.f}+\mathtt{stator.q.f}=\mathtt{q.f}\,,\,\\ &\mathtt{rotor.q.e}=\mathtt{stator.q.e}=\mathtt{q.e}\,\bigr{\}}\,.\qed\end{split}$
+Given a hierarchy of interconnection patterns, we can first compose the patterns and then determine the relation assigned to the composed pattern.
+Alternatively, we can first determine the relation assigned to each individual pattern and then compose these relations.
+Both ways lead to the same result.
+The combinatorial syntax and its relational semantics hence compose in compatible ways.
+In the language of category theory, we say that the (assignment of) semantics is functorial.
+A future paper shall use directories and the theory of multicategories/operads to demonstrate that 4.2 indeed defines a suitable notion of functor.
 5 Exergy
-Every system is ultimately determined by a power-preserving interconnection of primitive subsystems, also called components, which represent elementary physical behaviors such as storage and reversible as well as irreversible exchange of energy.
-To achieve a thermodynamically consistent combination of reversible and irreversible dynamics, components are defined with respect to a fixed exergy reference environment that is discussed in this section.
+Every system is ultimately determined by a power-preserving interconnection of primitive subsystems, called components.
+These represent elementary physical behaviors, namely storage, and reversible as well as irreversible exchange of energy.
+To achieve a thermodynamically consistent combination of reversible and irreversible dynamics, components are defined with respect to an exergy reference environment.
+Before components can be defined in the next section, in this section we define the reference environment.
+We start with its physical interpretation.
 5.1 Intuition
-As briefly discussed in Section 1.4, thermodynamic models do not have all their physically-relevant degrees of freedom determined by their state variables, which leads to irreversibility.
-According to the first and second law of thermodynamics, internal energy, which is given by an imprecise phenomenological description, cannot be fully converted into mechanical or electromagnetic energy, which is precise, in the sense that the relevant degrees of freedom are determined by the state variables.
+As briefly discussed in Section 1.4, thermodynamic models have an irreversible dynamics, since not all physically-relevant degrees of freedom are determined by the state variables.
+According to the first and second law of thermodynamics, internal energy, which is given by an ‘imprecise’ phenomenological description, cannot be fully converted into mechanical or electromagnetic energy, which is ‘precise’, in the sense that all relevant degrees of freedom are fully determined by the state variables.
 The irreversible degradation of ‘mechanically precise’ forms of energy into ‘thermodynamically imprecise’ forms of energy can be quantified using the concept of exergy.
-A hypothetical conversion device, whose mode of operation is only restricted by the first and second law of thermodynamics is used to determine how much mechanical (or electromagnetic) energy can be extracted from a system before it reaches thermodynamic equilibrium with a reference environment, whose intensive variables, such as temperature and pressure, are constant (in time and space), see Figure 9.
+A hypothetical conversion device, whose mode of operation is only restricted by the first and second law of thermodynamics is used to determine how much mechanical (or electromagnetic) energy can be extracted from a system, before it reaches thermodynamic equilibrium with a reference environment, see Figure 9.
+The environment is defined by intensive variables, such as temperature and pressure, which are constant in time and space.
 Figure 9: To assess the exergy content of a system, it is combined with a reference environment.
 The exergy content is then defined as the maximum amount of work that can be extracted from the combined system via a hypothetical conversion device that operates completely reversibly.
-Crucially, the ideal conversion device is not restricted to interacting with the system through its interface.
+The ideal conversion device is not restricted to interacting with the system through its interface.
 Instead, the device operates in a counterfactual where all energy domains of both the system and the environment can be accessed.
-This means that the exergy content of a (hierarchically flattened) composite system can be obtained by summing the exergy stored in each of its storage components.
-So, in particular, the exergy content of a system does not depend on how its subsystems are interconnected.
-The exergy content of a component storing internal energy is assessed based on an ideal conversion device and the reference environment, while the exergy content of a component storing only mechanical or electromagnetic forms of energy is equal to its energy content.
-As also remark that a similar concept of available energy could be defined intrinsically, meaning without relying on a reference environment.
-The reversible conversion device would then interact only with the system of interest until it reaches equilibrium with itself.
-However, the use of a reference environment with fixed intensive quantities has the important advantage that the exergy content of a system is given explicitly by its exergy storage function, which follows systematically from its energy storage function and the definition of the reference environment, as detailed in Section 6.1.
-In contrast, the intrinsic concept in general would lead to an optimization problem without a simple closed-form solution.
-More information and references about exergy, its history, and its links to port-Hamiltonian as well as GENERIC systems can be found in Lohmayer et al. (2021).
+The exergy content of a composite system is hence obtained by (recursively) summing the exergy stored in each of its storage components.
+The exergy content of a component storing purely mechanical or electromagnetic forms of energy is simply equal to its energy content.
+For components storing internal energy, the exergy content is assessed based on the ideal conversion device and the reference environment.
+A concrete statement of this follows in 6.3, after the definition of the exergy storage function in 6.2.
+For more information about exergy, its history, and its links to port-Hamiltonian as well as metriplectic or GENERIC systems, we refer to [3].
 5.2 Reference environment
+The internal energy of a thermodynamic system is a function of its entropy, and possibly further extensive quantities such as volume, mass of chemical species, etc.
+To assess the exergy content of the system, we need for each such variable an energy-conjugate, intensive variable attributed to the reference environment.
+Since these intensive variables are constant, they determine the equilibrium state of the combined system, see Figure 9.
 Definition 5.1.
-A reference environment $(I_{R},\,\lambda_{R})$ is given by an interface $I_{R}=(N_{R},\,\tau_{R})$ such that each port $p\in I_{R}$ is a power port with an unique extensive quantity in $\mathbf{Q}_{\mathbb{R}}$ and an associated value $\lambda_{R}(p)\in\mathbb{R}^{*}$ of the energy-conjugate intensive quantity.
+A reference environment $(I_{R},\,\lambda_{R})$ is defined by an interface $I_{R}=(N_{R},\,\tau_{R})$ whose ports are all power ports and have a unique extensive quantity in $\mathbf{Q}_{\mathbb{R}}$.
+One port must have the quantity $(\mathbb{R},\,\mathtt{entropy})$.
+Moreover, each port $p\in I_{R}$ has an associated value $\lambda_{R}(p)\in\mathbb{R}^{*}$ of the energy-conjugate intensive quantity.
 ∎
@@
 The minimal definition of a reference environment represents an isothermal heat bath.
-This amounts to an interface with a single entropy port and an associated reference temperature, henceforth denoted by ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}$:
-$N_{R}\>$ =$=\>\{\mathtt{s}\}$
-$\tau_{R}(\mathtt{s})\>$ $=\>\left((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p}\right)$
-[…]
-To determine the exergy content of pressurized fluids with respect to an isobaric atmosphere, the definition of the environment can be extended to $N_{R}=\{\mathtt{s},\,\mathtt{v}\}$ with
-$\tau_{R}(\mathtt{v})\>$ $=\>\left((\mathbb{R},\,\mathtt{volume}),\,\mathsf{p}\right)$
-[…]
-The minus sign in front of the reference pressure ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\pi_{0}}$ stems from the fact that more volume means less energy.
-To model for instance a mass diffusion process, a port $\mathtt{m}$ representing the mass of a chemical species with a fixed chemical potential ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\mu_{0}}$ should be included, i.e.
+This amounts to an interface with a single entropy port and an associated reference temperature, henceforth denoted by ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}$:
+$N_{R}$ =$=\>\{\mathtt{s}\}$
+$\tau_{R}(\mathtt{s})$ $=\>\left((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p}\right)$
+$\lambda_{R}(\mathtt{s})$ $=\>{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\,.$
+To determine the exergy content of pressurized fluids with respect to an isobaric atmosphere, the definition of the environment is extended to $N_{R}=\{\mathtt{s},\,\mathtt{v}\}$ with
+$\tau_{R}(\mathtt{v})$ $=\>\left((\mathbb{R},\,\mathtt{volume}),\,\mathsf{p}\right)$
+$\lambda_{R}(\mathtt{v})$ $=\>{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}-\pi_{0}}\,.$
+The minus sign in front of the reference pressure ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pi_{0}}$ stems from the fact that more volume means less energy.
+To model for instance a mass diffusion process, a port $\mathtt{m}$ representing the mass of a chemical species with a constant chemical potential ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\mu_{0}}$ can be included, i.e.
 let $N_{R}=\{\mathtt{s},\,\mathtt{v},\,\mathtt{m}\}$ with
 […]
-A software implementation should include a comprehensive definition of a reference environment and allow the user to redefine the values of the intensive quantities.
+A software implementation should already include a comprehensive definition of a reference environment.
+Allowing the user to change the values of the intensive quantities would be of interest for exergy analysis and thermodynamic optimization.
 6 Primitive systems
-The lowest level of any hierarchical model specification is defined by primitive systems, also called components. There are four classes of components:
+At the lowest level, any hierarchically-structured model is defined by primitive systems, called components. There are four kinds of them:
   • 1.
@@
     Reversible components represent reversible energy exchange
-    (or a transformation between different forms of describing an energy domain).
+    or a transformation between different forms of describing an energy domain.
   • 3.
@@
   • 4.
-    Environment components are used to model energy exchange with
+    Environment components may be used to model energy exchange with
     the reference environment.
@@
 Further, every component has an interface $I=(N,\,\tau)$ with port variables $(x,\,f,\,e)\in\mathcal{P}_{I}$.
-To define the exergy storage function of storage components and to state fundamental thermodynamic properties of reversible as well as irreversible components, we define the following notation: For any port $p\in I$, let
+To define the exergy storage function of storage components and to state conditions that reversible and irreversible components have to satisfy, we define the following notation: For any port $p\in I$, let
 $\lambda(p)\>=\>\begin{cases}\lambda_{R}(p_{R})&\exists\,p_{R}\in I_{R}\,\colon\,\tau_{R}(p_{R})\>=\>\tau(p)\\ 0\in\mathrm{T}^{*}\mathcal{X}_{p}&\text{else.}\end{cases}$
-For instance, if $p$ represents a thermal energy domain then $\lambda(p)={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}$, whereas if $p$ represents a kinetic energy domain then $\lambda(p)=0$, since the environment does not have a kinetic energy domain.
-This can also be applied to the entire interface.
-E.g.
-for $N=\{\mathtt{p_{1}},\,\ldots,\,\mathtt{p_{n}}\}$, let
+With regard to the existential quantifier, this is well-defined due to the injectivity of $\tau_{R}$.
+For instance, if $p$ represents a thermal energy domain then $\lambda(p)={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}$, whereas if $p$ represents a kinetic energy domain then $\lambda(p)=0$, since the environment does not have a kinetic energy domain.
+We extend the above notation to an entire interface: Given for example $N=\{\mathtt{p_{1}},\,\ldots,\,\mathtt{p_{n}}\}$, let
 $\lambda(I)\>=\>\bigl{(}\lambda(\mathtt{p_{1}}),\,\ldots,\,\lambda(\mathtt{p_{n}})\bigr{)}\,.$
-We also define a similar notation that references a specific environment port: For any port $p\in I$ and $p_{R}\in I_{R}$, let
+We also need a similar notation that references a specific environment port: For any port $p\in I$ and $p_{R}\in I_{R}$, let
 $\lambda(p,\,p_{R})\>=\>\begin{cases}\lambda_{R}(p_{R})&\tau_{R}(p_{R})\>=\>\tau(p)\\ 0\in\mathrm{T}^{*}\mathcal{X}_{p}&\text{else.}\end{cases}$
-Again, given e.g. $N=\{\mathtt{p_{1}},\,\ldots,\,\mathtt{p_{n}}\}$, let
+Again, given for example $N=\{\mathtt{p_{1}},\,\ldots,\,\mathtt{p_{n}}\}$, let
 $\lambda(I,\,p_{R})\>=\>\bigl{(}\lambda(\mathtt{p_{1}},\,p_{R}),\,\ldots,\,\lambda(\mathtt{p_{n}},\,p_{R})\bigr{)}\,.$
 6.1 Storage components
-Storage components are primitive systems representing energy storage. Examples include elastic springs, movable masses, capacitors, inductors, and thermal capacities.
+Storage components are primitive systems that represent energy storage. Examples include elastic springs, movable masses, capacitors, inductors, and thermal capacities.
 Definition 6.1.
 A storage component $(I,\,E)$ is defined by an interface $I$ containing only power ports and an energy storage function $E:\mathcal{X}_{I}\rightarrow\mathbb{R}$. ∎
-The function $E$ simply yields the stored energy for any state $x\in\mathcal{X}_{I}$. Based on the reference environment $I_{R}$, the exergy content is determined as follows.
+The function $E$ yields the stored energy for any state $x\in\mathcal{X}_{I}$. Based on the reference environment $I_{R}$, the exergy content is determined as follows.
 Definition 6.2.
-Let $(I,\,E)$ be a storage component. Its exergy storage function $H:\mathcal{X}_{I}\rightarrow\mathbb{R}$ is given by
+Let $(I,\,E)$ be a storage component. Its exergy storage function $H:\mathcal{X}_{I}\rightarrow\mathbb{R}$ is defined by
 $H(x)\>=\>E(x)-\sum_{p\,\in\,I}\sum_{\begin{subarray}{c}p_{R}\,\in\,I_{R}\\ \tau_{R}(p_{R})\,=\,\tau(p)\end{subarray}}\langle\lambda_{R}(p_{R})\mid p\mathtt{.x}\rangle\,.\qed$
+The exergy function $H$ follows directly from the energy function $E$ by subtracting the pairing between each state variable representing an extensive quantity also present in the environment, and the constant, energy-conjugate quantity from the environment.
+We come back to this in 6.3.
 […]
-$\forall\,p\in I\>:\>p\mathtt{.e}\>=\>\left.\partialderivative{H}{(p\mathtt{.x})}\right|_{x}$
-In case the constant offset should matter for thermodynamic analysis, it can be added to the total exergy function once the equilibrium state of the system under study has been determined.
-Finally, the trajectory of the state variables is obtained by integrating the flow variables:
+The effort variables are given by the differential of the exergy function evaluated at the current state $x\in\mathcal{X}_{I}$, i.e.
+$\forall\,p\in I\>:\>p\mathtt{.e}\>=\>\left.\partialderivative{H}{(p\mathtt{.x})}\right|_{x}\,.$
+The evolution of the state variables is obtained by integrating the flow variables, i.e.
+[…]
 […]
@@
 Consider a storage component $(I_{\text{gas}},\,E_{\text{gas}})$ representing a gas-filled compartment whose entropy and volume can change.
-Its interface $I_{\text{gas}}$ is defined as in 2.18 and we assume knowledge of its energy storage function $E_{\text{gas}}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$.
-Writing $x=(s,\,v)$, the exergy storage function is then given by
-$H_{\text{gas}}(s,\,v)\>=\>E_{\text{gas}}(s,\,v)\,-\,{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}\,s\,+\,{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\pi_{0}}\,v\,.$
-With the temperature being defined as $\theta=\partialderivative{E_{\text{gas}}}{s}$ and the negative pressure being defined as $-\pi=\partialderivative{E_{\text{gas}}}{v}$, the effort variables are given by
-$\begin{split}\mathtt{s.e}\>&=\>\theta-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}\\ \mathtt{v.e}\>&=\>-(\pi-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\pi_{0}})\,.\end{split}$
-Since the entropy rate $\dot{s}=\mathtt{s.f}$ is related to the heat flow rate $\Phi$ via $\dot{s}=\frac{1}{\theta}\,\Phi$, the exergetic power at the entropy port is $\langle\mathtt{s.e}\mid\mathtt{s.f}\rangle=\frac{\theta-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}}{\theta}\,\Phi=\eta_{\text{Carnot}}\,\Phi$, where $\eta_{\text{Carnot}}$ is the efficiency of an ideal heat engine operating between a hot source with temperature $\theta$ and a cold sink with temperature ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}$.
-Considering a piston with the pressurized gas on one side and the atmosphere on the other, $\langle\mathtt{v.e}\mid\mathtt{v.f}\rangle=-(\pi-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\pi_{0}})\,\dot{v}$ is the power that can be extracted from the pressurized gas when its volume grows at rate $\dot{v}$.
+Its interface $I_{\text{gas}}$ is defined in 2.18, and we assume knowledge of its energy storage function $E_{\text{gas}}:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$.
+With $x=(s,\,v)$, the exergy storage function is then given by
+$H_{\text{gas}}(s,\,v)\>=\>E_{\text{gas}}(s,\,v)\,-\,{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\,s\,+\,{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pi_{0}}\,v\,.$
+As the temperature is defined by $\theta=\partialderivative{E_{\text{gas}}}{s}$ and the (negative) pressure is defined by $-\pi=\partialderivative{E_{\text{gas}}}{v}$, the effort variables are given by
+$\begin{split}\mathtt{s.e}\>&=\>\theta-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\\ \mathtt{v.e}\>&=\>-(\pi-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pi_{0}})\,.\end{split}$
+Since the entropy rate $\dot{s}=\mathtt{s.f}$ is related to the heat flow rate $\Phi$ via $\dot{s}=\frac{1}{\theta}\,\Phi$, the exergetic power at the entropy port is $\langle\mathtt{s.e}\mid\mathtt{s.f}\rangle=\frac{\theta-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}}{\theta}\,\Phi=\eta_{\text{Carnot}}\,\Phi$, where $\eta_{\text{Carnot}}$ is the efficiency of an ideal heat engine operating between a hot source with temperature $\theta$ and a cold sink with temperature ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}$.
+Considering a piston with the pressurized gas on one side and the atmosphere on the other, $\langle\mathtt{v.e}\mid\mathtt{v.f}\rangle=-(\pi-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pi_{0}})\,\dot{v}$ is the power that can be extracted from the pressurized gas when the volume expansion rate is $\dot{v}$.
 ∎
-Next, we define the semantics of a storage component with interface $I$ in terms of a relation $\mathcal{S}:\mathrm{T}\mathcal{X}_{I}\nrightarrow\mathcal{P}_{I}$.
+Next, we formally state the semantics of a storage component with interface $I$ in terms of a relation $\mathcal{S}:\mathrm{T}\mathcal{X}_{I}\nrightarrow\mathcal{P}_{I}$.
 This is used in Section 7, where we define the semantics of a composite system in terms of the semantics of its primitive subsystems and the semantics of its (hierarchy of nested) interconnection pattern(s).
@@
 6.2 Environment components
-While the primary role of the reference environment is to enable the thermodynamically consistent combination of reversible and irreversible dynamics, it may also be represented by a component to model energy exchange with the environment.
-For instance, an environment component can represent an isothermal heat bath, which may absorb heat generated by irreversible components, or an isobaric atmosphere, which may exchange pressure-volume work with other components.
+From a user perspective, environment components provide a simple alternative to storage components, when dealing with low-grade energy.They may represent an isothermal heat bath, which absorbs heat generated by irreversible components, or an isobaric atmosphere, which exchanges pressure-volume work with other components, etc.
+Environment components are essentially storage components with a trivial energy and exergy function.
 Definition 6.5.
 An environment component is defined by a subinterface $I\subseteq I_{R}$ containing the exposed energy domains of the reference environment. ∎
-The state variables keep track of the exchanged entropy, volume, etc. As for storage components, their rate of change is given by the flow variables:
+The state variables keep track of the exchanged entropy, volume, etc. As for storage components, their rate of change is given by the flow variables, i.e.
 […]
-Since the exergy content of the environment is zero by definition, the effort variables are also zero:
+By definition, the environment has zero exergy content. Since the effort variables are given by the differential of the exergy function, they are also zero:
 […]
-Consequently, the semantics of an environment component is assigned as follows:
+In agreement with the definition, the chain rule also tells us that the rate of change of the stored exergy is zero.
 Definition 6.6.
-The semantics assigned to environment component $I\subseteq I_{R}$ is the relation $\mathcal{E}:\mathrm{T}\mathcal{X}_{I}\nrightarrow\mathcal{P}_{I}$ that is given by the subbundle
+The semantics assigned to environment component $I\subseteq I_{R}$ is the relation
+$\mathcal{E}:\mathrm{T}\mathcal{X}_{I}\nrightarrow\mathcal{P}_{I}$ that is given by the subbundle
 […]
@@
 In a generalized sense, reversible components represent gyrators, transformers, and/or kinematic constraints.
-Generalized gyrators describe the reversible coupling of different kinds of energy domains, such as a coupling of kinetic and potential energy domains in mechanics or an electro-magnetic coupling.
+Generalized gyrators may describe a coupling of kinetic and potential energy domains in mechanics, or a coupling of magnetic and electric energy domains in electromagnetism.
 Examples where one domain reversibly interacts with itself are gyroscopic forces in rigid body dynamics and self-advection of kinetic energy in fluid dynamics.
-The Lorentz force is another example where a kinetic energy domain interacts with itself in a way that depends on the state of a magnetic energy domain.
-Generalized transformers couple energy domains of the same kind.
-They may for instance represent mechanical levers or ideal electrical transformers.
-Frame transformations in multibody systems present an example where a transformation between different representations of the same energy domain is realized.
+The Lorentz force is an example where a kinetic energy domain interacts with itself in a way that depends on the state of a magnetic energy domain.
+Generalized transformers may for instance describe ideal mechanical levers or electrical transformers.
+Frame transformations in multibody systems are an example relating different representations of the same energy domain.
 Examples of kinematic constraints include joint constraints in multibody dynamics and the incompressibility constraint in fluid dynamics.
-Further, the parallel connection of capacitors, the series connection of inductors or springs, as well as rigid links between point masses are described by constraints.
+Further, the parallel connection of capacitors, the series connection of inductors or springs, and rigid links between point masses are described by constraints.
 The geometric structure behind reversible components is called Dirac structure.
-For our purposes, an (almost) Dirac structure $\mathcal{D}$ is a power-preserving, locally linear relation of port variables.
-More precisely, we could define $\mathcal{D}\subset\mathcal{P}_{I}$ as a maximally isotropic vector subbundle of the bundle of port variables $\mathcal{P}_{I}\twoheadrightarrow\mathcal{X}_{I}$, see Courant (1990); Bursztyn (2013); van der Schaft and Jeltsema (2014).
-However, to state fundamental properties of reversible components, we give a more practical definition of Dirac structures in terms of a particular representation, which may be called constrained hybrid input-output representation (cf.
-van der Schaft and Jeltsema (2014)):
+For our purposes, a Dirac structure $\mathcal{D}$ is a power-preserving, locally linear relation of port variables.
+More precisely, we could define $\mathcal{D}\subset\mathcal{P}_{I}$ as a maximally isotropic vector subbundle of the bundle of port variables $\mathcal{P}_{I}\twoheadrightarrow\mathcal{X}_{I}$, see [1, 4, 26].
+However, to state additional conditions that reversible components have to satisfy, we give a more practical definition of Dirac structures in terms of a particular representation, which might be called constrained hybrid input-output representation (cf.
+[4]).
 Definition 6.7.
-Let $I$ be an interface and let $\mathcal{X}_{I,p}\cong\mathcal{X}_{1}\times\mathcal{X}_{2}\ni x_{p}=(x_{1},\,x_{2})$ be a factorization of the state space of its power ports with flow and effort variables correspondingly denoted by $f=(f_{1},\,f_{2})$ and $e=(e_{1},\,e_{2})$.
+Let $I$ be an interface and let $\mathcal{X}_{I,p}\cong\mathcal{X}_{1}\times\mathcal{X}_{2}\ni x_{p}=(x_{1},\,x_{2})$ be a factorization of the state space of all power ports with flow and effort variables correspondingly denoted by $f=(f_{1},\,f_{2})$ and $e=(e_{1},\,e_{2})$.
 A Dirac structure $\mathcal{D}\subset\mathcal{P}_{I}$ admits the following representation:
 $\mathcal{D}\,=\,\left\{(x,\,f,\,e)\in\mathcal{P}_{I}\,\mid\,\exists\,\lambda_{c}\in V^{*}\,\colon\left[\begin{array}[]{c}f_{1}\\ \hline\cr e_{2}\\ \hline\cr 0\end{array}\right]\,=\,\left[\begin{array}[]{c | c | c}L(x)&-g(x)&C^{*}(x)\\ \hline\cr g^{*}(x)&0&0\\ \hline\cr-C(x)&0&0\end{array}\right]\,\left[\begin{array}[]{c}e_{1}\\ \hline\cr f_{2}\\ \hline\cr\lambda_{c}\end{array}\right]\right\}$
-Here, $L:\mathrm{T}^{*}\mathcal{X}_{1}\rightarrow\mathrm{T}\mathcal{X}_{1}$, $g:\mathrm{T}\mathcal{X}_{2}\rightarrow\mathrm{T}\mathcal{X}_{1}$, its linear dual $g^{*}:\mathrm{T}^{*}\mathcal{X}_{1}\rightarrow\mathrm{T}^{*}\mathcal{X}_{2}$, $C:\mathrm{T}^{*}\mathcal{X}_{1}\rightarrow\mathcal{X}_{I}\times V$, and its dual $C^{*}:\mathcal{X}_{I}\times V^{*}\rightarrow\mathrm{T}\mathcal{X}_{1}$ are vector bundle maps over the identity on $\mathcal{X}_{I}$.
+Here, $L:\mathrm{T}^{*}\mathcal{X}_{1}\rightarrow\mathrm{T}\mathcal{X}_{1}$, $g:\mathrm{T}\mathcal{X}_{2}\rightarrow\mathrm{T}\mathcal{X}_{1}$, its linear dual $g^{*}:\mathrm{T}^{*}\mathcal{X}_{1}\rightarrow\mathrm{T}^{*}\mathcal{X}_{2}$,
+$C:\mathrm{T}^{*}\mathcal{X}_{1}\rightarrow\mathcal{X}_{I}\times V$, and its dual $C^{*}:\mathcal{X}_{I}\times V^{*}\rightarrow\mathrm{T}\mathcal{X}_{1}$ are vector bundle maps over the identity on $\mathcal{X}_{I}$.
 Further, $V$ is a vector space and $L(x)$ is skew-symmetric for all $x\in\mathcal{X}_{I}$.
 ∎
-The maps $L$, $g$ and $C$ define a gyrator, a transformer and a constraint, respectively.
+The maps $L$, $g$ and $C$ respectively define a generalized gyrator, a generalized transformer and a constraint.
 The constraint is enforced via the Lagrange multiplier $\lambda_{c}$.
@@
 […]
+We remark that this actually defines an almost Dirac structure, as no integrability condition, akin to the Jacobi identity for Poisson structures, is required.
+However, in the unconstrained or holonomically constrained case, such a condition can be checked.
 Example 6.8.
-Let the interface $I_{\text{hk}}=(N_{\text{hk}},\,\tau_{\text{hk}})$ be given by
-$\begin{split}N_{\text{hk}}\>&=\>\{\mathtt{v_{1}},\,\mathtt{v_{2}},\,\mathtt{p}\}\\ \tau_{\text{hk}}(\mathtt{v_{1}})\>&=\>\tau_{\text{hk}}(\mathtt{v_{2}})\>=\>((\mathbb{R},\,\mathtt{volume}),\,\mathsf{p})\\ \tau_{\text{hk}}(\mathtt{p})\>&=\>((\mathbb{R},\,\mathtt{momentum}),\,\mathsf{p})\end{split}$
-and let $a\in\mathbb{R}$ be a constant surface area of a piston. Then, the following equation defines a Dirac structure $\mathcal{D}_{\text{hk}}\subset\mathcal{P}_{I_{\text{hk}}}$:
-$\left[\begin{array}[]{c}\mathtt{v_{1}.f}\\ \mathtt{v_{2}.f}\\ \mathtt{p.f}\end{array}\right]\>=\>\left[\begin{array}[]{rrr}0&0&a\\ 0&0&-a\\ -a&a&0\end{array}\right]\,\left[\begin{array}[]{c}\mathtt{v_{1}.e}\\ \mathtt{v_{2}.e}\\ \mathtt{p.e}\end{array}\right]$
-This describes a generalized gyrator that couples the two hydraulic energy domains at the front and the back side of a piston with a kinetic energy domain corresponding to the piston’s moving mass. ∎
-A reversible component is defined by a Dirac structure which satisfies two extra properties ensuring thermodynamic consistency.
+Let the interface $I_{\text{hkc}}=(N_{\text{hkc}},\,\tau_{\text{hkc}})$ be given by
+$\begin{split}N_{\text{hkc}}\>&=\>\{\mathtt{v_{1}},\,\mathtt{v_{2}},\,\mathtt{p}\}\\ \tau_{\text{hkc}}(\mathtt{v_{1}})\>&=\>\tau_{\text{hkc}}(\mathtt{v_{2}})\>=\>((\mathbb{R},\,\mathtt{volume}),\,\mathsf{p})\\ \tau_{\text{hkc}}(\mathtt{p})\>&=\>((\mathbb{R},\,\mathtt{momentum}),\,\mathsf{p})\end{split}$
+and let $a\in\mathbb{R}$ be a constant surface area of a piston. Then, the equation
+$\left[\begin{array}[]{c}\mathtt{v_{1}.f}\\ \mathtt{v_{2}.f}\\ \mathtt{p.f}\end{array}\right]\>=\>\left[\begin{array}[]{rrr}0&0&-a\\ 0&0&+a\\ +a&-a&0\end{array}\right]\,\left[\begin{array}[]{c}\mathtt{v_{1}.e}\\ \mathtt{v_{2}.e}\\ \mathtt{p.e}\end{array}\right]$
+defines a Dirac structure $\mathcal{D}_{\text{hkc}}\subset\mathcal{P}_{I_{\text{hkc}}}$.
+This describes a generalized gyrator that couples the two hydraulic energy domains at the front and the back side of a piston with the kinetic energy domain of the piston’s moving mass.
+∎
+A reversible component is defined by a Dirac structure that satisfies two additional conditions ensuring thermodynamic consistency.
 Definition 6.9.
-A reversible component $(I,\,\mathcal{D})$ is given by an interface $I$ and a Dirac structure $\mathcal{D}\subset\mathcal{P}_{I}$ such that all quantities of the reference environment are conserved and the time-reversal invariance property holds.
-The first condition requires that
-$\forall\,p_{R}\in I_{R}\>:\>\forall\,x\in\mathcal{X}_{I}\>:\>\Bigl{(}0\>=\>L(x)\,\lambda(I_{1},\,p_{R})\quad\text{and}\quad 0\>=\>g^{*}(x)\,\lambda(I_{1},\,p_{R})\Bigr{)}\,.$
+A reversible component $(I,\,\mathcal{D})$ is given by an interface $I$ and a Dirac structure $\mathcal{D}\subset\mathcal{P}_{I}$ such that all quantities present in the reference environment (in particular entropy) are conserved and the time-reversal invariance property holds.
+The first condition requires that $\forall\,p_{R}\in I_{R}$ and $\forall\,x\in\mathcal{X}_{I}\,\colon$
+$0\>=\>L(x)\,\lambda(I_{1},\,p_{R})\quad\text{and}\quad 0\>=\>g^{*}(x)\,\lambda(I_{1},\,p_{R})\quad\text{and}\quad 0\>=\>C(x)\,\lambda(I_{1},\,p_{R})\,.$
 The time-reversal invariance property holds if $\forall\,x\in\mathcal{X}_{I}\,\colon$
@@
 ∎
+According to the chain rule, a Dirac structure $\mathcal{D}$ conserves the quantity corresponding to some port $p_{R}\in I_{R}$ if any resulting flow variable $f_{1}$ satisfies $0=\langle\lambda(I_{1},\,p_{R})\mid f_{1}\rangle$.
+Inserting the expression for $f_{1}$, this condition is satisfied if $0=\langle\lambda(I_{1},\,p_{R})\mid L(x)\,e_{1}\rangle$ is satisfied for any effort variable $e_{1}$, and $0=\langle\lambda(I_{1},\,p_{R})\mid g(x)\,f_{2}\rangle$ is satisfied for any flow variable $f_{2}$, and $0=\langle\lambda(I_{1},\,p_{R})\mid C^{*}(x)\,\lambda_{c}\rangle$ is satisfied for any Lagrange multiplier $\lambda_{c}$.
+Taking the linear adjoint (transpose), the first condition requires that $0=\langle L^{*}(x)\,\lambda(I_{1},\,p_{R})\mid e_{1}\rangle$ for any $e_{1}$.
+This simplifies to the first part of the condition in the definition, since $L$ is skew-adjoint.
+The other two parts follow analogously.
+Regarding details on the time-reversal invariance condition, we refer to [46] or [3].
 Example 6.10.
-The reversible component $(I_{\text{hk}},\,\mathcal{D}_{\text{hk}})$, as already defined in 6.8, describes the coupling between the kinetic energy domain of a piston with surface area $a$ and the two hydraulic energy domains on either side of the piston.
+The reversible component $(I_{\text{hkc}},\,\mathcal{D}_{\text{hkc}})$, as already defined in 6.8, describes the coupling between the kinetic energy domain of a piston with surface area $a$ and the two hydraulic energy domains on either side of the piston.
 The first condition here ensures that volume is conserved.
-This is satisfied since
-$\left[\begin{array}[]{c}0\\ 0\\ 0\end{array}\right]\>=\>\left[\begin{array}[]{rrr}0&0&a\\ 0&0&-a\\ -a&a&0\end{array}\right]\,\left[\begin{array}[]{c}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}-\pi_{0}}\\ {\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}-\pi_{0}}\\ 0\end{array}\right]\,.$
+It is satisfied since
+$\left[\begin{array}[]{c}0\\ 0\\ 0\end{array}\right]\>=\>\left[\begin{array}[]{rrr}0&0&-a\\ 0&0&+a\\ +a&-a&0\end{array}\right]\,\left[\begin{array}[]{c}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}-\pi_{0}}\\ {\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}-\pi_{0}}\\ 0\end{array}\right]\,.$
 The time-reversal invariance property holds since
-$\begin{split}-\mathrm{P}(\mathtt{(\mathbb{R},\,volume}))\>&=\>\mathrm{P}(a)\,\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\\ -\mathrm{P}((\mathbb{R},\,\mathtt{volume}))\>&=\>\mathrm{P}(-a)\,\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\,.\end{split}$
-Checks for the other two non-zero entries of $L$ must pass due to its skew-symmetry. ∎
-The semantics of a reversible component is simply given by its Dirac structure interpreted as a relation:
+$\begin{split}-\mathrm{P}(\mathtt{(\mathbb{R},\,volume}))\>&=\>\mathrm{P}(-a)\,\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\\ -\mathrm{P}((\mathbb{R},\,\mathtt{volume}))\>&=\>\mathrm{P}(+a)\,\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\,.\end{split}$
+Checks for the other two non-zero entries of $L$ pass due to skew-symmetry of $L$. ∎
+The semantics of a reversible component is simply given by its Dirac structure interpreted as a relation.
 Definition 6.11.
@@
 Such models can be defined in terms of a function that maps thermodynamic forces, which drive the relaxation process, to thermodynamic fluxes, which seek to restore equilibrium.
-Based on the assumption that the microscopic dynamics are reversible, Onsager found that such maps must satisfy a symmetry property Onsager (1931).
-This motivates our definition of an Onsager structure:
+Based on the assumption that the microscopic dynamics are reversible, Onsager found that such maps must satisfy a symmetry property [15].
+This motivates our definition of an Onsager structure.
 Definition 6.12.
 Let $I$ be an interface. An Onsager structure $\mathcal{O}\subset\mathcal{P}_{I}$ is a fiber subbundle that admits the following representation:
-$\mathcal{O}\,=\,\left\{(x,\,f,\,e)\in\mathcal{P}_{I}\,\mid\,f\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}\,M(e)\,e\right\}\,,$
+$\mathcal{O}\,=\,\left\{(x,\,f,\,e)\in\mathcal{P}_{I}\,\mid\,f\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\,M(e)\,e\right\}\,,$
 where $M$ is a smooth function that yields a symmetric non-negative definite linear operator $M(e)\colon\mathrm{T}^{*}\mathcal{X}_{I,p}\to\mathrm{T}\mathcal{X}_{I,p}$ for all effort variables $e\in\mathrm{T}^{*}\mathcal{X}_{I,p}$.
 ∎
-In contrast to a Dirac structure, an Onsager structure is neither power-preserving, nor is it a locally linear relation of the port variables, since destruction of exergy, or equivalently production of entropy, is described in general by a non-linear relation.
+In contrast to a Dirac structure, an Onsager structure is neither power-preserving, nor is it a locally linear relation of the port variables, since destruction of exergy, or equivalently production of entropy, is usually described by a non-linear relation.
 Hence, $\mathcal{O}\subset\mathcal{P}_{I}$ is not a vector subbundle of $\mathcal{P}_{I}$, but merely a fiber subbundle.
-The symmetry of $M(e)$ corresponds to Onsager’s reciprocal relations and its non-negative definiteness implies a non-negative exergy destruction rate.
+The symmetry of $M(e)$ corresponds to Onsager’s reciprocal relations and its non-negative definiteness implies a non-negative exergy destruction rate, i.e.
+$\langle e\mid f\rangle\geq 0$.
 Example 6.13.
@@
 and let $d\in\mathbb{R}_{>0}$ be a friction coefficient. Then, the following defines an Onsager structure $\mathcal{O}_{\text{mf}}\subset\mathcal{P}_{I_{\text{mf}}}$ representing mechanical friction:
-$\begin{bmatrix}\mathtt{p.f}\\ \mathtt{s.f}\end{bmatrix}\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}\,d\,\begin{bmatrix}\theta&-\upsilon\\ -\upsilon&\frac{\upsilon^{2}}{\theta}\end{bmatrix}\,\begin{bmatrix}\mathtt{p.e}\\ \mathtt{s.e}\end{bmatrix}$
-Here, $\upsilon=\mathtt{p.e}$ is the velocity and $\theta={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}+\mathtt{s.e}$ is the absolute temperature at which kinetic energy is dissipated into the thermal energy domain.
+$\begin{bmatrix}\mathtt{p.f}\\ \mathtt{s.f}\end{bmatrix}\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\,d\,\begin{bmatrix}\theta&-\upsilon\\ -\upsilon&\frac{\upsilon^{2}}{\theta}\end{bmatrix}\,\begin{bmatrix}\mathtt{p.e}\\ \mathtt{s.e}\end{bmatrix}$
+Here, $\upsilon=\mathtt{p.e}$ is the velocity and $\theta={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}+\mathtt{s.e}$ is the absolute temperature at which kinetic energy is dissipated into the thermal energy domain, since adding ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}$ cancels the shift $-{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}$ in the effort variable, see 6.3.
 The phenomenological parameter $d$ could also be a function of the effort variables $e$, i.e.
@@
 The above can be simplified to $\mathtt{p.f}=d\,\upsilon$ and $\mathtt{s.f}=-\frac{1}{\theta}\,d\,\upsilon^{2}$.
-Hence, the environment temperature ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}$ cancels out in the flow variables, which here are the friction force and the entropy production rate (with a minus sign since thermal energy is leaving the component).
+Hence, the environment temperature ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}$ cancels out in the flow variables, which here are the friction force and the entropy production rate (with a minus sign since thermal energy is leaving the component).
+The exergy destruction rate is given by $\langle\mathtt{p.e}\mid\mathtt{p.f}\rangle+\langle\mathtt{s.e}\mid\mathtt{s.f}\rangle={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\,\frac{d\,\upsilon^{2}}{\theta}\geq 0$, with $\frac{d\,\upsilon^{2}}{\theta}\geq 0$ being the entropy production rate.
 ∎
-An irreversible component is defined by an Onsager structure which satisfies two extra properties ensuring thermodynamic consistency:
+An irreversible component is defined by an Onsager structure that satisfies two additional conditions ensuring thermodynamic consistency.
 Definition 6.14.
 An irreversible component $(I,\,\mathcal{O})$ is given by an interface $I$ and an Onsager structure $\mathcal{O}\subset\mathcal{P}_{I}$ such that energy and all quantities of the reference environment, except for entropy, are conserved and the time-reversal invariance property does strictly not hold.
-The condition for conservation of energy requires that
-[…]
-The condition that all quantities of the reference environment, except for entropy, are conserved requires that
-[…]
+The condition for conservation of energy requires that $\forall\,e\in\mathrm{T}^{*}\mathcal{X}_{I,p}\,\colon$
+$0\>=\>M(e)\,\bigl{(}\lambda(I)+e\bigr{)}\,.$
+The condition that all quantities present in the reference environment, except for entropy, are conserved requires that $\forall\,p_{R}\in I_{R}$ with $\tau_{R}(p_{R})\neq((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p})$, we have that $\forall\,e\in\mathrm{T}^{*}\mathcal{X}_{I,p}\,\colon$
+$0\>=\>M(e)\,\lambda(I,\,p_{R})\,.$
 The time-reversal invariance property does strictly not hold if $\forall\,x\in\mathcal{X}_{I}\,\colon$
@@
 Here, we use the same notation as in 6.9 with indices $i$ and $j$ ranging over all power ports of $I$. ∎
+From 6.4 and 6.3 we know that efforts variables are given by the differential of the energy function and additional shifts by the constant, intensive quantities of the environment, whenever the quantity of the respective port matches with a quantity present in the envirionment.
+We thus have $e=e^{\prime}-\lambda(I)$, where $e^{\prime}$ is precisely the part that stems from the differential of the energy.
+From this and the chain rule, it follows that energy is conserved if $0=\langle e^{\prime}\mid f\rangle=\langle e+\lambda(I)\mid M(e)\,e\rangle$.
+Taking the linear adjoint, this results in the condition in the definition, since $M(e)$ is symmetric.
+Regarding the condition that all quantities that are also present in the environment, except for entropy, are conserved, follows as already shown for reversible components, following 6.9.
+Again, regarding the time-reversal invariance condition, we refer to [46] or [3].
 Example 6.15.
@@
 $\begin{bmatrix}0\\ 0\end{bmatrix}\>=\>d\,\begin{bmatrix}\theta&-\upsilon\\ -\upsilon&\frac{\upsilon^{2}}{\theta}\end{bmatrix}\,\begin{bmatrix}\upsilon\\ \theta\end{bmatrix}\,.$
-The condition that all quantities of the reference environment, except for entropy, are conserved is trivially satisfied. The time-reversal invariance property does strictly not hold since
-$\begin{split}\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\>&=\>\mathrm{P}(d)\,\mathrm{P}(\theta)\,\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\\ \mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\>&=\>\mathrm{P}(d)\,\mathrm{P}(-\upsilon)\,\mathrm{P}((\mathbb{R},\,\mathtt{entropy}))\\ \mathrm{P}((\mathbb{R},\,\mathtt{entropy}))\>&=\>\mathrm{P}(d)\,\mathrm{P}(-\upsilon)\,\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\\ \mathrm{P}((\mathbb{R},\,\mathtt{entropy}))\>&=\>\mathrm{P}(d)\,{(\mathrm{P}(\upsilon))}^{2}\,/\,\mathrm{P}(\theta)\,\mathrm{P}((\mathbb{R},\,\mathtt{entropy}))\,,\end{split}$
-where we used that temperature has positive parity and velocity has negative parity. ∎
+The condition that all quantities of the reference environment, except for entropy, are conserved is trivially satisfied.
+As required, the time-reversal invariance property does strictly not hold since
+$\begin{split}\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\>&=\>\mathrm{P}(d)\,\mathrm{P}(\theta)\,\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\\ \mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\>&=\>\mathrm{P}(d)\,\mathrm{P}(-\upsilon)\,\mathrm{P}((\mathbb{R},\,\mathtt{entropy}))\\ \mathrm{P}((\mathbb{R},\,\mathtt{entropy}))\>&=\>\mathrm{P}(d)\,\mathrm{P}(-\upsilon)\,\mathrm{P}((\mathbb{R},\,\mathtt{momentum}))\\ \mathrm{P}((\mathbb{R},\,\mathtt{entropy}))\>&=\>\mathrm{P}(d)\,{(\mathrm{P}(\upsilon))}^{2}\,/\,\mathrm{P}(\theta)\,\mathrm{P}((\mathbb{R},\,\mathtt{entropy}))\,.\end{split}$
+Here, we used that temperature has positive parity and velocity has negative parity, which follows from the parities of entropy and momentum, knowing that energy has positive parity. ∎
+Irreversible components, and hence EPHS in general, respect Onsager’s reciprocal relations because the symmetry of $M(e)$ in 6.12 implies equal cross-effect coefficients (see, e.g., [48]).
+In the terminology of Linear Irreversible Thermodynamics (LIT), thermodynamic fluxes are related to thermodynamic forces by a symmetric operator.
+The following example shows this in detail.
 Example 6.16.
@@
 let $\alpha\in\mathbb{R}_{>0}$ be a heat transfer coefficient, and let the Onsager structure $\mathcal{O}_{\text{ht}}\subset\mathcal{P}_{I_{\text{ht}}}$ be given by
-$\begin{bmatrix}\mathtt{s_{1}.f}\\ \mathtt{s_{2}.f}\end{bmatrix}\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}\,\alpha\,\begin{bmatrix}\frac{\theta_{2}}{\theta_{1}}&-1\\ -1&\frac{\theta_{1}}{\theta_{2}}\end{bmatrix}\,\begin{bmatrix}\mathtt{s_{1}.e}\\ \mathtt{s_{2}.e}\end{bmatrix}\,,$
-where $\theta_{1}={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}+\mathtt{s_{1}.e}$ and $\theta_{2}={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}+\mathtt{s_{2}.e}$ are the absolute temperatures of the two thermal domains.
-Then, we can check that $(I_{\text{mf}},\,\mathcal{O}_{\text{mf}})$, defines an irreversible component.
+$\begin{bmatrix}\mathtt{s_{1}.f}\\ \mathtt{s_{2}.f}\end{bmatrix}\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\,\alpha\,\begin{bmatrix}\frac{\theta_{2}}{\theta_{1}}&-1\\ -1&\frac{\theta_{1}}{\theta_{2}}\end{bmatrix}\,\begin{bmatrix}\mathtt{s_{1}.e}\\ \mathtt{s_{2}.e}\end{bmatrix}\,,$
+where $\theta_{1}={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}+\mathtt{s_{1}.e}$ and $\theta_{2}={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}+\mathtt{s_{2}.e}$ are the absolute temperatures of the two thermal energy domains.
+It can be easily checked that $(I_{\text{ht}},\,\mathcal{O}_{\text{ht}})$ defines an irreversible component.
 Again, the phenomenological parameter $\alpha$ could for instance be a function of the temperatures.
 The above can be simplified to $\mathtt{s_{1}.f}=-\frac{1}{\theta_{1}}\,\alpha\,(\theta_{2}-\theta_{1})$ and $\mathtt{s_{2}.f}=-\frac{1}{\theta_{2}}\,\alpha\,(\theta_{1}-\theta_{2})$.
-Hence, the environment temperature ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}$ cancels out in the flow variables, which are rates of entropy change due to heat transfer.
+Hence, the environment temperature ${\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}$ cancels out in the flow variables, which are rates of entropy change due to heat transfer.
+Since $M(e)$ is symmetric and non-negative definite, an Onsager structure can be represented in a factorized form:
+$\underbrace{\begin{bmatrix}\mathtt{s_{1}.f}\\ \mathtt{s_{2}.f}\end{bmatrix}}_{f}\>=\>\underbrace{\begin{bmatrix}+\frac{1}{\theta_{1}}\\ -\frac{1}{\theta_{3}}\end{bmatrix}}_{C(e)}\,\underbrace{\begin{bmatrix}\alpha\,\theta_{1}\,\theta_{3}\end{bmatrix}}_{D(e)}\frac{1}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}}\,\underbrace{\begin{bmatrix}+\frac{1}{\theta_{1}}&-\frac{1}{\theta_{3}}\end{bmatrix}}_{C^{*}(e)}\,\underbrace{\begin{bmatrix}\mathtt{s_{1}.e}\\ \mathtt{s_{2}.e}\end{bmatrix}}_{e}$
+The factorization has a physical interpretation in terms of LIT (see, e.g., [13]). The thermodynamic force that drives the irreversible process is given by
+$X\>=\>\frac{1}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}}\,C^{*}(e)\,e\>=\>\frac{1}{\theta_{2}}-\frac{1}{\theta_{1}}\,.$
+Application of the symmetric operator $D(e)$ yields the thermodynamic flux
+$J\>=\>D(e)\,\frac{1}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}}\,C^{*}(e)\,e\>=\>\alpha\,(\theta_{1}-\theta_{2})\,.$
+Since $M(e)$ models a single process, $D(e)$ has rank $1$.
+Application of $C(e)$ yields the flow variables $f$.
+The net power $\langle\mathtt{s_{1}.e}\mid\mathtt{s_{1}.f}\rangle+\langle\mathtt{s_{2}.e}\mid\mathtt{s_{2}.f}\rangle=\langle e\mid f\rangle={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\,\frac{\alpha\,{(\theta_{1}-\theta_{2})}^{2}}{\theta_{1}\,\theta_{2}}\geq 0$ is the exergy destruction rate, and $\frac{\alpha\,{(\theta_{1}-\theta_{2})}^{2}}{\theta_{1}\,\theta_{2}}=\langle X\mid J\rangle\geq 0$ is the entropy production rate.
 ∎
-The semantics of an irreversible component is simply given by its Onsager structure interpreted as a relation:
+The semantics of an irreversible component is simply given by its Onsager structure interpreted as a relation.
 Definition 6.17.
@@
 7 Composite systems
-A composite system is formed by interconnecting finitely many systems, be they primitive systems or other composite systems, according to an interconnection pattern.
-The data defining a composite system is therefore given by an interconnection pattern and for each of its inner boxes a system with the corresponding interface.
+A composite system is formed by interconnecting a finite number of (primitive or composite) systems according to an interconnection pattern.
+The data defining a composite system consists of an interconnection pattern and, for each of its inner boxes, a system with the corresponding interface.
 Definition 7.1.
-Let $\mathbf{System}$ refer to the set of all systems defined with respect to a fixed reference environment. A composite system $(I,\,\gamma,\,J)$ is defined by the following data:
+Let $\mathsf{System}$ refer to the set of all systems defined with respect to a fixed reference environment. A composite system $(I,\,\gamma,\,J)$ is defined by the following data:
   1. 1.
@@
   2. 2.
-    A package of systems $\gamma\in\mathbf{Pkg}(\mathbf{System})$, which can be reduced to a package of interfaces $\beta\in\mathbf{Pkg}(\mathbf{Interface})$ by forgetting further data.
+    A directory of systems $\gamma\in\mathsf{Dtry}(\mathsf{System})$, which can naturally be reduced to a directory of interfaces $\beta\in\mathsf{Dtry}(\mathsf{Interface})$.
   3. 3.
     A partition $J$ of the combined interface $P=\sum\left[\mathtt{inner}\mapsto\sum\beta,\,\mathtt{outer}\mapsto I\right]$ such that $(\beta,\,I,\,J)$ defines an interconnection pattern. ∎
-class of system interface relation definition
+class of systems interface relation definition
 […]
@@
 Table 2: Overview of systems and their semantics given by a relation.
-Table 2 recalls the different classes of components defined in Section 6.
-In order to write the relations that determine their semantics in a consistent form, the Cartesian unit is inserted a few times.
-As already shown in the last row of the table, we can state the semantics of a composite system in the same form.
-The idea is that we just need to compose the relation assigned to the interconnection pattern with the Cartesian product of the relations assigned to the subsystems filling its inner boxes.
+Table 2 recalls the different kinds of components defined in Section 6.
+In order to write the relations that determine their semantics in a unified form, the Cartesian unit is inserted a few times.
+As shown in the last row of the table, the semantics of a composite system takes the same form.
+We just need to compose the relation assigned to the interconnection pattern with the Cartesian product of the relations assigned to the subsystems filling its inner boxes.
 Definition 7.2.
@@
 Further, let $\mathcal{P}_{P}$ be the relation assigned to pattern $P$ according to 4.2.
-For any inner box $b\in\operatorname{dom}(\gamma)$, let $\left\llbracket{s}\right\rrbracket$ denote the relation assigned to the subsystem $s=\gamma(b)$.
-In case system $s$ is itself a composite system, this definition applies recursively.
+For any filled inner box $b\in\operatorname{dom}(\gamma)$, let $\left\llbracket{s}\right\rrbracket$ denote the relation assigned to the subsystem $s=\gamma(b)$.
+In case system $s$ is itself a composite system, the present definition applies recursively.
 Then, the semantics assigned to composite system $(I,\,\gamma,\,J)$ is the relation $\mathcal{C}:\mathrm{T}\mathcal{X}_{I_{s}}\times\mathrm{T}\mathcal{X}_{I_{e}}\nrightarrow\mathcal{P}_{I}$ given by
 $\mathcal{C}\>=\>\mathcal{P}_{P}\circ\left(\prod_{b\,\in\,\operatorname{dom}(\gamma)}\left\llbracket{\gamma(b)}\right\rrbracket\right)\,.$
-Here, $\mathcal{X}_{I_{s}}$ denotes the state space of all storage components and $\mathcal{X}_{I_{e}}$ denotes the state space of all environment components, which are subsystems of the composite system in a recursive sense.
+Here, $\mathcal{X}_{I_{s}}$ denotes the state space of all storage components and $\mathcal{X}_{I_{e}}$ denotes the state space of all environment components that are subsystems of the composite system in a recursive sense.
 ∎
@@
 from $\mathrm{T}\mathcal{X}_{\text{tc}}\times 1$ to $\mathcal{P}_{I}$. ∎
-For an isolated composite system the assigned relation $\mathcal{C}\colon\mathrm{T}\mathcal{X}_{I_{s}}\times\mathrm{T}\mathcal{X}_{I_{e}}\nrightarrow 1$ simply amounts to a subbundle of $\mathrm{T}\mathcal{X}_{I_{s}}\times\mathrm{T}\mathcal{X}_{I_{e}}$, which should implicitly define a vector field on $\mathcal{X}_{I_{s}}\times\mathcal{X}_{I_{e}}$.
+For an isolated composite system the assigned relation $\mathcal{C}\colon\mathrm{T}\mathcal{X}_{I_{s}}\times\mathrm{T}\mathcal{X}_{I_{e}}\nrightarrow 1$ simply amounts to a subbundle of $\mathrm{T}\mathcal{X}_{I_{s}}\times\mathrm{T}\mathcal{X}_{I_{e}}$, which implicitly defines a vector field on $\mathcal{X}_{I_{s}}\times\mathcal{X}_{I_{e}}$.
 Given an initial condition, one can integrate this vector field to determine the evolution of the system.
 We simply say that $(x,\,\dot{x})\in\mathcal{C}$, where $x=(x_{s},\,x_{e})$, gives the dynamics of the isolated system.
-This is generalized by the following definition:
+The following definition generalizes this statement.
 Definition 7.4.
-Let $s\in\mathbf{System}$ be a system with interface $I$ and let the relation $\mathcal{R}\colon\mathrm{T}\mathcal{X}_{I_{s}}\times\mathrm{T}\mathcal{X}_{I_{e}}\nrightarrow\mathcal{P}_{I}$ define its semantics, according to one of the definitions mentioned in Table 2.
+Let $s\in\mathsf{System}$ be a system with interface $I$ and let the relation $\mathcal{R}\colon\mathrm{T}\mathcal{X}_{I_{s}}\times\mathrm{T}\mathcal{X}_{I_{e}}\nrightarrow\mathcal{P}_{I}$ define its semantics, according to the definitions mentioned in Table 2.
 Further, let $(x_{s},\,x_{e})\in\mathcal{X}_{I_{s}}\times\mathcal{X}_{I_{e}}$ denote the state variables associated to all (nested) storage and environment components.
-If $s$ is a primitive system this may be vacuous.
+If the system is a reversible or an irreversible component or given by an interconnection of only such components this is vacuous.
 Moreover, let $(x,f,e)\in\mathcal{P}_{I}$ denote the port variables of its (outer) interface.
-Again, if the system is isolated this is vacuous.
-Then, the dynamics of system $s$ is determined by the implicit differential-algebraic equation
+If the system is isolated this is vacuous.
+Then, the dynamics of system $s$ is implicitly determined by
 $\bigl{(}(x_{s},\,\dot{x}_{s}),\,(x_{e},\,\dot{x}_{e}),\,(x,\,f,\,e)\bigr{)}\in\mathcal{R}\,.\qed$
@@
 […]
-For programming languages it usually holds that not everything that is syntactically correct is also semantically meaningful.
-There are also cases where a compiler for the EPHS language should throw an error, despite of syntactic correctness.
-Every junction must have either zero or one connected port belonging to a storage or environment component.
-Also, one cannot connect more than two power ports with a quantity such as $(\mathbb{R},\,\mathtt{displacement})$ to the same junction, since velocities don’t balance.
-As of now, we don’t have a comprehensive set of restrictions that should be imposed on the interconnection of systems in order to get something meaningful, but we hope that this can be worked out in the future.
+In the context of programming languages, syntactically correct code does not necessarily define a semantically valid program.
+Similarly, there are cases where a compiler for the EPHS language must raise an error.
+For example, no junction can have two or more connected ports belonging to storage or environment components.
+A more comprehensive discussion of the conditions required for a composite system to be well-defined is left for future work.
 8 Thermodynamic consistency
-As already touched upon in the introduction, the GENERIC establishes a general structure, also called metriplectic structure, for combining reversible and irreversible dynamics in a way that respects fundamental principles from thermodynamics, namely the first law, the second law, and Onsager reciprocal relations.
-The vector field which determines the dynamics of a metriplectic system is given as the sum of a reversible Hamiltonian vector field that conserves energy due to the skew-symmetry of the Poisson bivector and a vector field of an irreversible gradient flow that has non-negative entropy production due to the non-negative definiteness of the metric tensor.
-Further, so-called non-interacting or degeneracy conditions ensure that the reversible dynamics conserves entropy and the irreversible dynamics conserves energy.
-The first law then holds since both parts conserve energy and the second law holds because the irreversible part has non-negative entropy production, while the reversible part conserves entropy, see e.g.
-Morrison (1984); Grmela and Öttinger (1997).
-Onsager reciprocal relations hold, as the symmetry of the metric tensor implies equal cross-effect coefficients, see e.g.
-Mielke et al.
-(2016).
-Usually, the reversible part and the irreversible part both satisfy further degeneracy conditions that ensure the conservation of further quantities such as volume or mass.
-Given that (some of) these quantities, together with entropy, are present in an exergy reference environment that is adjoined to the metriplectic system, it follows that the reversible dynamics conserve exergy and the irreversible dynamics are characterized by a non-negative exergy destruction.
-This immediately follows from the definition of the exergy storage function analogous to 6.2.
-The argument for EPHS is very similar in nature, although it is in a sense reversed. To keep it short and simple, we focus on isolated systems.
+Akin to metriplectic systems or the GENERIC, the structure of EPHS ensures thermodynamic consistency. To simplify things a bit, we focus on isolated systems.
 Proposition 8.1.
-The dynamics of any isolated EPHS respects the first law, the second law, and Onsager reciprocal relations. ∎
+The dynamics of an isolated EPHS respect the first law and the second law. ∎
+To prove this, one could first show that the semantics of any (isolated) composite system can be represented by some abstract system of evolution equations.
+Analogous to metriplectic systems or the GENERIC, one could then use the structural properties of this abstract system of equations to demonstrate thermodynamic consistency.
+The following proof, however, does not require this intermediate step.
 Proof.
-We argue in terms of components and their interconnection.
+As the semantics associated to interconnection patterns is functorial, it suffices to show thermodynamic consistency for composite systems, where all subsystems are primitive.
+In terms of semantics, environment components are a special case of storage components.
+Any finite number of storage components can be combined into a single component by taking the named sum of their interfaces and by summing their energy functions.
+Similarly, any finite number of reversible/irreversible components can be combined into a single such component by taking the named sum of their interfaces and by taking the Cartesian product of the relations that define them.
+It hence suffices to show thermodynamic consistency for a composite system with one storage component, one reversible component, and one irreversible component.
   1. 1.
-    According to 6.9, all reversible components conserve exergy as well as all quantities of the reference environment.
-    Considering the definition of exergy in 6.2, it follows immediately that all reversible components also conserve energy.
+    First, we consider the storage component and focus specifically on the structure of its exergy function $H$, see 6.2. The following two facts follow immediately:
+    1. (a)
+      The energy $E$ must be conserved if the exergy $H$ and all quantities also present in the reference environment are conserved.
+    2. (b)
+      Entropy must be growing at a non-negative rate if the exergy $H$ is decreasing at a non-negative rate, while the energy $E$ and all quantities also present in the reference environment, except for entropy, are conserved.
   2. 2.
-    According to 6.14, all irreversible components are defined by a non-negative exergy destruction rate and they conserve energy as well as all quantities of the reference environment, except for entropy.
-    Considering the definition of exergy, it follows that all irreversible components also have a non-negative entropy production rate.
+    Regarding the interconnection of the components based on 4.2, we know that at each junction the flow variables balance, while the effort variables are equal. We can conclude the following:
+    1. (a)
+      The interconnection is power-preserving, i.e. it conserves exergy.
+    2. (b)
+      The interconnection conserves the quantities associated with the ports because each flow variable represents the rate of change of the corresponding quantity, and the flow variables balance.
+    3. (c)
+      The interconnection conserves energy. This follows from the previous two points and point 1. (a)
   3. 3.
-    According to 4.2, any interconnection of the components conserves exergy and all quantities (of the reference environment), since flows balance at equal efforts.
-    Considering the definition of exergy, it follows that any interconnection of the components also conserves energy.
-The first law holds since all three aspects conserve energy.
-The second law holds because the irreversible components have a non-negative entropy production rate, while the reversible components and the interconnection conserve entropy.
-The reciprocal relations hold because, according to 6.12, thermodynamic fluxes (flows) depend on thermodynamic forces (efforts) via a symmetric locally-linear operator.
-∎
-We note that the argument is reversed because, according to Definitions 6.4 and 7.2, the semantics of EPHS is defined in terms of a storage function that is interpreted as the exergy of the system, as opposed to two separate functions representing the energy and entropy of a system.
-The concept of a storage function originates in Willems’ theory of dissipative systems Willems (1972).
-In this context, reversible components as well as all EPHS without any irreversible components are said to be lossless, while (all systems with) irreversible components are said to be dissipative.
-9 Example
-This section compares a bond graph and an EPHS model of a DC shunt motor.
-A direct current (DC) motor has a fixed outer part, called stator, and a rotating inner part, called rotor.
+    Next, we consider the reversible component defined by a Dirac structure $\mathcal{D}$, see 6.7.
+Since $\mathcal{D}$ is a power-preserving relation, it conserves the exergy $H$.
+One of the conditions in 6.9 guarantees that $\mathcal{D}$ conserves all quantities also present in the reference environment, in particular entropy.
+Considering also point 1.
+(a), it follows that $\mathcal{D}$ must also conserve the energy $E$.
+  4. 4.
+    Finally, we consider the irreversible component defined by an Onsager structure $\mathcal{O}$, see 6.12.
+As a consequence of the non-negative definiteness property, $\mathcal{O}$ destroys exergy at a non-negative rate.
+One of the conditions in 6.14 guarantees that $\mathcal{O}$ conserves the energy $E$.
+Another condition guarantees that $\mathcal{O}$ conserves all quantities also present in the reference environment, except for entropy.
+Considering also point 1.
+(b), it follows that $\mathcal{O}$ must produce entropy at a non-negative rate.
+The composite system must respect the first law, since the interconnection, the reversible component, and the irreversible component all conserve energy.
+The composite system must respect the second law, since the interconnection and the reversible component conserve entropy, while the irreversible component produces entropy at a non-negative rate. ∎
+The concept of a storage function was introduced in Willems’ theory of dissipative systems [49].
+In this framework, EPHS without irreversible components are classified as lossless, whereas EPHS that include irreversible components are considered dissipative or passive.
+The additional conditions defining reversible and irreversible components in EPHS extend the notion of dissipativity, ensuring full thermodynamic consistency.
+9 Example: shunt motor
+This section compares a bond graph and an EPHS model of a direct current (DC) shunt motor.
+A DC motor has a fixed outer part, called stator, and a rotating inner part, called rotor.
 Both parts contain a coil of wire through which current flows.
-In the case of a shunt motor, these two inductors are connected in parallel (aka in shunt) to a source of electric energy.
-The working principle of such motors is based on a coupling of the electro-magnetic and the kinetic energy domains via the Lorentz force and the commutation of the rotor coil after every half turn.
-However, both aspects are not directly taken into account by the two models, which instead consider the coupling as being lumped into a gyrator element.
+In the case of a shunt motor, these two inductors are connected in parallel to a source of electric energy.
+The working principle of such motors is based on a coupling of the electromagnetic and the kinetic energy domains via the Lorentz force and the commutation of the rotor coil after every half turn.
+However, both aspects are not directly resolved by the considered models.
+Instead, the coupling is considered to be lumped into a gyrator element.
 Figure 10: Bond graph model of a DC shunt motor.
-We start with the bond graph model that is taken from Borutzky (2010) and redrawn in Figure 10.
+We start with the bond graph model that is taken from [8] and reproduced, with slightly different naming of variables, in Figure 10.
 The half arrows are called bonds and each bond has two associated variables, called flow and effort.
@@
 However, not all variables are explicitly named in the example.
-At any 00-junction the effort variables of all connected bonds are equal, while the flow variables balance, with their sign given by the bond direction.
-Dually, at $1$-junctions flows are equal and efforts balance.
-Specifically, at the 00-junction on the left the current from the $\mathrm{Se}$-element (effort source) with effort/voltage $E$ splits into the current $i_{s}$ through the stator coil and the current $i_{r}$ through the rotor coil.
+At any 00-junction, the effort variables of all connected bonds are equal, while the flow variables balance, with their sign given by the bond direction.
+Dually, at $1$-junctions, flows are equal and efforts balance.
+Specifically, at the 00-junction on the left, the current from the $\mathrm{Se}$-element (effort source) with effort/voltage $E$ splits into the current $i_{s}$ through the stator coil and the current $i_{r}$ through the rotor coil.
 The inductance of the rotor coil is represented by the $\mathrm{I}$-element (generalized inductor) with inductance $L_{r}$, while an $\mathrm{I}$-element representing the stator coil inductance is omitted.
 The (steady-state) stator coil current $i_{s}$ is instead entirely determined by the $\mathrm{R}$-element (generalized resistor) at the top representing the coil’s electrical resistance $R_{s}$.
-The magnetic flux of the stator coil $b_{s}$ is instead obtained as the signal that results from multiplying the signal $i_{s}$, which is the shared effort variable at the upper $1$-junction, with the constant gain $L_{s}$.
-The bond graph hence combines energy-based modeling with signal-based modeling known from block diagrams.
+The magnetic flux of the stator coil $b_{s}$ is then obtained as the signal that results from multiplying the signal $i_{s}$, which is the shared effort variable at the upper $1$-junction, with the constant gain $L_{s}$.
+The bond graph hence combines energy-based modeling with signal-based modeling based on block diagrams.
 At the $1$-junction on the left, the source voltage $E$ is equal to the sum of the voltage over the rotor coil inductance $L_{r}$, the voltage $u_{R}$ over the rotor coil resistance $R_{r}$, and the induced (back EMF) voltage $u_{r}$.
@@
 As already shown in 4.3, the EPHS model is specified as an interconnection of two subsystems named $\mathtt{stator}$ and $\mathtt{rotor}$.
-The two systems share an electric energy domain represented by the junction with connected ports named $\mathtt{q}$.
+The two systems have a common electric energy domain represented by the junction with connected ports named $\mathtt{q}$.
 The kinetic energy domain of the rotor is represented by the junction with connected ports named $\mathtt{p}$.
@@
 The interconnection pattern of the stator model is shown in Figure 11.
-The box $\mathtt{em}$ is filled by a reversible component that represents the coupling of the electric energy domain with port $\mathtt{q}$ and the magnetic energy domain with port $\mathtt{b_{s}}$.
-The box $\mathtt{coil}$ is filled by a storage component that represents the inductance of the stator coil, while the box $\mathtt{emloss}$ is filled by an irreversible component that represents its electric resistance.
-The box $\mathtt{env}$ is filled by an environment component whose thermal energy domain with port $\mathtt{s}$ absorbs the dissipated electromagnetic energy.
-The reversible component $(I_{\text{em}},\,\mathcal{D}_{\text{em}})$ is defined by its interface $I_{\text{em}}=(\{\mathtt{q},\,\mathtt{b_{s}}\},\,\tau_{\text{em}})$ with $\tau_{\text{em}}(\mathtt{q})=((\mathbb{R},\,\mathtt{charge}),\,\mathsf{p})$, $\tau_{\text{em}}(\mathtt{b_{s}})=((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{p})$ and its Dirac structure $\mathcal{D}_{\text{em}}\subset\mathcal{P}_{I_{\text{em}}}$ given by
+Box $\mathtt{emc}$ is filled by a reversible component that represents the coupling of the electric energy domain associated with port $\mathtt{q}$ and the magnetic energy domain associated with port $\mathtt{b_{s}}$.
+Box $\mathtt{coil}$ is filled by a storage component that represents the inductance of the stator coil.
+Further, box $\mathtt{res}$ is filled by an irreversible component that represents the resistance of the coil.
+Finally, box $\mathtt{env}$ is filled by an environment component whose thermal energy domain associated with port $\mathtt{s}$ absorbs the dissipated electromagnetic energy.
+The reversible component $(I_{\text{emc}},\,\mathcal{D}_{\text{emc}})$ is defined by its interface
+$I_{\text{emc}}=(\{\mathtt{q},\,\mathtt{b_{s}}\},\,\tau_{\text{emc}})$ with
+$\tau_{\text{emc}}(\mathtt{q})\>$ $=\>((\mathbb{R},\,\mathtt{charge}),\,\mathsf{p})$
+$\tau_{\text{emc}}(\mathtt{b_{s}})\>$ $=\>((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{p})$
+and its Dirac structure $\mathcal{D}_{\text{emc}}\subset\mathcal{P}_{I_{\text{emc}}}$ given by
 $\begin{bmatrix}\mathtt{q.f}\\ \mathtt{b_{s}.f}\end{bmatrix}\>=\>\begin{bmatrix}0&1\\ -1&0\end{bmatrix}\,\begin{bmatrix}\mathtt{q.e}\\ \mathtt{b_{s}.e}\end{bmatrix}\,.$
-The storage component $(I_{\text{coil}},\,E_{\text{coil}})$ is defined by its interface $I_{\text{coil}}=(\{\mathtt{b_{s}}\},\,\tau_{\text{coil}})$ with $\tau_{\text{coil}}(\mathtt{b_{s}})=((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{p})$ and its energy function $E_{\text{coil}}:\mathcal{X}_{I_{\text{coil}}}\rightarrow\mathbb{R}$ with $\mathcal{X}_{I_{\mathtt{coil}}}=\mathbb{R}\ni\mathtt{b_{s}.x}=b_{s}$ given by
+The storage component $(I_{\text{coil}},\,E_{\text{coil}})$ is defined by its interface $I_{\text{coil}}=(\{\mathtt{b_{s}}\},\,\tau_{\text{coil}})$ with $\tau_{\text{coil}}(\mathtt{b_{s}})=((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{p})$ and its energy function $E_{\text{coil}}:\mathcal{X}_{I_{\text{coil}}}\rightarrow\mathbb{R}$ with $\mathcal{X}_{I_{\mathtt{coil}}}=\mathbb{R}\ni\mathtt{b_{s}.x}=b_{s}$.
+The function is given by
 $E_{\mathtt{coil}}(b_{s})\>=\>\frac{1}{2\,L_{s}}\,b_{s}^{2}\,,$
-where the parameter $L_{s}$ is the inductance. Since $H_{\mathtt{coil}}=E_{\mathtt{coil}}$, the effort variable $\mathtt{b_{s}.e}=\frac{b_{s}}{L_{s}}$ is the current through the coil.
-The irreversible component $(I_{\text{emloss}},\,\mathcal{O}_{\text{emloss}})$ is defined by $I_{\text{emloss}}=(\{\mathtt{b_{s}},\,\mathtt{s}\},\,\tau_{\text{emloss}})$ with $\tau_{\text{emloss}}(\mathtt{b_{s}})=((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{p})$, $\tau_{\text{emloss}}(\mathtt{s})=((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p})$ and its Onsager structure $\mathcal{O}_{\text{emloss}}\subset\mathcal{P}_{\text{emloss}}$ given by
-$\begin{bmatrix}\mathtt{b_{s}.f}\\ \mathtt{s.f}\end{bmatrix}\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}\,R_{s}\,\begin{bmatrix}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}+}\mathtt{s.e}&-\mathtt{b_{s}.e}\\ -\mathtt{b_{s}.e}&\frac{{(\mathtt{b_{s}.e})}^{2}}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}+}\mathtt{s.e}}\end{bmatrix}\,\begin{bmatrix}\mathtt{b_{s}.e}\\ \mathtt{s.e}\end{bmatrix}\>=\>\begin{bmatrix}R_{s}\,\mathtt{b_{s}.e}\\ -\frac{R_{s}\,{(\mathtt{b_{s}.e})}^{2}}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}+}\mathtt{s.e}}\end{bmatrix}\,,$
-where the parameter $R_{s}$ is the electric resistance of the coil.
+where the parameter $L_{s}$ is the inductance. The effort variable $\mathtt{b_{s}.e}=\frac{b_{s}}{L_{s}}$ is the current through the coil.
+The irreversible component $(I_{\text{res}},\,\mathcal{O}_{\text{res}})$ is defined by its interface $I_{\text{res}}=(\{\mathtt{b_{s}},\,\mathtt{s}\},\,\tau_{\text{res}})$ with
+$\tau_{\text{res}}(\mathtt{b_{s}})\>$ $=\>((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{p})$
+$\tau_{\text{res}}(\mathtt{s})\>$ $=\>((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p})$
+and its Onsager structure $\mathcal{O}_{\text{res}}\subset\mathcal{P}_{\text{res}}$ given by
+$\begin{bmatrix}\mathtt{b_{s}.f}\\ \mathtt{s.f}\end{bmatrix}\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\,R_{s}\,\begin{bmatrix}\theta&-i\\ -i&\frac{i^{2}}{\theta}\end{bmatrix}\,\begin{bmatrix}\mathtt{b_{s}.e}\\ \mathtt{s.e}\end{bmatrix}\>=\>\begin{bmatrix}R_{s}\,i\\ -\frac{1}{\theta}\,R_{s}\,i^{2}\end{bmatrix}\,,$
+where the parameter $R_{s}$ is the resistance of the coil, $i=\mathtt{b_{s}.e}$ is the current through the coil, and $\theta={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}+}\mathtt{s.e}$ is the absolute temperature at which heat is dissipated into the thermal energy domain.
 The environment component is defined by the subinterface $I_{\text{env}}=\{\mathtt{s}\}\subseteq I_{R}$.
@@
 The interconnection pattern of the rotor model is shown in Figure 12.
-The components filling the boxes $\mathtt{coil}$, $\mathtt{em}$, $\mathtt{emloss}$, and $\mathtt{env}$ are defined as for the stator model, except for a different port name, namely $\mathtt{b}$ instead of $\mathtt{b_{s}}$, and different parameters, namely $L_{r}$ instead of $L_{s}$ and $R_{r}$ instead of $R_{s}$.
-The box $\mathtt{mk}$ is filled by a reversible component that describes the coupling of the magnetic energy domain with port $\mathtt{b}$ and the (rotational) kinetic energy domain with port $\mathtt{p}$.
+The components filling the boxes $\mathtt{coil}$, $\mathtt{emc}$, $\mathtt{res}$ and $\mathtt{env}$ are defined as for the stator model, except for a different port name, namely $\mathtt{b}$ instead of $\mathtt{b_{s}}$, and different parameters, namely $L_{r}$ instead of $L_{s}$ and $R_{r}$ instead of $R_{s}$.
+Further, box $\mathtt{mkc}$ is filled by a reversible component that describes the coupling of the magnetic energy domain associated with port $\mathtt{b}$ and the (rotational) kinetic energy domain associated with port $\mathtt{p}$.
 This coupling depends on the magnetic flux $\mathtt{b_{s}.x}$ that is generated by the stator coil.
-The box $\mathtt{mass}$ is filled by a storage component that represents the rotor’s angular mass and the box $\mathtt{mloss}$ is filled by an irreversible component that represents its mechanical friction.
-The reversible component $(I_{\text{mk}},\mathcal{D}_{\text{mk}})$ is defined by $I_{\text{mk}}=(\{\mathtt{b},\,\mathtt{p},\,\mathtt{b_{s}}\},\,\tau_{\text{mk}})$ with $\tau_{\text{mk}}(\mathtt{b})=((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{p})$, $\tau_{\text{mk}}(\mathtt{p})=((\mathbb{R},\,\mathtt{angular\_momentum}),\,\mathsf{p})$, $\tau_{\text{mk}}(\mathtt{b_{s}})=((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{s})$ and its Dirac structure $\mathcal{D}_{\mathtt{mk}}\subset\mathcal{P}_{I_{\text{mk}}}$ given by
+Box $\mathtt{mass}$ is filled by a storage component that represents the rotor’s angular mass and box $\mathtt{mf}$ is filled by an irreversible component that represents its mechanical friction.
+The reversible component $(I_{\text{mkc}},\mathcal{D}_{\text{mkc}})$ is defined by its interface
+$I_{\text{mkc}}=(\{\mathtt{b},\,\mathtt{p},\,\mathtt{b_{s}}\},\,\tau_{\text{mkc}})$ with
+$\tau_{\text{mkc}}(\mathtt{b})\>$ $=\>((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{p})$
+$\tau_{\text{mkc}}(\mathtt{p})\>$ $=\>((\mathbb{R},\,\mathtt{angular\_momentum}),\,\mathsf{p})$
+$\tau_{\text{mkc}}(\mathtt{b_{s}})\>$ $=\>((\mathbb{R},\,\mathtt{flux\_linkage}),\,\mathsf{s})$
+and its Dirac structure $\mathcal{D}_{\mathtt{mkc}}\subset\mathcal{P}_{I_{\text{mkc}}}$ given by
 $\begin{bmatrix}\mathtt{b.f}\\ \mathtt{p.f}\end{bmatrix}\>=\>\begin{bmatrix}0&\mathtt{b_{s}.x}\\ -\mathtt{b_{s}.x}&0\end{bmatrix}\,\begin{bmatrix}\mathtt{b.e}\\ \mathtt{p.e}\end{bmatrix}\,.$
-The storage component $(I_{\text{mass}},\,E_{\text{mass}})$ is defined by $I_{\text{mass}}=(\{\mathtt{p}\},\,\tau_{\text{mass}})$ with $\tau_{\text{mass}}(\mathtt{p})=((\mathbb{R},\,\mathtt{angular\_momentum}),\,\mathsf{p})$ and $E_{\text{mass}}:\mathcal{X}_{I_{\text{mass}}}\rightarrow\mathbb{R}$ with $\mathcal{X}_{I_{\text{mass}}}=\mathbb{R}\ni\mathtt{p.x}=p$ given by
+The storage component $(I_{\text{mass}},\,E_{\text{mass}})$ is defined by its interface $I_{\text{mass}}=(\{\mathtt{p}\},\,\tau_{\text{mass}})$ with $\tau_{\text{mass}}(\mathtt{p})=((\mathbb{R},\,\mathtt{angular\_momentum}),\,\mathsf{p})$ and its energy function
+$E_{\text{mass}}:\mathcal{X}_{I_{\text{mass}}}\rightarrow\mathbb{R}$ with $\mathcal{X}_{I_{\text{mass}}}=\mathbb{R}\ni\mathtt{p.x}=p$. The function is given by
 $E_{\text{mass}}(p)\>=\>\frac{1}{2\,J_{r}}\,p^{2}\,,$
 where the parameter $J_{r}$ is the angular mass.
-The irreversible component $(I_{\text{mloss}},\mathcal{O}_{\text{mloss}})$ is defined by $I_{\text{mloss}}=(\{\mathtt{p},\,\mathtt{s}\},\,\tau_{\text{mloss}})$ with $\tau_{\text{mloss}}(\mathtt{p})=((\mathbb{R},\,\mathtt{angular\_momentum}),\,\mathsf{p})$, $\tau_{\text{mloss}}(\mathtt{s})=((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p})$ and its Onsager structure $\mathcal{O}_{\text{mloss}}\subset\mathcal{P}_{I_{\text{mloss}}}$ given by
-$\begin{bmatrix}\mathtt{p.f}\\ \mathtt{s.f}\end{bmatrix}\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}\,d_{r}\,\begin{bmatrix}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}+}\mathtt{s.e}&-\mathtt{p.e}\\ -\mathtt{p.e}&\frac{{(\mathtt{p.e})}^{2}}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}+}\mathtt{s.e}}\end{bmatrix}\,\begin{bmatrix}\mathtt{p.e}\\ \mathtt{s.e}\end{bmatrix}\>=\>\begin{bmatrix}d_{r}\,\mathtt{p.e}\\ -\frac{d_{r}\,{(\mathtt{p.e})}^{2}}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}+}\mathtt{s.e}}\end{bmatrix}\,,$
-where the parameter $d_{r}$ is the mechanical friction coefficient.
+The irreversible component $(I_{\text{mf}},\mathcal{O}_{\text{mf}})$ is defined by its interface $I_{\text{mf}}=(\{\mathtt{p},\,\mathtt{s}\},\,\tau_{\text{mf}})$ with
+$\tau_{\text{mf}}(\mathtt{p})\>$ $=\>((\mathbb{R},\,\mathtt{angular\_momentum}),\,\mathsf{p})$
+$\tau_{\text{mf}}(\mathtt{s})\>$ $=\>((\mathbb{R},\,\mathtt{entropy}),\,\mathsf{p})$
+and its Onsager structure $\mathcal{O}_{\text{mf}}\subset\mathcal{P}_{I_{\text{mf}}}$ given by
+$\begin{bmatrix}\mathtt{p.f}\\ \mathtt{s.f}\end{bmatrix}\>=\>\frac{1}{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}\,d_{r}\,\begin{bmatrix}\theta&-\upsilon\\ -\upsilon&\frac{\upsilon^{2}}{\theta}\end{bmatrix}\,\begin{bmatrix}\mathtt{p.e}\\ \mathtt{s.e}\end{bmatrix}\>=\>\begin{bmatrix}d_{r}\,\upsilon\\ -\frac{d_{r}\,\upsilon^{2}}{\theta}\end{bmatrix}\,,$
+where the parameter $d_{r}$ is the mechanical friction coefficient, $\upsilon=\mathtt{p.e}$ is the angular velocity, and $\theta={\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}+}\mathtt{s.e}$ is the absolute temperature at which heat is dissipated into the thermal energy domain.
 We write $(b_{s},\,b,\,p,\,s)\in\mathbb{R}^{4}$ for the state of the interconnected motor model, where we identify the entropy state variables $s$ of both environment components.
-In future work, the identification of environment components can be formalized by composing with a suitable relation.
+In a future work, the identification of environment components can be formalized by composing with a suitable relation.
 The EPHS model then yields the equations
 $\dot{b}_{s}\>$ $=\>\mathtt{q.e}-R_{s}\,i_{s}$
-$\dot{b}\>$ $=\>\mathtt{q.e}-b_{s}\,\omega-R_{r}\,i_{r}$
+$\dot{b}\>$ $=\>\mathtt{q.e}-R_{r}\,i_{r}-b_{s}\,\omega$
 $\dot{p}\>$ $=\>b_{s}\,i_{r}-d_{r}\,\omega+\mathtt{p.f}$
-$\dot{s}\>$ $=\>\frac{1}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\pgfsys@color@rgb@stroke{.5}{0}{.5}\pgfsys@color@rgb@fill{.5}{0}{.5}\theta_{0}}}\left(R_{s}\,i_{s}^{2}+R_{r}\,i_{r}^{2}+d_{r}\,\omega^{2}\right)$
-$\mathtt{q.f}\>$ $=\>i_{s}+i_{r}\,,$
+$\dot{s}\>$ $=\>\frac{1}{{\color[rgb]{.5,0,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,0,.5}\theta_{0}}}\left(R_{s}\,i_{s}^{2}+R_{r}\,i_{r}^{2}+d_{r}\,\omega^{2}\right)$
+$\mathtt{q.f}\>$ $=\>i_{s}+i_{r}$
+$\mathtt{p.e}\>$ $=\>\omega\,,$
 where
@@
 $i_{r}\>$ $=\>\frac{1}{L_{r}}\,b$
-$\omega\>$ $=\>\frac{1}{J_{r}}\,p\,.$
+$\omega\>$ $=\>\frac{1}{J_{r}}\,p\,,$
+which are similar to the equations for the bond graph model, except that (1) the source of electrical energy and the mechanical load are regarded as external dynamical systems, rather than internal constants, and (2) no quasi-static assumption is hence made for the magnetic field of the stator.
 10 Discussion
-While it still seems reasonable to package a motor model as a single bond graph, integrating the model as part of a more complex system is not supported by the bond graph notation.
-Hence, bond graph users have to reproduce models as parts of more complex models.
-EPHS improves upon this, making it nearly trivial to encapsulate, reuse, and swap out submodels.
-Another main feature of EPHS is their built-in thermodynamic consistency.
-Restricting the space of models in this way not only provides important guarantees, it also makes models compatible with each other, as a close link between their mathematical structure and physical interpretation is established.
-This is reflected by the correspondence of junctions and energy domains, which in our opinion contributes substantially to the improved readability of the EPHS syntax.
-The reason why EPHS users can easily integrate existing models without first needing to adapt them to fit into the larger system is that EPHS simply compose by sharing energy domains.
-In this way, we hope to improve upon port-Hamiltonian systems, which lack an intuitive graphical syntax and require that users define the ports of a system in a somewhat arbitrary manner.
+10.1 Comparison with bond graphs
+Bond graphs provide a graphical notation for representing physical systems in terms of subsystems that exchange energy.
+However, the notation does not naturally support the hierarchical nesting of systems.
+Relying exclusively on primitive subsystems limits the applicability of bond graphs to complex systems, where a hierarchical breakdown into simpler parts would facilitate both model construction and reuse.
+Although the notation is not inherently composable, one can introduce conventions that allow for subsystems defined by open bond graphs.
+For example, The convention in [50] considers bond graphs that share a specific element in common as a modules.
+These modules can be ‘glued’ together in a parent-level bond graph via an additional 00-junction.
+In this approach, the common element in each individual bond graph is replaced by an external port; the element is instantiated once in a parent-level bond graph, where it is connected to a 00-junction.
+The modules are represented by additional elements, and their ports are also connected to the 00-junction.
+The direction of these bonds must match the direction of the corresponding bond (or port) at the module level.
+While port-Hamiltonian systems can be seen as a geometric formalization of the dynamical systems derived from generalized bond graphs, the algebraic formalization of the bond graph notation itself is not straightforward (see, e.g., [51, 32]).
+Some inherent complexity in the notation is evident also in practice, where the direction in which exchange of power is counted as positive, and sometimes also the role of a variable as flow versus effort, must be decided on a bond-by-bond basis.
+These decisions must either be left to the user or managed through heuristic algorithms or conventions.
+This can impede seamless hierarchical decomposition and reuse of subsystems.
+Interconnection patterns provide a composable, graphical syntax for the EPHS modeling language.
+Compared to bond graphs, interconnection patterns have an outer box, and they exclusively rely on 00-junctions.
+Based on this simplification, the mentioned issues are resolved:
+  • 1.
+    Interconnection patterns have a clear mathematical structure, which supports their hierarchical (de-)composition.
+  • 2.
+    Interconnection patterns are relatively easy to interpret, owing to their simple structure, the direct correspondence between junctions and energy domains, and the clear role of a power variables as either flow or effort.
+  • 3.
+    The direction in which exchange of power is counted positive is predetermined in agreement with the convention used in thermodynamics, i.e. energy supplied to a system has a positive sign.
+In contrast to bond graphs, EPHS models are guaranteed to be thermodynamically consistent.
+Bond graphs may incorporate elements from block diagrams.
+In addition to bonds, which represent energy exchange, there may be edges that transmit signals according to their orientation.
+Elements may have signal ports to receive signals, which modulate their behavior, and there may be ‘blocks’ that perform arbitrary signal processing.
+In contrast, the formalism presented here does not cover arbitrary signal processing; system behavior can be modulated only by state variables of other systems.
+To express this, every port has a state variable, which is shared with all other ports connected to the same junction.
+We distinguish between state ports, which share only state, and power ports, which additionally model energy exchange.
+10.2 Comparison with port-Hamiltonian systems
+Port-Hamiltonian Systems (PHS) provide a mathematical framework for open dynamical systems with guaranteed passivity.
+While bond graphs are frequently used to depict port-Hamiltonian models, establishing a rigorous link between bond graph syntax and port-Hamiltonian semantics remains challenging.
+On the one hand, bond graphs offer a network representation of individual systems, but the representation is not inherently composable.
+On the other hand, PHS provide a composable representation of individual systems, albeit without an explicit internal network structure.
+In particular, the PHS framework provides no built-in support for referencing subsystems and their ports.
+Previous work has focused on algorithms for converting certain classes of bond graphs to PHS (see, e.g., [52]).
+A well-known issue concerns the relationship between port-Hamiltonian structure, which implies passivity, and its physical interpretation.
+Owing to its roots in mechanics, the ‘Hamiltonian’ storage function is typically interpreted as the system’s energy.
+However, the differential of the ‘Hamiltonian’ generally also drives dissipative dynamics, resulting in an apparent loss of ‘energy’.
+This raises questions, especially outside the realm of electromechanical systems, where thermodynamic considerations are often overlooked.
+In response, some members of the port-Hamiltonian community have proposed alternative port-based frameworks that explicitly encode thermodynamic structure (see, e.g., [53, 54, 55]).
+EPHS provide a mathematical framework for open dynamical systems with guaranteed thermodynamic consistency.
+In contrast to PHS, this framework is directly based on a composable, graphical syntax used to specify how more complex systems are formed by interconnecting simpler systems.
+An explicit understanding of the translation between graphical syntax and relational semantics guarantees that syntax and semantics compose in compatible ways.
+This enables a modular and hierarchical modeling approach, which inherently supports referencing nested subsystems and their ports.
+In a sense, this approach unites the strengths of both bond graphs and PHS.
+The dynamical systems defined by EPHS models essentially possess port-Hamiltonian structure.
+Beyond passivity, additional structural properties imposed on the primitive systems guarantee that all EPHS share a uniform and consistent physical interpretation, aligned with nonequilibrium thermodynamics.
+Since EPHS compose simply by sharing energy domains, users typically do not need to explicitly define external ports.
+As a result, users can typically integrate existing models into more complex systems without any prior adaptation.
+10.3 Comparison with metriplectic and GENERIC systems
+Metriplectic systems and the GENERIC provide a framework for (usually isolated) dynamical systems with guaranteed thermodynamic consistency.
+The dynamical systems defined by (isolated) EPHS models are basically metriplectic systems or instances of the GENERIC, since the built-in thermodynamic consistency follows from essentially the same ideas.
+Regarding the reversible part, EPHS appear to be more general, since they use Dirac structures, rather than the less general Poisson structures used in the definition of metriplectic systems or the GENERIC.
+This allows EPHS to easily model kinematic constraints, which is important e.g.
+for multibody systems (see [56]).
+However, it should be no problem to generalize the metriplectic or GENERIC framework accordingly.
+Regarding the irreversible part, EPHS appears to be less general, since it remains unclear if the single-generator formulation based on exergy is useful beyond local thermodynamic equilibrium.
+In contrast to the metriplectic or GENERIC framework, EPHS enable a modular and hierarchical approach to thermodynamic modeling, based on a graphical syntax for building more complex models from simpler parts.
+10.4 Future work
+Future work shall address the details omitted in this work.
+A more technical paper will define interconnection patterns and their functorial semantics using a categorical framework based on directories and the theory of operads/multicategories.
+As presented here, the definition of (composite) EPHS exhibits a recursive structure.
+It would be interesting to study the ‘closure’ of all EPHS models by deriving representations of EPHS as structured systems of differential-algebraic equations.
+The analysis could benefit further developments and further clarify the relationship with the metriplectic or GENERIC framework as well as port-Hamiltonian systems.
+As an important groundwork for further practical developments, we have implemented the framework within the Julia programming language, see [57].
+The reader is invited to review its application in several examples.
+Based on this, future work can explore topics such as advanced applications (see, e.g., [56, 11]), structure-preserving numerical methods, semi-automated control design, and the integration with scientific machine learning methods.
 Author contribution statement
-Markus Lohmayer: Conceptualization, Investigation, Writing – Original Draft, Writing – Review & Editing, Visualization; Owen Lynch: Investigation, Writing – Original Draft; Sigrid Leyendecker: Supervision
-Acknowledgements
-We thank David Spivak for help when working out the concept of namespaces.
+Markus Lohmayer: Conceptualization, Investigation, Writing – Original Draft, Writing – Review & Editing, Visualization; Owen Lynch: Investigation, Writing – Original Draft; Sigrid Leyendecker: Supervision, Review & Editing, Funding
+Acknowledgments
+We thank David Spivak for help when working out the concept of namespaces. The second author acknowledges support from the DARPA ASKEM program under grant HR00112220038.